A Bibliography of Publications of John W. Tukey

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 June 2024 Version 1.105

Abstract This bibliography records publications of John Wilder Tukey (1915–

2000).

Title word cross-reference

2 [DT68]. 2^{ℵ_{i}} [Max40a]. 2^{n} [ST82]. $49.50 [Won00]. 6 [AT73b]. $69.95 [Zie01]. α [Kel40a]. B [Ars40, Die46, Kel40a, Kel40b, Kel41, Kos40, Max39, Nov39, Otc39, Tuk49h, Tuk57c]. χ^{2} [Ban49, Bañ46, GvdW43, GL46, Ste45b, Tho47, TW49, Tuk57c, Wis47a]. D^{2} [Bos47]. δs [Lia39b]. E_{α} [Kel40a]. F [MPT91]. g [YG19]. H [Kat40, YG19]. HHH [MT00]. HQ [MT00]. HR [MT00]. K [Rid39, Tuk48i, MT50c]. λ [Tuk62f]. t [DT68, GL46, KT88, Wis47a]. Z [Wis47a, Aro47, Mac47].

*39 [Ano01a].

-abgeschlossene [Kat40]. -and- [YG19]. -Distribution [Aro47]. -Kriteriums [GvdW43]. -line [AT73b]. -opérations [Lia39b]. -sample [MT50c, Tuk48i]. -sets [Nov39]. -statistic [Bos47]. -statistics [Rid39].

0 [Won00]. 0-471-38491-7 [Zie01]. 0-691-05782-6 [Won00].

1 [KLTW65b]. 11 [Asp84]. 112 [MT54a, MT54b]. 113 [Tuk55b]. 11th [Tuk52c]. 13-15 [Cra87]. 16th [Gol77]. 1946 [LT46]. 1954 [Mey56]. 1958 [Lan59]. 1962 [Ros63]. 1964 [Boa85, Pri85]. 1976 [Tuk79d, Tuk78f]. 1977 [Efr79]. 1980 [Bar81]. 1983 [You95]. 1984 [Zie91]. 1987 [Cra87]. 1990 [T^{+}90a, T^{+}90b]. 1993 [BCEP94]. 19th [Gol77].

2 [KLTW65c, Tuk84q]. 2000 [Kaf01, McC03, Mos05]. 20th [Cip00]. 24-27 [Bar81]. 2nd [Tuk47j].

3R [BP92, Mal79].

48 [Bri06, Buj06, Efr06, Hub06, Lan06].

596 [Tuk63e, Tuk84g].

6 [Won00]. 61 [AT73a, Ley77]. 63 [Hoa61a]. 64 [Hoa61b]. 65 [Hoa61c].

77V26 [AT73a].

80th [Tuk97a]. 81 [Sim87].

9 [FFT74a, FFT74b, FFT75, FFT88].

A. [Tuk82c]. A.S [Ley77]. abbreviations [TT66]. abbreviato [Lov45]. abgeschlossene [Kat40]. Ability [Tuk47i]. Abraham [PT50]. abstract [Tuk58a, Tuk59f, Tuk68c, Tuk70f, Tuk79e]. Acceptance [F.71, Coo87a, Coo90, Tuk50b]. Access [PCT91, TP98a, TP99a]. account [Tuk95b]. accuracies [Tuk48e, Tuk84d]. Accuracy [Tuk79c, Rao45, Tuk75c, Tuk76d]. achievement [WJT99]. Acid [T^{+}91a]. ACM [Cra87]. Acoustic [ACMT76, CMAT76, ACMT78, TACM77]. Acoustic-Articulatory [ACMT76]. actions [Cla54]. Acute [FSG^{+}78, Tuk78f, TL87]. Adaptations [CHNT93a, CHNT92a]. Additional [Fré48]. additive [Mas41b]. additivity [FRT92, Tuk49p]. Adjective [EHTW88b]. adjectives [EHTW88a]. Adjusting [EKT89]. adjustment [KT93, Tuk69b, Tuk79h, T^{+}90b, Tuk90p, Tuk93c, Tuk86d]. advance [Tuk81c]. Advanced [DGP40, TP40]. Advances [TBBM78, ABH^{+}72, Hin73b, Rob72, Tak73, Hin73a, Hoa74, Law75, Pre75]. advantage [Tuk85h]. Adventures [Lor18]. Adverb [EHTW88b]. Adverbs [EHTW88a]. Advisory [T^{+}62a, TAB^{+}65]. affect [Tuk71d, Tuk72b]. after [Bri06, Buj06, Efr06, Hub06, Lan06, Mal06, To78, Tuk92a]. Age [Tho01, BT78]. agent [GT69a]. Agglomeration [TT81a]. aggregate [MT78]. ahead [Tuk80c, Tuk89]. aid [BT97, TT66]. Air [FSG^{+}78, T^{+}70b, Tuk76b, Tuk78f]. Air-Pollution [FSG^{+}78]. airplane [PT56a, PT56b]. Alamos [Mac91]. alanysis [BHT63]. Albert [Asp84]. algebra [Bal40, Boc38]. Algebraic [Wat10]. algebriche [Cas43, Rom46, Rom47]. Algorithm [CT65, Goo58, CT75, Coo87b, FT73, FT74, HT92, AT73a, Hoa61c, Hoa61a, Hoa61b, Ley77]. Algorithms [AT73a, Cip00, Ley77, CBB^{+}99]. alle [Lov42]. Allowances [Tuk84k]. alphabet [Tuk86a]. alternagraphic [Tuk73g]. alternate [Tuk91c]. Alternative [Tuk86e, Tuk86i, Tuk52a]. alternatives [Tuk93b]. Ambient [FSG^{+}78, Tuk78f]. Ambiguities [ACMT76]. Amer [Sim87]. amounts [Tuk60c, Tuk90l]. Analiz [MT82b, Tuk81a]. Analyse [Stu39]. analyses [GT60, Tuk49d, Tuk51c]. Analysis [AB99, AT63b, BT99, BJM^{+}20, Coo00, De 64, FS48, FFT75, FSG^{+}78, GWD17, Goo58, Hec01, HMT83c, HMT91, HMT00, Hoa03, Hol79, Ile93a, Ile93b, Kan93, MT68, MT77, Olm63, Osb78a, Osb78c, Ros63, Sch93, Spe02, Tal00, Thu47, Tuk47i, Tuk47j, Tuk47k, Tuk65e, Tuk70a, Tuk70b, Tuk71c, Tuk81a, Tuk84a, TT85, Tuk85j, Tuk87e, Tuk87a, Won00, Yat37, Zie88, AT63a, AT73b, BK46, Bar47, BLA18, BT66, BT78, BT84, BT85, BCG^{+}93, BFM97, Bri02b, CHT89, Dan40, DL42a, DL42b, Fer09, Fis54, FFT74a, FFT74b, FFT88, FT73, FT74, FTT80, GGT69, Gol77, HT66, JT87, Jon86a, Jon86b, KT56, LD39, MT82a, MT78, MT90, MT91e, Nai48b, PT56a, PT56b, ST01, Ste45b, Stu39, Tuk48d, Tuk49c, Tuk49g, Tuk53c, Tuk54a, Tuk54c, Tuk61b]. analysis [Tuk61c, Tuk61d, Tuk62c, Tuk63a, Tuk63c, Tuk63d, Tuk63e, Tuk63h, Tuk65b, Tuk65c, TW65, Tuk65d, Tuk65f, TW66, Tuk66c, Tuk67c, Tuk68b, Tuk71a, Tuk72a, Tuk73a, Tuk73b, Tuk73d, Tuk74a, Tuk74b, Tuk74c, Tuk76f, Tuk76g, Tuk76h, Tuk77c, Tuk77f, TBBM78, Tuk78f, TFT79, TM80, Tuk80a, Tuk81c, Tuk82e, Tuk83, Tuk84b, Tuk84c, Tuk84f, Tuk84g, Tuk84p, TV85, Tuk86b, Tuk86f, Tuk88d, TMH91, Tuk93a, Tuk93c, Tuk97b, Hea83, Kem01, Osb78d, Tuk47g, Tuk47b, Col84, Ehr79, Hea84, Osb78b, SG80, Stu84, Zie01, EHTW88a, PT50, Tuk62b, Tuk87b, Tuk88a]. analyst [Tuk78e]. analytical [MT50a]. analytiques [Kos40]. analyzer [Tuk90s]. Analyzing [Tuk69a, Gut45, MT91e, ST82, Tuk51d, Tuk77h]. anatomy [Tuk57d]. Andrews [Hin73a, Hin73b, Hoa74, Law75, Pre75, Rob72, Tak73]. Anecdotes [Buc81, TB84]. angular [FT50, TF49]. annual [Tuk62a, Tuk63g]. anomaly [Tuk90f]. ANOVA [Cox92, Tuk49g, TMH91]. Anscombe [De 64, Olm63, KLTW65a, Tuk67b]. Answer [MT54a, MT54b, Tuk49a, Tuk55b]. Answers [MT49c, PTM47]. Antiaircraft [MBT^{+}45]. Antiquadrature [Tuk57a]. Antithesis [Tuk57b]. apparatus [PKCT95, PT97, TP98a, TP99a]. apparent [Cla54, GT69a]. apparente [Cla54]. appartenance [Kel40a]. appendices [Tuk73f]. Appendix [GT69c, GT69b, GT69a, Tuk59a, Tuk69b, Tuk91a, Tuk83]. Application [CLW67a, LD39, Ste45b, Dan40, DL42a, DL42b, Dwi47, Har46, Mas04, PT56a, PT56b]. Applications [GKT98, Now94, Blu41, Eps48, HR10, MT49a, MT49b, MT50a, MT50b, Tuk68b, Tuk47k, Tuk52b]. Applied [Tuk50f, Bri64, Tuk59a, Tuk59b, Tuk59c, Kem01]. apply [MT91e]. Approach [CPKT92, PCT91, Tuk78a, Tuk65a, Tuk68b, Tuk78c, Tuk84r, Tuk91a]. Approaches [FTT80, TFT79, TW65, Tuk88d]. appropriate [Tuk86d]. Approximate [DT68, GL46, PT65, Que49, Tho47, Tuk48a, Bri64, MT84a, Tuk49b]. Approximation [Ber46, TW46a, Lan59, Tuk49h, TS50]. Approximations [KLTW65c, Tuk55c, Tuk57c, Tuk90o]. April [Asp84, Lan59]. archaeological [BLA18]. Arcsine [Mil78]. arithmétiques [Kel40a]. ARMA [Tuk78c, Tuk78a]. Army [JDC20]. articles [Tuk63f, Tuk63g]. Articulatory [ACMT76, CMAT76, ACMT78, TACM77]. articulatory-to-acoustic [ACMT78, TACM77]. Aspects [BC93, MT82a, TDD^{+}72, TM80, Tuk93b]. assays [MTCH82]. Assessing [MTY91, PT81]. Assessment [GT69c, T^{+}91a, WJT94, Tuk90f]. Assoc [Sim87]. Associated [BCH^{+}86, Bos47, Tuk86c]. Association [KMT99, OT47]. Asymptotic [Tuk48b, BST38, Bri64, WBST38]. Asymptotics [Mas04]. asymptotique [Eyr47a]. Atkinson [Tuk82c]. attainable [Rao45]. attention [WJT94]. attribute [Tuk49b]. Aufgaben [Stu39]. augmented [CT89b]. Aurel [BST38, WBST38]. Author [KLTW65a, Tuk93b]. Authors [RT73, J.77b]. autocovariance [BHT63]. automated [CT89a]. Automatic [CBT98, CT98, PT97]. autour [Cla54]. aux [Cla54]. Average [CT56]. averages [Tuk49t]. ayant [Cla54].

B [Hea84, Sim87, Tuk86k]. B. [Tuk86g]. Background [MT91c]. bad [Tuk86h]. badmandments [Tuk86f]. Bagplot [RRT99]. Balanced [KLTW65b, KLTW65c, Tuk57j, Tuk56c]. band [BT74a]. band-spectroscopic [BT74a]. Based [BCT56, FSG^{+}78, PT65, RK57, CPKT92, EHTW88a, FMT04, PKCT95, TM63, TM75, Tuk78f, Tuk86g, Tuk90c, Tuk95b]. Basford [Coo00, Hec01, Tal00]. Basic [Tuk57e]. basis [Gut45]. batches [HT89]. Bayes [McG11]. Bayesian [Fie06, Fie06]. be [AT73b, CT89a, GMT76, Tuk79g, Tuk84r, Tuk98]. bear [Tuk86f]. Beaton [Tuk90f]. Becker [Tuk93b]. become [Fie06]. Bedford [Tuk74e]. been [Tuk86o]. Before [Coo94b]. Behavior [McN55, CMT54a, DT68, MT78, PT81]. Behavioral [SB01, T^{+}62b, Tuk63c, TDD^{+}72, Tuk80c, Tuk86f]. behaviour [Bri64]. Bell [Ake02, Mal03, Tuk80c]. Berkson [Tuk48h, Tuk58b]. Bessel [Bos47]. Best [Cip00]. Beta [Tho47, TS50, Ban49, TW46a]. Beta-Function [TS50]. beta-functions [Ban49]. Better [Tuk90a, MT84a, Tuk84o]. between [Fes46, HMT50, Mas41a, PT65, Tuk49f, Tuk60c, Tuk61b, Tuk61c, Tuk63d, Tuk65f, Tuk90d, Tuk90l, Wol46]. Beyond [LSZ07, Tuk85f, She07, Tuk62d, Tuk62e, Tuk69b, Tuk85b]. Bias [Que56, Tuk58a, CT02, Tuk86e]. biased [Tuk57a]. Bicentennial [Bla47, LT46]. Bickel [Hin73a, Hin73b, Hoa74, Law75, Pre75, Rob72, Tak73]. bicompacts [ŜB40]. bidec [KT88]. bien [Kur40a]. bifurqués [Eyr40b]. Big [Pét38]. bikompakte [Kat40]. bikompakten [Ste39, ŠB41]. Binomial [MT49d, Fis54]. Biographical [Bri09, Mos05]. Biographies [Wei88]. Biography [Lor18, Mos84, Tuk90b]. Biology [Tuk47j]. biometric [Tuk48c]. Biometrical [T^{+}48]. birthday [Tuk97a]. Bit [TB84]. Bits [BBB59]. Bivariate [RRT99]. Blocks [BT95, Tuk47e, Tuk48f]. Board [T^{+}91a]. bodies [KT56]. bois [Ste47]. Book [AB99, Ano07, Bea86, Boa85, Cha92, Col84, Coo00, Dav88a, Ehr79, Hea84, Hec01, Hin73a, Hol79, Ile93a, J.77a, J.77c, J.77b, Kem01, Law75, Osb78b, Osb78a, Pre75, Pri85, Qua74, Sch93, SG80, Stu84, Tho86, TW46c, Won00, You95, Zie88, Zie89, Zie91, Zie95, Zie01]. Bootstrap [Efr79, Tuk92a]. bootstraps [Tuk87f]. both [Tuk75c, Tuk76d, Tuk80e, Tuk79c]. box [MTL78, TML76, CT89b]. Boxplot [GWD17, RRT99]. Bradley [Tuk92a]. Braun [Zie95, You95]. Breakthroughs [KJ97]. Breeding [BT99, Coo00, Hec01, Tal00, BT97]. brief [Cro01]. Brillinger [AB99, Boa85, Won00, Pri85]. broad [Tuk90p]. Brown [Ros63]. Brownlee [TW46d]. Browsing [CPKT92, PKCT95]. Bryan [MBB^{+}75]. Budne [YKT^{+}59]. Buja [Cha92]. burden [Tuk86f]. Byte [Buc81].

C [Ano07, Bea86, Bri06, Buj06, Col84, Dav88a, Dav88b, Efr06, F.71, Hea83, Hub06, Ile93a, Ile93b, J.77c, J.77b, Kan93, Kem87, Kem01, Lan06, Sch93, Sim87, Stu84, Tho86, Tuk49s, Tuk74d, Tuk82c, Tuk87d, Zie01]. C. [Tuk78b, Tuk85d]. calcolo [Lov45, Odo42, Rol43]. Calcul [Lag65, Ver47b]. calculating [Har46]. Calculation [Bal39, CT65, FS48, Rud66, CT75, Egu46, RK24, Stu39]. calculations [Har46, Tuk67c]. calculus [Lag65]. Call [TB66]. Can [Tuk67e, Tuk80b, AT73b, Tuk63h]. cancer [KT93, KFGT96]. canonique [Kel40a]. Capture [KMT99]. Capture-Recapture [KMT99]. caratteristiche [Lov45]. carcinogens [MBB^{+}75]. cardiac [To78]. Career [Ans03]. Carlo [ABTT56, Mey56, TT56, Tuk85j, TV85, TT90]. Carolina [BCEP94]. cas [Lia39a]. case [AT63a, MT96, MT01b, MT82c, MT84b, Tuk47e, Tuk48f, TM82, Tuk86g, Tuk90e, Vil47]. catalytic [Bal39]. Cathode [PST48]. Cathode-Ray [PST48]. Causation [Tuk54a, Tuk78b]. cause [Tuk76b]. celebration [Tuk97a]. Census [EKT89, Tuk85f, Tuk86d, Tuk85b, T^{+}90a, T^{+}90b]. Centenary [BCEP94]. Center [CHNT93b, CHNT92b]. centers [Tuk81b]. Central [Tuk57c, Egu46, BP92]. centuries [McG11]. Century [Lid01, Gol77, Cip00, Sal01]. cepstrum [BHT63]. cercles [Ste45a]. Certain [Rid39, GT69b, Har46]. certaines [Obr47]. certains [Ars40]. Certainty [TCH85]. Chairman [Tuk50b]. Challenges [Cha98, Cha99, Tuk93c]. Chambers [Hea84]. changes [MTY91, Tuk77h, Tuk90f]. Chapter [GGT69]. Chapters [Tuk63c]. character [BTW95, CBT98, CT98]. characteristics [Tuk49b]. characterization [BT40a]. charting [CT89a]. Chatfield [Tuk85d]. Checking [HT85]. chemical [KT56]. Chemicals [T^{+}73]. Chemistry [Tuk47c, Bal40, T^{+}67]. chi [Ber46, Rea93]. chi-square [Ber46]. chi-squared [Rea93]. chlorofluoromethane [TS76]. choices [MT91b, Tuk93a]. Choosing [Tuk86b]. Chronicle [Lid01]. Churchman [Tuk49s]. Circuits [Tuk61e]. circular [MT96, Tuk91a]. circularly [TT66]. circularly-shifted [TT66]. citable [Tuk63f, Tuk63g]. Citation [J.77a, Tuk62d, Tuk62e, Tuk62a, Tuk63b, Tuk73c]. City [Tuk78f, FSG^{+}78]. Civic [Ans03]. classe [Eyr44, Kel40a]. classes [Kel40a, Kre46, Obr47, Tuk90p]. classical [Ber45, Bos47]. classification [Tuk57i, Tuk57j]. Classifications [KLTW65b, KLTW65c, Tuk90p]. Cleaner [T^{+}70b]. Cleveland [Hea84, Tuk93b, Zie89]. Cliff [EHTW88a]. climb [Tuk73h]. Clinic [T^{+}48]. Clinical [CHNT93a, Tuk78d, CHNT92a, ST91, Tuk77g, Tuk91e, Tuk93e]. Clinical-Trial [CHNT93a]. Clinician [Tuk50c]. closure [Tuk50b]. Cluster [CPKT92, PKCT95]. Cluster-based [CPKT92, PKCT95]. clustering [HT92, NS56]. clusters [TP98a, TP98b, TP99a, TP99b]. CO [ST91]. Cochran [McN55, CMT55]. code [DRT69, McG11]. codes [BTW95]. coefficient [Mas41a, Mas41b]. coefficiente [Lov42]. coefficients [GT69c, Jur47, Ric46, Tol47]. Coined [Leo00]. Colin [Zie91]. Collaboration [Tuk82i]. collaudo [Odo42]. Collected [Boa85, Pri85, You95, Zie88, Zie89, Zie95, Bra94, Bri84a, Bri84b, CT89b, Cle88, Cox92, Jon86a, Jon86b, Mal90, Zie91]. Collections [CPKT92, PKCT95]. collision [Tuk63d, Tuk65f]. color [TH49]. Columbia [Tuk47k]. column [Tuk49j]. Combination [MT82c, MT84b, TM82, Tuk79h, Tuk85g]. combinations [Tuk90a]. combined [Tuk49b]. come [MT91e, Tuk86l, Tuk90c]. comme [ŜB40]. commencement [Tuk74e]. Comment [Sim87, Tuk78a, Tuk85f, Tuk85i, Tuk85j, TV85, JV85, Tuk67d]. Comments [Ans65, BT74b, EHTW88a, Jac65, JP61, KLTW65a, ST91, Tuk76b, Tuk76c, Tuk77a, Tuk78c, Tuk78b, Tuk79a, Tuk84b, Tuk84a, Tuk85b, Tuk85c, Tuk86d, Tuk86c, Tuk86e, Tuk87b, Tuk88a, Wai93, Tuk78e, Tuk80d, Tuk93b]. commercial [Har46]. Commission [RWM^{+}72]. Committee [T^{+}62a, TAB^{+}65]. common [Tuk60c, Tuk90l]. Communication [Sha48]. Communications [BT58a, BT58b, BT59]. commutative [Boc38]. Compact [GWD17, Mor40, Tuk59g]. Compactness [Tuk41]. Company [Leh08]. comparabilité [Die46]. comparative [GT58, HMTW47, Tuk57d]. Comparing [ALTY00, Tuk48d, Tuk49c, Tuk49d, Tuk54b]. Comparison [HP94, LT01, Som93, Tuk50a, Tuk95a]. Comparisons [BC93, BB02, Hay84, KLTW65b, KLTW65c, Wai93, You95, BT83, Bra94, GT69a, Tuk53b, T^{+}70a, Tuk84i, Tuk84l, Tuk91d, Tuk92d, Tuk93b, Tuk93f, WJT94, WJT99, Tuk84m, Zie95]. Compensation [ACMT76]. complet [Die39c, Die39b]. completeness [Coh40]. compléter [Ver47a]. complets [Die39a]. Complex [CT65, GT60, LSZ07, ABTT56, CT75, Mas41a]. Components [Tuk51a, Tuk51b, Tuk56c, Tuk57i, Tuk57j]. composition [KT56]. compound [BPT81, Mac48]. Computation [CLW67a, Cra87, Wat10, Tuk63d, Tuk65f, Tuk71a, Tuk72a]. Computational [Tuk55a]. Computations [Hod58, Tuk56a]. Computer [MACT76, TT85, ACMT78, Buc62, Tho63, TACM77]. Computer-Sorting [MACT76, ACMT78]. Computers [T^{+}67, GK93]. Computing [Cha98, Cha99, GKT98, Nas90, Sin67, BGvN46, Tuk65a, Tuk68a, Tuk71d, Tuk72b, Tuk80a, Tuk86m, Cha92]. concentrate [FTT80, TFT79]. concept [Blu41]. Concepts [Cha98, Cha99, TMH91, Pea55]. Conclusions [Tuk60a, GT69a, T^{+}78]. Concrete [Tuk58e]. Conditional [TT56]. Conference [Bla47, BCEP94, Cra87, LT46, Bar81]. Confidence [Tuk49e, Tuk84o, Wol46, Bri64, HT91, Tuk49w, Tuk58a, TM63, Tuk81b, Tuk90r]. Configural [MT91d, Ros83, Tuk87c, PT81, Tuk81b, Tuk81c]. confirmatory [Tuk80e]. Conjecture [Hay84]. Conjoint [LT64]. connection [Tuk61b, Tuk61c, JK92]. Connections [Mil97]. Conservation [GWD17, T^{+}78]. conservatism [Som93]. Conservative [Hay84]. considerations [GT69a, Tuk77h]. constant [Tol47]. constructed [Tuk82c]. constructing [Fré46, Fré48]. Constructive [Mil97]. Consultants [Tuk55a]. Consumer [Tuk91b]. contact [Tuk62d, Tuk62e]. contains [Tuk65j]. contaminated [Tuk49v, Tuk59i, Tuk60d]. contamination [HT49]. contemporary [Fer09]. Continu [Eyr47c, Eyr43, Eyr47b]. continuous [HR08, Tuk47e]. continuum [Max40a]. contrails [Tuk82f]. Contribution [Bañ46, Pét38]. Contributions [BB02, Fis50, Hub02, Spe02, Bri02a]. Contributor [Ans03]. Control [CT89a, Tuk82b, Tuk88b, Tuk88c, TM75, Tuk90p]. Control-charting [CT89a]. Controlled [Tuk78d]. Controlling [Tuk95a, WJT94, WJT99, LT01]. controversy [McG11]. Convergence [Tuk40, Ber45, Ste41]. convergent [Sam45]. Conversation [Ans88, FM00, FM03, FM97]. conversion [AT59]. converting [BTW95]. convex [Tuk42]. Convolution [CLW67a]. convolutions [BST38, WBST38]. Cooley [BJM^{+}20, JK92]. coordinates [Tuk90a]. Cornelius [BCEP94]. corner [OT47]. Corporation [Tuk55d]. Correction [De 64, Olm63, Tuk62b, BT40b]. Correlated [KLTW66]. Correlation [KLTW66, Mas41a, ETW71, Fré46, Fré48, Mas41b, Que49, Sim46, TEW72, Vas47]. correlazione [Lov42, Lov45]. Correspondence [BLA18]. Corrigenda [ST61]. Costs [ZM08]. count [CHT00, TCH00]. Counted [MT49b, Tuk88e]. counties [KFGT96]. counts [CHT00, TCH00]. Course [MT77]. Covariance [TT90]. covariates [Tuk91e]. coverage [T^{+}90a, T^{+}91b]. cracked [McG11]. cracking [BHT63]. Criterion [BCT56, Ber45]. criticism [Tuk77d, Tuk90q]. cross [BHT63]. cross-cepstrum [BHT63]. crucial [Tuk85g]. cubature [Ste47, Tuk91a]. Cult [ZM08]. Cum [Qua74]. CumIndex [DT73a, DT73b]. Cumulants [Aro47, Mac47, Tuk51b, Wis47a]. curly [Tuk82f]. Curves [PT65, Tuk60b, Tuk61a, Mol46, Tor39]. Cut [KLTW65b, KLTW65c].

D [AB99, Bea86, Dav88a, Dav88b, Hea83, Hin73a, Hin73b, Hoa74, Ile93a, Ile93b, Kan93, Kem87, Kem01, Law75, Pre75, Pri85, Rob72, Sim87, Tak73, Tuk63c, Tuk86j, Won00]. d’être [Die39b]. Daily [FSG^{+}78, Tuk78f]. damit [Kno43]. Daniel [KFT^{+}60]. Dannykh [MT82b]. dans [Lus40a, Max39, Max40c]. Data [Ake02, AB99, Ano07, Bar81, BT99, Bea86, BBB59, Cha98, Cha99, CHNT93a, Col84, Coo00, Dav88a, Dav88b, Don17, Ehr79, EHTW88b, FFT75, Hea83, Hea84, Hec01, HMT83c, HMT85, HMT00, Hoa03, Hol79, Kem87, Kem01, MT68, MT77, Osb78b, Osb78c, Osb78d, Par79, SG80, Stu84, Tal00, Tho86, Tuk47k, Tuk61e, Tuk63c, Tuk65b, Tuk65c, TW65, Tuk65e, TW66, Tuk70a, Tuk70b, Tuk71a, Tuk71c, Tuk72a, Tuk73a, Tuk74a, TT81a, Tuk84c, TT85, Tuk86f, Tuk88d, Tuk90c, Won00, Zie88, Zie01, AT73b, BT74a, BLA18, BT66, BFM97, CHNT92a, CHT00, Fer09, FFT74a, FFT74b, FFT88, FT73, FT74, FTT80, HT66, HMT06, JT87, Jon86a, Jon86b, MT82a, MT90, MT49b, MT91e, Tuk51d, Tuk61d, Tuk62b, Tuk62c, Tuk63a, Tuk63d, Tuk63h, Tuk65d, Tuk65f, Tuk69a, Tuk73b, Tuk73d, Tuk74c]. data [Tuk75b, Tuk77c, Tuk77f, TBBM78, Tuk78e, TFT79, Tuk79a, TM80, Tuk80a, TT81b, Tuk81c, Tuk82e, Tuk82g, TT82, Tuk84f, Tuk84h, Tuk86b, Tuk87b, Tuk87f, Tuk88a, Tuk88f, Tuk95b, Tuk97b, TCH00, Tuk85d, Osb78a]. Data-based [Tuk90c, Tuk95b]. Data-driven [TT81a]. data-modification [Tuk88f]. Datasets [LSZ07]. datesware [Tuk91b]. David [Ano07, Boa85, Col84, Sch93, Stu84, Tho86, Zie01, Tuk49r]. Davies [Tuk95b]. death [KT93, To78]. decades [Tuk90c]. December [BCEP94, LT46]. Decision [Kem01, RK57]. Decisions [Tuk76a, Tuk60a, Tuk76c]. Deconstructing [NGL^{+}94]. Deduction [Gau66]. degree [Tuk49p, Tuk90k]. del [Lov42]. dell’ortogonalizzazione [Rom46, Rom47]. Demographic [MT75]. Démonstration [Kel40a]. dénombrables [Die46]. d’ensembles [Kel40a, Kra39, Kur40a, Max40b, Max40c, Otc39]. Denumerability [Tuk39b]. denumerable [BS39]. depth [DGC11, HR08, Mas04]. derivations [Tuk86l, KLTW65c]. Deriving [Tuk86d]. Design [Yat37, BGvN46, Tuk49j, Tuk53a]. designs [Tuk51c, Tuk56c]. d’espace [Die39c, Die39b]. detailed [Tuk85h]. Detecting [BTW95]. detection [HIT81, THI81, Tuk90k]. detective [Tuk69a]. Determination [Tuk49f, Tuk90d, Tuk59a, Tuk59b, Tuk59c]. déterminés [Ver47a]. determining [Jur47]. Development [Coo94a, GT69b, HT49, Mac91, Now94, Thu47, Tuk47k, Sle74]. développements [Ver47a]. Deviate [Nai48a]. Deviates [KLTW66, Fin46, Lev47, Tuk55d]. deviation [Tuk50a]. Deviations [PT65, Tuk46]. Device [PST48, Tol47]. Diagnosis [MT75, Tuk48c]. diagnostics [Tuk82c]. Diagrams [Ber83, Ber10]. dialing [TM75]. did [Fie06]. Difference [GWD17, Tuk50f, Che43, Fes46]. differences [CT89a, Tuk69b, WJT99]. different [Lag65]. differential [Tol47]. differentiation [Tuk58c]. Différents [Lag65]. difficulty [Gar69]. Digits [Tuk55d]. dimension [Mor40]. dimensional [FMT06, FFT75, FTT80, TFT79, Tuk82g, TT82]. dimensions [TT81b].

Dipole [TM47]. directe [Kel40a]. directed [Tuk81c]. direction [MT96]. Directions [Tuk85i, BCG^{+}93, Tuk95a, Tuk85c]. dirty [Tuk51c]. discontinuous [Tuk48f]. discontinus [ŜB40]. Discoverer [Wai03]. discoveries [Tuk95a]. Discovery [Coo94b, Coo87b, LT01]. discrete [HT85]. discrimination [Pen47, Smi47]. discussed [Tuk87f]. Discussion [BC93, BCH^{+}86, Bri06, Buj06, Efr06, FSG^{+}78, GH93, HBK^{+}78, Hub06, KFT^{+}60, Lan06, NGL^{+}94, SGT^{+}78, Tuk47c, Tuk50c, Tuk61b, Tuk61c, Tuk67b, Tuk67a, Tuk77b, Tuk78f, Tuk82c, Tuk82d, Tuk84l, Tuk85d, Tuk85e, Tuk85a, Tuk86g, Tuk86h, Tuk86i, Tuk86j, Tuk86k, Tuk87d, Tuk87e, Tuk87a, Tuk92a, YKT^{+}59, BGvN46, TGH^{+}xx]. Discussions [JP61, Pon38]. Disentangling [Tuk90f]. dispersion [dTP46]. Display [FFT75, FFT74a, FFT74b, FFT88, TT81b, Tuk82b, Tuk88b, Tuk88c, Tuk88f, Tuk90c]. Displays [ALTY00, Tuk70f, Tuk72c, Tuk73g, Tuk91c]. distance [Blu41]. Distances [PT65]. distributed [Tuk92b]. Distribution [Aro47, Nai48a, Tho47, Tuk57c, Ban49, Bos47, BT78, DT68, Fes46, Fin46, HMT50, HR08, Jon48, Lev47, dTP46, Tuk38, TW46a, Tuk49q, Tuk90s, YG19]. Distributions [JK73, JR71, MT75, BST38, BT46, HT85, Mac48, MT00, RT72, Ste42, Tuk49m, Tuk49n, Tuk49v, Tuk52a, Tuk59i, Tuk60d, Tuk62f, WBST38, Wis47a]. divisibilité [Eyr47a]. division [Ste45a]. Do [Tuk60e, Tuk71b, Tuk86l, Tuk67e]. Document [CPKT92, BTW95, CBT98, CT98, PKCT95, PT97, TP98b, TP99b]. Dodge [Tuk79d, Tuk79b]. Does [JDC20, Tuk50f, Tuk76b]. doivent [Ver47a]. Dolby [Qua74]. Doses [TCH85]. Double [KLTW65b, KLTW65c, Mil78, MT83, Tuk85g]. doubling [MT00]. down [McG11]. driven [TT81a]. drop [CT98]. Drug [TCH85]. Duckworth [Tuk59g]. d’un [Kel40a, Kre46, Lus40b]. d’une [Die39c, Die39b, Ver47b, Eyr40a]. Dyadic [Tuk49g]. Dynamic [CM88, Tuk87b, Tuk88a, Tuk88f]. dynamics [PT56a, PT56b, Tuk47a, Wie56].

each [Tuk71d, Tuk72b]. Earth [Cla54, TB66]. easy [Tuk53c]. echoes [BHT63]. ed [Odo42, Tuk47j, Tuk52c]. edited [Zie01]. Edition [Tuk47i]. Editors [Cip00]. eds [Won00]. education [BMTW49, T^{+}67]. Educational [WJT94, WJO95, WJT99]. EEG [Tuk68b]. effect [Tuk90f, Tuk90k]. effective [Egu46]. Effects [FSG^{+}78, TS76, Tuk78f]. Efficacy [CHT00, TCH00]. efficiency [Tuk49a]. Efficient [AT59, AT63a]. effort [Tuk78g]. Efron [Tuk92a]. égard [Cla54]. eigenvectors [MT96]. elections [MT78]. electrical [Tol47]. electronic [BGvN46]. élément [Kel40a]. elementary [Tuk48j, Tuk48k]. Elizabeth [FM00, GH93]. elongation [MT96]. Emanuel [Tuk79a]. emerged [McG11]. Emeritus [Haf00]. emphasizing [MT82a, Tuk61b, Tuk61c, TM80]. Empirical [MT75, BT78, Run05, Tuk49h, Tuk83]. empirically [Run03]. empirically-given [Run03]. empirisch [Kno43, Run03]. empirischen [Run05]. employing [PHC^{+}94, TP98a, TP99a]. Encyclopedia [SB01]. Energy [T^{+}78]. Engineering [BT58a, BT58b, Tuk52b, Tuk55a, BT59, T^{+}62a]. Engineers [TW52]. Enhanced [GWD17]. ENIAC [TB84]. Enigma [McG11]. enquiry [MT84a]. Ensembles [Eyr47c, Ars40, Die46, Eyr40a, Kel40b, Kel41, Kon40, Kos40, Kur40b, Kur41, Max39, Ste45a]. entitled [Bar81]. Entomological [T^{+}48]. environment [TAB^{+}65]. environmental [TS76, TAB^{+}65]. EPA [ST91]. Equality [Ram56]. Equalization [Tuk59b, Tuk59c, Tuk59a]. Equalizing [Tuk48e, Tuk84d]. equations [Ber45, Ham46, Tol47]. equazioni [Cas43, Rom46, Rom47]. equilibria [Bal39]. équivalence [Otc39]. equivalent [Tuk47e, Tuk48f]. Errata [LLT54]. Error [BJM^{+}20, ZM08, Tuk84k, WJT94, WJT99]. errors [Tuk57e, Tuk58c, Tuk92a]. Erweiterungen [Fom40]. espace [Lus40b, Max39, Max40c, Ste45a]. Espaces [ŜB40, Ale40, Die39b, Die39a, Lus40a]. especially [Tuk77g, Tuk85g]. Essays [AB99, Won00, BFM97]. Estadistica [Tuk47h]. Estatística [Kin45]. estimate [Tuk57a]. Estimated [TM46]. Estimates [Hin73a, Hin73b, Hoa74, Law75, Pre75, Rob72, Tak73, ABH^{+}72, Bri64, CHT00, MPT91, PT81, Tuk49a, Tuk50d, Tuk57f, Tuk81c, Tuk90j, TCH00]. Estimating [Tuk85b, Tuk85f]. Estimation [Hub64, Que56, Tuk52a, BGT67, Hal46, Jef46, MT91b, Rao45, ST45, Tuk47e, Tuk48f, Tuk59d, Tuk60b, Tuk61a, TB82, Vil47]. estimators [GT73, HMT83b, Tuk58b, HMT83a]. evaluating [Tuk86e, Tuk86i]. Evaluation [CHNT92a, CHNT93a, Tuk84a, GMT76, T^{+}90a, T^{+}91b, Tuk86g]. events [HMT50]. Evidence [FSG^{+}78, Tuk78f, ST91]. evolution [Tuk88f]. Exact [Jon48]. Examination [AT63b, De 64, Olm63, Tuk85d]. example [Gar69]. examples [BT84, BT85, Smi47, Tuk57g, Tuk74b, TMH91, Tuk93a, WJT99]. Exbrids [Tuk92b]. Excellence [Lid01]. exemple [Die39c]. exemples [Kel40a]. exercise [Tuk58c]. exercises [Stu39]. Expansion [Thu47]. expectations [Egu46, Tuk48b]. experience [T^{+}91a]. Experimental [Tuk47c, Tuk49s, Tuk54d, TT90, FRT92, Tuk53a]. Experimentation [TW46d]. Experiments [Yat37, BT97, ST82, Tuk85g]. Expert [Tuk85i, Tuk85c, Tuk86a]. Explaratory [Ile93a, Ile93b]. exploration [Tuk74d]. Exploratory [Col84, Ehr79, Hea83, HMT83c, HMT91, HMT00, Kan93, Kem01, MT78, Osb78d, Sch93, Stu84, Tuk70a, Tuk70b, Tuk71c, Tuk73b, Tuk77c, TT85, Tuk93a, Zie01, FT73, FT74, JT87, MT82a, TM80, Tuk80e, Osb78b, SG80]. Exploring [Dav88b, EHTW88b, HMT85, HMT06, Kem87, Ano07, Bea86, Dav88a, Tho86]. exponential [Tuk49t, Vil47]. exponentially [Tuk92b]. expository [TW66]. expression [Tuk79e, Tuk82h, Tuk90h]. expressions [Tuk92b]. Extended [JK73]. Extension [Kre46]. Extrapolation [Tuk52b]. Extreme [Nai48a, Nai48b].

F [Bea86, Dav88a, Dav88b, De 64, F.71, Hea83, Hin73a, Hin73b, Hoa74, Ile93a, Ile93b, Kan93, Kem87, Kem01, Law75, McN55, Olm63, Osb78a, Osb78c, Pre75, Rob72, Tak73, Tuk49r, Tuk63c, Tuk79b, Tuk79d]. faceless [Tuk74d]. Factor [Thu47, Tuk47g, Tuk47b]. Factorial [Tuk47i, Yat37, ST82, Cox92]. factorials [CT56, Tuk59f]. false [LT01, Tuk95a]. familles [Kur40a, Otc39]. family [Tuk52a]. Fast [BJM^{+}20, Sin67, AT73b, CLW67a, CLW67b, Coo87b, Gar69, GS66, HJB84, HJB85, Hua71, Sch12]. faster [Sch12]. Federal [RWM^{+}72, Tuk49k, Tuk49l, Tuk67e, T^{+}71]. females [KFGT96]. fermés [Eyr40a]. Fernholz [AB99, Won00]. ferromagnetism [ST49]. fertility [BT78]. FFT [Coo87a, Coo90, CT93, Coo94a, Coo94b, GK93, JK92, Now94]. fiducial [Tuk57g, Tuk90m]. fields [ST49]. financial [VVT16]. Find [BT95, Hoa61c]. Fine [LT46]. Finite [GWD17, Tuk49i, Tuk51b, Rid39, Tuk90e]. First [Tuk70c, Tuk84e, Tuk74b]. Fisher [Aro47, Mac47, Rid39, Tuk49h, Tuk52c, Tuk57c]. Fission [Whe62]. fit [Rea93]. fits [Tuk91c]. Fitted [Tuk85j, Tuk67b, TV85]. Fitting [MT00, BT74b, BT74a]. flexible [Tuk82b, Tuk88b, Tuk88c]. Florida [Mey56]. flow [Tuk47d]. fluctuations [Tuk57e, Tuk58c]. fluid [Tuk47a]. focused [Tuk80d]. following [TDS50]. fonctionnel [Lus40b]. fonctions [Obr47]. foray [Cro01]. Force [T^{+}70b]. forests [Tuk90i]. Foreword [Cas03, Tuk90f]. forgetting [Tuk86o]. Form [Nai48a, Jef46, MT91e, Nai48b, Tol47]. formation [PHC^{+}94]. formula [Che43, ST44]. Formulae [Tho47, Bos47]. formulas [GT69b, GL46, Tuk57a, Tuk57e]. formulation [JT00, JT01, TJ00]. forthcoming [ST91]. Foundation [Tuk63g, Stu37]. foundations [Tuk97b]. four [Tuk82g, TT82]. four-dimensional [Tuk82g, TT82]. Fourier [BJM^{+}20, CLW67a, CLW67b, Coo87b, Gar69, GS66, HJB84, HJB85, Hua71, Sch12, BT66, CT65, CLW67a, CT75, Dan40, DL42a, DL42b, Goo58, HR10, Rud66, Sin67, Tuk84f]. fraction [Wol46]. fractional [Tuk38]. fractions [Tuk60c, Tuk90l]. Frank [Tuk82i]. Fréchet [Sim46]. Frederick [Ano07, Col84, Sch93, Stu84, Tho86, Zie01, Ans88, Hol79, Tuk90b]. freedom [Tuk49p, Tuk90k]. Freeman [BLA18, Mil78, Rea93, SC00]. French [Cla54, Lag59, Lag65]. Frequency [PT65, BT66, Fes46, Fin46, Tuk63e, Tuk68b, Tuk84f, Tuk84g]. frontiers [Tuk65b, Tuk65c]. Fun [GS66]. Function [Pon38, Tho47, TS50, BTW95, Bos47, BPT81, CT02, Fes46, Run05]. Function-Theoretical [Pon38]. functional [Ham46]. functionals [BT38, BT40b]. Functions [MACT76, Ban49, Rid39, Run03, Tuk92a, BL01]. Fundamental [LT64, Tuk59a, Mol46]. Fundamentals [HMT91, Ile93b, Kan93, Ile93a, Sch93]. Fundamentalsatz [Ste39]. Fundamentos [CMT56]. Funktion [Run05]. Funktionen [Run03]. Further [TBS77, Tuk69b, Tuk70e, Tuk74b, Tuk90n]. Future [Tuk65e, Tuk85c, Tuk85i, Tuk95a, MT90, Tuk61d, Tuk62b, Tuk62c, Tuk65d, Tuk71b, Tuk74f, Tuk82a].

G [McN55, Tuk63c]. gained [Coo87a, Coo90]. gaining [Gar69]. galactic [ST49]. Galaxies [McV56, NS56]. Galois [Mil97]. Game [Pét38]. gather [PKCT95, CPKT92]. gathering [MT50b]. Gauss [HJB84, HJB85, Tuk75a]. gegebenen [Kno43]. gegebener [Run03]. General [FSG^{+}78, GT69a, SGT^{+}78, Bri64, Fré46, Fré48, GT60, Tuk41, Tuk90o, AT63a]. generalist [BMTW49]. generalization [Cot40]. Generalized [ST42, Mac48, Tuk57e]. generating [CBT98]. genetics [Cro01, TDD^{+}72]. geometry [Bal39, Blu41]. geophysical [HT66]. Geophysics [Now94, Tuk65b, Tuk65c, Tuk66c]. German [Gau66, Run03, Run05, RK24, Stu37, Stu39]. German/Latin [Gau66]. Germany [Wat10]. Given [ALTY00, Run03, TP99b, Wol46]. giving [Tol47]. Glynn [Tuk86j]. GN [Tuk63g]. Go [Tuk60e, Tuk80b, Tuk92d, Tuk93f]. Godfrey [Tuk47i]. going [Tuk98]. Goodman [JP61]. goodness [Rea93]. goodness-of-fit [Rea93]. Got [Hua71]. Government [Tuk49k, Tuk49l, Tuk67e]. Graduate [JDC20, T^{+}62a, Tuk82d]. grafico [Rol43]. Granger [Tuk78b]. Grant [Tuk63g]. Graphic [BC93, Tuk93b, Tuk70f, Tuk72c, Tuk88b, Tuk88c]. Graphical [BT97, BT99, Coo00, HP94, JT87, Tal00, TT81b, Tuk91c, Wai90, CHT89, MT49b, Tuk82b, Hea84, Hec01]. Graphics [Ber83, Ber10, Cha92, FS02, TT85, Zie89, CM88, Tuk82g, TT82, Tuk87b, Tuk88a, Tuk88f, Tuk90c, Cle88]. Graphing [Wai93]. Group [Tuk47k, Tuk49m, Tuk49n, T^{+}70a, Ake02]. Groups [Tuk85j, TV85]. growth [Tuk53a, Tuk74f]. Grundlagen [Stu37]. Guided [MBT^{+}45, Tuk90h]. guidelines [T^{+}90b]. guiding [Tuk82h]. günstigste [Wis47b].

H [Hin73a, Hin73b, Hoa74, Kem01, Law75, Pre75, Rob72, Tak73, Tuk47i, Tuk63c, You95, Zie95]. H. [Tuk86h]. half [DGC11]. half-space [DGC11]. Hall [LT46]. Halocarbons [TS76]. Halothane [T^{+}66]. Hampei [Pre75, Rob72]. Hampel [Hin73a, Hin73b, Hoa74, Law75, Tak73]. Hampshire [MT78]. Handbook [BL01]. handed [Tuk82b, Tuk88b, Tuk88c]. Handout [FRT92]. Handouts [Tuk90g]. harmonic [Stu39]. Harmonischen [Stu39]. Harold [Tuk79b, Tuk79d]. Hartley [Ram56]. Health [FSG^{+}78, T^{+}73, Tuk76a, Tuk76c, Tuk78f]. hearing [T^{+}91b]. heaven [Tuk86l]. Heckman [Tuk86e, Tuk86i]. held [Bar81, Mey56, Ros63]. Henry [Tuk84l]. Herbert [Tuk78f]. Here [Tuk60e]. Heterogeneity [KMT99, BK46]. Heterogeneous [CDT93]. heuristics [HT89]. High [FMT06, Tuk74e, Tuk81c]. High-dimensional [FMT06]. Higher [MT75, Tuk77d, BT78, Tol47, Tuk53d, TFT79, TT81b]. higher-dimensional [TFT79]. Higher-Order [MT75]. higher-rank [BT78]. Highlights [T^{+}78]. Historical [CLW67b]. History [Cra87, Nas90, Tuk84c, Asp84, CBB^{+}99, Gol77, HJB84, HJB85]. Hoaglin [Ano07, Bea86, Col84, Dav88a, Dav88b, Hea83, Ile93a, Ile93b, Kan93, Kem87, Kem01, Sch93, Sim87, Stu84, Tho86, Zie01, HIT87]. Homogeneity [TM46]. homogeneous [Tol47]. homogènes [Kel40b]. honest [Tuk95b, Tuk97b].

Honor [AB99, Won00, BFM97, Fer09]. Hospital [Tuk47l]. Housing [EKT89, T^{+}78]. Howard [Tuk93b]. HSNC’87 [Cra87]. Huber [Hin73a, Hin73b, Hoa74, Law75, Pre75, Rob72, Tak73]. Human [McN55, Tuk47i, CMT54a, TDD^{+}72]. humor [Haf00]. hunted [McG11]. Hunting [Pét38]. Hybrid [GWD17]. Hyperbolic [GWD17]. hypothèse [Kra39]. Hypothesis [JLT01].

I. [TV85]. idea [Tuk86h]. ideas [MT91a]. identification [TT66, TM47, Tuk81b, TB82]. identifying [CT98]. Iglewicz [Sim87, HIT87]. Ignored [KMT99]. II [BT58b, BJT78, Bri84b, FSG^{+}78, MT49b, MT84b, SGT^{+}78, Tuk47e, Tuk49l, Tuk51c, Tuk57i, T^{+}71]. III [Jon86a, MT50a, Tuk48f, Tuk57j, Tuk77b]. Illustrated [BT99, Coo00, Hec01, MT75, Tal00, BT74a, JT87]. illustrations [T^{+}70a]. image [CBT98, CT98]. images [NS56, ŜB40]. immediate [Tuk82b, Tuk88b, Tuk88c]. immensed [Tuk54c]. impact [Tuk86e, Tuk86i]. implications [Tuk74f, Tuk80a, Tuk78b]. importance [Tuk48j, Tuk48k]. important [Tuk77h]. Improved [LT01, MBB^{+}75]. Improvement [Dan40, Tuk73h, Tuk79h]. improvements [DL42a, DL42b]. Improving [Tuk85g]. in-place [HR10]. inadequacy [Tuk90p]. include [CHT00, TCH00]. Including [MT68, AT73b, MTCH82, Per47]. Incomplete [Tho47, Ban49]. Increasing [TCH85]. Index [DT73a, RT73, RT75, Tuk73c, Sim46, Tuk62a, Tuk63b, J.77a, Qua74, J.77a, J.77c, J.77b]. indexes [TT66]. indicator [Tuk86h]. indices [Fré46, Fré48, Tuk62d, Tuk62e, dTP46]. indifference [Per47]. individual [Tuk48d, Tuk49c, Tuk77d]. Inducão [Kin45, Tuk47h]. induction [Fis55]. Industrial [TW46b, Tuk47c, TW46c, TW46d]. inequality [Gut48, Tuk46]. inevitable [Tuk63d, Tuk65f]. infarction [To78, TL87]. Inference [MT96, Fie06, Tuk95b, Tuk49s]. Infinite [Tuk47f, BST38, WBST38]. Influence [Mac91, CT02, Tuk92a]. Information [Rao45, Tuk63g, AT59, AT63a, MT50b, PHC^{+}94, Sle74, Tuk62a, Tuk65j, TP98a, TP99a]. Ingenieure [TW52]. inhomogeneous [Tol47]. initial [Tuk59a, Tuk59b, Tuk59c, Tuk73h, Tuk74d, Tuk85d]. inputs [Tuk70e]. inquiry [Tuk69b]. insensitive [HT49]. Instead [Tuk75a]. institutional [Tuk69b]. instrument [BGvN46]. instruments [CT89a]. Intégral [Lag65, Cot40, Dwi47, Lag65]. Integrals [CLW67a, Kan39]. integration [FMT06]. Intelligent [Tuk85i, Tuk85c]. Interaction [CHNT93b, Goo58, Tuk49j, CHNT92b]. Interactive [FFT75, FS02, FFT74a, FFT74b, FFT88]. Intercomparing [BT95]. Interdisciplinary [BT99, Hec01]. interests [TP98b, TP99b]. interface [Tuk86m]. Interim [Tuk77e]. intermediate [Bal39, FTT80]. intermediate-dimensional [FTT80]. International [BCEP94, SB01, Wat10]. Interpolation [Gau66, Tuk52b, Ban58, LD39, Ric46]. Interpolationis [Gau66]. Interpolations [Tuk55c]. interpretation [Tuk78b]. Interpreting [Bar81].

Intersections [Lus40a]. Interstellar [ST49, ST51]. intervals [Tuk81b]. intervention [Tuk86i]. Interview [AT85, Asp84]. intriguing [DGC11]. intrinsic [Tuk39a]. Introduction [HMT83b, Tuk73d, Tuk74c, Tuk82e, HMT83a, TT85, Tuk90h, PT81, Tuk59e, Tuk63e, Tuk67c, Tuk84g]. invariant [Jef46]. Inventor [Wai03]. Inventory [LLT54]. Inverse [LSZ07, Mil78, Ban58, Per47]. Inversion [ACMT78, CMAT76, TACM77]. invertibility [Tuk56b, Tuk58d]. Inverting [MACT76]. investigation [Tuk67b]. inviting [MT91e]. Involved [BJM^{+}20, Tuk81c]. ionization [Tuk37]. ISBN [Won00, Zie01]. ISSAC [Wat10]. issue [Ano01b, Fer09]. Issues [Tuk95b, BT84, BT85, Tuk78e]. items [Tuk54b, Tuk63f, Tuk63g]. iteration [Ham46, Tuk48g]. Iterative [PHC^{+}94, Ber45, Sam45]. IV [Jon86b, MT50b].

J [Bea86, Coo00, Dav88a, Dav88b, De 64, F.71, Hea83, Hea84, Hec01, Hin73a, Hin73b, Hoa74, Ile93a, Ile93b, J.77c, J.77b, Kan93, Kem87, Kem01, Law75, McN55, Olm63, Pre75, Rob72, Ros83, Sim87, Tak73, Tal00, Tuk47h, Tuk63c, Tuk78b]. J. [Ste41, Tuk86e, Tuk86i]. Jackknife [Efr79, Mil74, Tuk92a, Bri64, MTCH82]. Jackknife-after-bootstrap [Tuk92a]. jackknives [Tuk87f]. James [Qua74, Tuk84j]. January [Tuk47a]. Jersey [Cra87]. Jobs [ZM08]. John [AB99, Ano07, BC93, Boa85, Col84, De 64, GH93, Hoa02, J.77a, Olm63, Osb78b, Osb78a, Osb78c, Osb78d, Pri85, Qua74, SG80, Stu84, Tho86, Won00, You95, Zie88, Zie89, Zie91, Zie95, Zie01, Ake02, Ano01a, Ano02, Ans88, Ans03, Asp84, AT85, BB02, Bla47, Bra94, Bri84a, Bri84b, BFM97, Bri02a, Bri02b, Bri02c, Bri09, Cas03, Cle88, Cox92, Dem02, Ehr79, FM97, FM00, FM03, Fer03, Fer09, FS02, Har03, Hoa03, Hol79, Hub02, JDC20, JV85, Jon86a, Jon86b, Kaf01, Kaf03, Leo00, Lor18, Mal90, Mal03, McC03, Mor03, Mos84, Mos05, Spe02, TGH^{+}xx, Wai90, Wai03]. Johnson [JT87]. Johnston [Tuk85j]. Johnstone [TV85]. Jones [Tuk87d, Zie88]. Jordan [Tor39]. journal [Tuk55e, Tuk63f, Tuk63g]. judgment [GT58]. July [Kaf01, McC03, Mos05, Wat10]. June [Kaf01, McC03, Mos05, Ros63]. Justice [ZM08]. JWT [Tuk97a].

Keeping [Tuk56a, Tuk62d, Tuk62e, Hod58]. Kept [Leh08]. Key [MT91a, Sle74, Tuk85h]. Kinds [Tuk81b, Tuk87f]. kinetics [Bal39]. Kingston [Tuk47h]. Kinsey [CMT55, McN55, CMT53, CMT54a]. kit [HT89]. kleine [GvdW43]. Kleiner [Hea84]. knowledge [TP98b, TP99b]. Korrekturformeln [Kno43]. Kramer [Hay84, Som93]. Kriteriums [GvdW43]. kurtosis [Gut48]. Kurtz [Ans65, Jac65].

L [AB99, Qua74, Tuk47g, Tuk47b, Won00]. labeling [HIT86, Sim87]. Laboratories [Ake02, Mac91, Tuk80c]. laboratory [CT89a, Tuk53a]. Labs [Mal03]. Lack [TDS50]. Lady [Sal01]. Lagrangian [Ric46]. Lags [Tuk71e, Tuk73e]. Laird [Tuk86j]. Lambda [JR71, JK73]. Lanczos [BCEP94, Coo94b]. Large [CPKT92, HT49, PKCT95, Tuk49y, Tuk58a, Tuk73b, Tuk78g, Tuk86m]. large-sample [HT49]. larger [MT96]. Latin [Gau66]. Laurie [Tuk95b]. law [GT58]. Laws [GWD17]. learning [Tuk86f]. least [Tuk67b, Tuk75a, Tuk90j]. Lebesgue [Cot40]. Lecons [Eyr47c]. Lecture [Tuk73f, Tuk70c, Tuk70d, Tuk84e, Tuk84n, Efr79]. Lectures [BST38, Boc38, WBST38, Tuk90g, RK24]. legacy [T^{+}91a, VVT16]. Less [TM63, Tuk57a, Tuk74b]. levels [MT50c, Ric46, Tuk48i]. librement [Ver47a]. lies [Wol46]. life [Bri02a]. light [ST91, Tuk95b]. Like [Hod58, Tuk56a, TA71]. likelihood [Bri64, Tuk49a]. Limit [BP92]. limite [Rol43, Ver47b]. Limited [Tuk85h]. limités [Ver47a]. limits [Bri64, ST44, Tuk49e, Wol46]. Linder [TW52]. Line [AT73a, Ley77, Tuk85j, AT73b, JV85, MT01b, TV85]. linéaires [Kur40a, Lus40a]. linear [Ber45, BT38, BT40b, Tol47, Tuk49f, Tuk67b, Tuk77a, Tuk90d]. lineari [Rom46]. Linearization [Tuk47d]. Lines [Tuk85j, Fin46, TV85]. Link [Ans65, Jac65]. Linked [BC93, Tuk93b]. Linking [VVT16]. liquids [Dan40, DL42a, DL42b]. literature [Tuk62d, Tuk62e]. lithologic [KT56]. Little [Tuk59f]. lively [Haf00]. Livermore [Mac91]. Lives [ZM08]. localement [Lus40a]. Location [Hin73a, Hin73b, Hoa74, Hub64, Law75, MT49a, Pre75, Rob72, Tak73, ABH^{+}72, HT49, MT91b, PT81, TM63, Tuk78g]. Locations [J.77b, RT73]. logarithmic [BK46, Ric46, Wis47a]. logical [BGvN46]. logits [Ber46]. long [RT72]. long-tailed [RT72]. Look [Efr79, Tuk82a]. Looking [Bar81, Tuk80c]. looks [Tuk66a]. Lotteries [JDC20]. Low [HMTW47, Tuk78g]. low-effort [Tuk78g]. lower [Jon48]. Luminaries [JDC20]. lune [Cla54]. lung [KFGT96]. Lyle [Zie88].

M [Hea84]. M. [Kra39, Sim46, Tuk87d]. Machine [CT65, CT75, JK92]. machines [Har46]. Madison [Lan59, Tuk84j]. magnetic [ST49]. Mahalanobis [Tuk74d]. main [Cla54]. Maintained [Ake02]. Major [Tuk93c, HT92]. male [CMT54a, McN55]. males [KFGT96]. Mallows [Zie91, Bri06, Buj06, Efr06, Hub06, Lan06]. man [Tuk86f, TB66]. Management [BJT78]. manifolds [FTT80, TFT79]. manpower [T^{+}62a]. Mantel [MBB^{+}75]. Many [ALTY00, CT89a, Tuk54b, Tuk91e]. Mapping [ACMT76, Tuk79g]. Maps [Ber83, Ber10]. March [Bar81, Mey56]. Markov [Tuk75a]. Master [Tuk97c]. Mathematical [BL01, Fis50, Pét38, Sha48, Ste42, Tuk55a, Zie91, Egu46, Mal90]. Mathematicians [Tuk67e]. Mathematics [LT46, Tuk55a, Tuk58e, Tuk65g, T^{+}67, Tuk75b, T^{+}62a, Tuk63i, Tuk71a, Tuk72a, Tuk63e, Tuk84g, Bla47]. Mather [Tuk47j]. maxima [Tuk49e]. Maximalkorrelation [FM48, Ric49]. Maximum [Tuk49a, BT83, Bri64, CT02]. May [Cra87, CT89a]. Mean [MPT91, Nai48a, Tuk50f, Che43, CT56, Lev47, Mas04, Nai48b, Tuk49y]. Mean-Square-Successive-Difference [Tuk50f]. meaning [BT74a]. Means [PT65, Ram56, BT46, Fes46, Fin46, Tuk48d, Tuk48e, Tuk49c, Tuk69b, Tuk84d]. measurable [Cot40]. Measure [Tuk50f, HT49]. measured [Tuk77h]. Measurement [BT58a, BT58b, LT64, BT59, CT89b, Tuk59e, VVT16]. Measurements [Lid01, CHT89]. Measures [GKT98]. Measuring [TH49]. mechanically [TT66]. Medal [Tuk84j]. median [CT02]. medians [Tuk46]. medie [Lov42]. Mediziner [TW52]. Meeting [T^{+}62a, FRT92, Tuk47a]. Mellin [Eps48]. Memoir [Bri09, Mos05, Cla54]. Mémoire [Cla54]. Memorandum [Tuk49k, Tuk49l]. memorial [Ano01b]. Memoriam [Kaf01]. Memories [Ake02, Bri02c]. Memory [Har03, Tuk74d]. mence [Tuk49e]. Mercy [Tuk47l]. Messrs. [JP61, KFT^{+}60, YKT^{+}59]. mesurables [Ars40, Die46, Kel40a, Kel40b, Kel41, Kos40, Max39, Otc39]. Method [GWD17, HP94, PT97, TP98a, TP98b, TP99a, TP99b, Ber45, Bri64, CBT98, CT98, Egu46, FMT04, Fré46, Fré48, Gau66, PKCT95, Thu46, Tuk67b]. Methoden [Stu37, TW52]. methodo [Gau66]. Methodological [FSG^{+}78, Tuk80d, Tuk78f]. Methodology [Tuk75c, Tuk76d, Tuk79c, Tuk59h, Tuk61f]. Methods [Efr79, Hea84, Mey56, Tuk37, Tuk47c, TW52, Tuk52c, Tuk85j, AMT46, BT66, Fer09, Fis55, JV85, MT49b, PT56a, PT56b, Stu37, Tuk51c, Tuk51d, Tuk53c, Tuk84f, Tuk84h, TV85, Tuk86e, Tuk86i, Tuk49r]. Metodo [Lov45, Cas43, Rom46, Rom47]. metoprolol [TL87]. metric [AT59, Tuk39a]. métrisation [Ale40]. microchip [CBB^{+}99]. Microwave [Tuk61e]. might [Tuk98]. Million [Tuk55d]. Mind [Thu47]. Mineral [Tuk84a, Tuk84b]. Mineral-Resource [Tuk84a]. mineralogic [KT56]. minima [Tuk49e]. minimum [Tuk58b]. Mining [Ake02]. Missile [MBT^{+}45]. mixed [Tuk59f]. Mixture [Tuk86j]. model [BT78, EHTW88a]. model-based [EHTW88a]. modeled [Tuk83]. Modeling [Par79, Tuk78a, Tuk86j, Tuk78c, Tuk79a]. models [Tuk77a]. moderately [Tuk49y]. Modern [BGT67, Tuk77f, Cox16, Tuk85h]. modification [Tuk85g, Tuk88f]. modified [Bar46]. molecular [Tuk73a, Tuk74a]. Moment [Hod58, Rid39, Tuk56a]. Moment-Like [Hod58, Tuk56a]. Moments [HK47, Tuk49m, Tuk49n, Egu46, HMTW47, Jon48, TM47, Tuk48b, Tuk53d, Tuk57j]. monotones [Kur40b, Kur41]. Monte [ABTT56, Mey56, TT56, Tuk85j, TV85, TT90]. Moon [Cla54]. Morgenthaler [Ros83, Won00, AB99]. Mortality [FSG^{+}78, KFGT96, Tuk76b, Tuk78f]. Mosteller [Ano07, Bea86, Col84, Dav88a, Dav88b, F.71, Hea83, Ile93a, Ile93b, Kan93, Kem87, Kem01, McN55, Osb78a, Osb78c, Sch93, Stu84, Tho86, Zie01, Ans88, CHNT92a, CHNT93a, Hol79, Tuk48i, Tuk90b]. move [TM75]. move-then-take [TM75]. MR2248365 [Bri06, Buj06, Efr06, Hub06, Lan06]. muestreo [CMT56]. Multi [FFT75, GWD17]. Multi-dimensional [FFT75]. Multi-resolution [GWD17]. Multicenter [CHNT93a, CHNT92a]. Multicentre [TL87]. multidimensional [FFT74a, FFT74b, FFT88]. multihalver [FMT04]. multipla [Lov45]. Multiple [BB02, BT83, Hay84, HP94, KLTW65b, KLTW65c, Thu47, Tuk47g, Tuk47b, Tuk84m, Wai93, You95, Zie95, LT01, Lia39a, Som93, Tuk53b, Tuk84i, Tuk84l, Tuk91d, Tuk92d, Tuk93c, Tuk93f, Tuk95a, WJT94, WJT99, Bra94]. multiple-adjustment [Tuk93c]. Multiple-Factor [Thu47]. multiple-response [Tuk93c]. multiplication [HR10]. multiplicity [JLT01, Tuk77g]. Multipolishing [MT01a]. Multiresponse [BT99, Coo00, Hec01, Tal00]. Multivariate [Bar81, Bar47, JK73, KT56, Fin46, Mas04, Tuk48f, Tuk90n]. Munich [Wat10]. munis [Die39b]. myocardial [To78, TL87].

N [Tuk49r, Tuk86j]. N. [EHTW88a]. Nablyudeniyi [Tuk81a]. Nachlass [Gau66]. NAEP [Tuk90f]. Näherungsverfahren [Kno43]. Name [Cip00, Hua71]. Named [Tuk74d]. NAPAP [T^{+}91a]. narrowly [Tuk83]. Nation [T^{+}70b]. National [Mac91, Tuk63g, T^{+}70a, Tuk76a, T^{+}91a, Tuk76c, Tuk90f, T^{+}66, WJT94]. National-Health [Tuk76a]. Natural [BCH^{+}86, TW52, Tuk86c]. Nature [Lag59, Ars40]. Naturwissenschafter [TW52]. NBS [Lid01]. NBS/NIST [Lid01]. near [FTT80, TFT79]. nearly [Tuk92b]. Need [Tuk59a, Tuk80e]. needs [T^{+}62a]. Networks [Ber83, Ber10, LSZ07]. neues [Wis47b]. New-York-City [FSG^{+}78]. next [MT90, Tuk80b, Tuk92d, Tuk93f]. Nile [Tuk90e]. ninther [Tuk78g]. NIST [Lid01]. NkNS [Tuk90i]. No [T^{+}62a, Won00, Bri06, Buj06, Efr06, Hub06, Lan06, Sim87]. noise [BT84, BT85, TH49, TB82]. Nombres [Eyr47c]. nomination [FMT04]. Non [Mil97, ST45, Tuk47e, Tuk57c, AT59, AT63a, Cot40, Die39c, FRT92, Tuk49p]. non-additivity [FRT92, Tuk49p]. Non-Central [Tuk57c]. Non-Constructive [Mil97]. non-measurable [Cot40]. non-metric [AT59]. non-numerical [AT63a]. Non-parametric [ST45, Tuk47e]. nonignorable [Tuk86j]. Nonlinear [Tuk84h, Wie56]. Nonparametric [Par79, ST60, ST61, Tuk48f, Tuk79a]. nonresponse [Tuk86j]. nonsuperposable [Tuk84h]. nonuniqueness [Tuk49o]. Norbert [Tuk52b]. Normal [Tuk55c, Tuk55d, ABTT56, Die39c, Fin46, Jon48, Lev47, TT56, Tuk49i, Tuk49z, Wol46]. normalization [GL46]. normalizations [TW49]. North [BCEP94]. not-quite [Tuk58a]. Notch [Tuk88e]. Note [Aro47, Rud66, Sim46, TW49, BT38, BT40b, Fré48, Tuk48g, Tuk90j]. Noted [Tuk54c]. Notes [Cue47a, Pon38, Que56, Rud66, Tuk62b, Tuk73f, CLW67b, Pen47, Tuk42]. nova [Gau66]. NSF [Tuk63g]. NSF-GN-297 [Tuk63g]. Number [Cue47b]. Numeric [Cra87]. numerica [Cas43, Rom46, Rom47]. Numerical [CMAT76, HK47, AT63a, Gol77, Lan59, RK24, Tuk66c, Tuk67c]. Numerisches [RK24]. nuovo [Cas43]. nurses [Tuk47l].

o [Cas43]. Obituary [Tuk79d]. observation [Tuk73h]. Observations [Tuk81a, HMT50, Per47, Tuk47f]. obtain [Ric46]. offer [Tuk63h]. Office [Tuk47k, Tuk63g]. Oh [Wei88]. One [Tuk49p, Tuk90k]. operating [Tuk49b]. opération [Lia39a]. Operations [BJM^{+}20, T^{+}90a, Lia39b]. opportunities [Tuk62a, Tuk80d]. optimistic [MT82c, TM82]. Oral [Asp84]. orbit [Cla54]. orbite [Cla54]. Order [MT75, AT63a, HMTW47, Jon48, ST45, Tol47, Tuk90a]. ordering [Cue47b, TP98b, TP99b]. orderly [BT83, HT89, Tuk58b]. ordinary [Tol47]. ordinaux [Eyr44]. ordonnés [Kra39, Kur40a, Kur40b, Kur41]. Origin [Buc81, CT93, TB84]. other [Ric46, Tuk63f, Tuk63g, Tuk71d, Tuk72b]. our [TAB^{+}65]. outcomes [Tuk86e]. outlier [FMT04, HIT81, HIT86, Sim87, THI81]. Outliers [BT95]. outline [MT91a]. overlapping [TP98a, TP99a]. Oversight [T^{+}91b, T^{+}91a]. overview [MT82a, TW66, TM80]. Ozone [T^{+}77].

P [Cha92, F.71, Hea84, Hin73a, Hin73b, Hoa74, Law75, Pre75, Rob72, Tak73]. P. [TV85]. p368 [AT73a]. pages [Won00]. Panel [HBK^{+}78, TAB^{+}65]. Paper [Ans65, CT93, Jac65, Mal06, MT49d, Bri06, Buj06, Efr06, Hub06, Lan06, Tuk67b, Tuk78b, Tuk78f, Tuk82c, Tuk84l, Tuk85d, Tuk86e, Tuk86g, Tuk86h, Tuk86i, Tuk86j, Tuk87d, Tuk92a, Tuk95b]. Papers [KFT^{+}60, Cra87, Sle74, Tuk47c, YKT^{+}59, Rud66]. parabola [Tuk49e]. Paradox [KMT99, Pon38]. parallel [GK93]. Parameter [Hub64]. Parameters [CMAT76, Rao45, Tuk60b, Tuk61a]. parametric [ST45, Tuk47e]. paramétrique [Kon40]. parasite [CHT00, TCH00]. Parsimony [Tuk90k]. Part [BCG^{+}93, HT92, Tuk38, Tuk65j, Tuk73b, BT58a, BT58b, KLTW65b, KLTW65c, MT49a, MT49b, MT50a, MT50b, Tuk49k, Tuk49l, Tuk51c]. partial [Tuk58c]. partially [Tuk95b]. particular [NS56]. Particularly [Tuk79h, TB82]. partiellement [Kur40b, Kur41]. Partition [Hoa61a]. partitions [BT83, Tuk57h]. parts [Tuk49f, Tuk77d, Tuk90d]. Parzen [Tuk79a]. passion [Haf00]. path [Tuk54a]. pattern [Tuk82f]. pattern-recognition [Tuk82f]. patterns [Tuk49d]. Pearson [Bañ46]. pebble [CBB^{+}99]. Percentage [PT65, Tho47, Ban49, GL46]. percentages [Tuk60c, Tuk90l]. percentiles [Tuk49t]. Performance [HIT86, Sim87, HIT81, Tuk81c, THI81, Tuk86g]. performance-based [Tuk86g]. performing [CBT98, CT98]. period [Stu37]. Periodenforschung [Stu37]. periodic [Run03]. Periodischer [Run03]. periodogram [Stu39]. Periodogrammrechnung [Stu39]. Permuted [J.77c, TT66, RT75]. personal [Cox16, Tuk82i]. Perspectives [WJO95]. perturbations [Cla54]. pg [JT01]. phase [Tuk74b]. phi [Jur47]. Philosopher [Dem02]. Philosophy [Zie88, Tuk82b, Tuk88b, Tuk88c, Tuk91d, Jon86a, Jon86b]. Phrase [PCT91, PHC^{+}94]. physical [T^{+}62a]. Physicians [TW52]. physics [Tho63, T^{+}67]. picturing [Tuk75b]. Pieces [BBB59, Tuk59f]. Pioneers [Wei88, Cox16]. pitfall [Vas47]. place [HR10]. plane [Thu46]. planètes [Cla54]. planets [Cla54]. Planning [Buc62]. plans [Kos40, Ste45a, Tuk49b]. Plant [BT99, Coo00, Hec01, Tal00, BT97]. Playfair [Wai90]. Plots [AT73a, Ley77, AT73b, CT89b, MTL78, MT01a, MT01b, TML76, Tuk88e]. plotted [Tuk79g]. plotting [TA71, Tuk79e]. Point [BT58a, BT58b, MT91b, BT59, Tuk49h, Tuk49z]. Points [PT65, Tho47, Tuk57c, Ban49, GL46, Tuk49w, Tuk78e, Tuk90r]. Poisson [Mac48]. polarization [ST49, ST51]. Policy [Tuk76a, Tuk76c]. polluted [TB66]. Pollution [FSG^{+}78, TAB^{+}65, Tuk66b, Tuk76b, Tuk78f, T^{+}70b]. polyconfidence [Tuk81b]. polykays [Tuk51b]. polynomial [HR10]. polynomials [BT74a]. Polysampling [MT91d, Ros83, Tuk87c, PT81, Tuk81b, Tuk81c]. polytope [Tuk39a]. Population [EKT89, Tuk85f, Fin46, Rid39, ST44, Tuk85b, Wol46]. populations [Bos47, Tuk49i]. portions [Lev47]. Possible [FSG^{+}78, Tuk78f]. possibly [CHT89]. potentials [Tuk37]. pour [Lia39a]. Power [BT58a, BT58b, PT56a, PT56b, BT74b, BT74a, BGT67, BT59, Max40a, Tuk50d, Tuk57f, Tuk59d]. pp [CMT55, Zie01]. PPD [TDS50]. PPD-S [TDS50]. Practical [Goo58, MT91d, MT49a, MT49b, MT50a, MT50b, Ros83, Tuk68b, Dan40, DL42a, DL42b, Tuk60c, Tuk90l]. Practice [AB99, Won00, BFM97, Tuk48j, Tuk48k]. practices [Tuk55e]. practicing [Tuk66a]. Prasanta [Tuk74d]. pratique [Ver47b]. prechosen [TT81c]. Precipitation [T^{+}91a]. precision [MT82c, MT84b, TM82]. predict [Tuk80b]. predicting [ST91]. prediction [Wie56]. Preliminary [BGvN46]. Preparation [TT81c]. prepared [TT66]. presence [BT84, BT85, TB82]. present [Tuk90m]. presented [Cra87]. President [TAB^{+}65, RWM^{+}72, T^{+}62a, T^{+}70b]. Presidential [MT78]. Press [Won00]. pretty [Tuk86h]. prevention [To78]. Price [Won00]. PRIM [FFT74a, FFT74b, FFT75, FFT88]. PRIM-9 [FFT74a, FFT74b, FFT75, FFT88]. Princeton [Bla47, Cra87, LT46, Won00, GT73, Tuk74b, Tuk78h]. principales [Cla54]. Principles [CMT54b, Zie88, Jon86a, Jon86b, Tuk90o, Tuk93b]. prior [Jef46]. probabilità [Odo42]. probabilities [Bar46]. Probability [DT73a, J.77a, J.77c, J.77b, MT49d, RT73, RT75, AT73b, HR08, Jef46, Per47, Ric46, Tuk63b, Tuk73c, Tuk90m, Tuk49r]. Probit [FS48, Bar46]. Probits [FS48, Ber46]. Problem [HK47, LLT54, Tak73, Tuk90e, ABTT56, Tuk48c, Tuk48h, Tuk49q, Tuk53b, Tuk58b, Tuk62a, Tuk82f, Tuk84i, Tuk86e]. Problème [Eyr47c, Ale40, Eyr43, Eyr45]. Problèmes [Lag65]. Problems [Bla47, DGP40, LT46, LSZ07, McN55, TP40, T^{+}48, Tuk54d, CMT53, CMT54a, GT60, Jef46, Lag65, Mol46, PT56a, PT56b, Tho63, Tuk48j, Tuk48k, Tuk77g]. Procedure [CHNT93a, CHNT93b, Hay84, MACT76, RK57, ST60, ST61, BT83, CHNT92a, CHNT92b, MBB^{+}75, Som93]. Procedures [HT89, LT01, ST82, TM63, Tuk81b, Tuk84o, Tuk90o, T^{+}91b]. Proceedings [Bar81, Tuk68b, Wat10, Cra87, WJO95, BCEP94, Lan59, Ros63]. process [Mas04, Sam45]. Processes [Tuk65e, Tuk65d]. Processing [BBB59, Tuk61e]. prodédés [Die46]. prodotti [Odo42]. produced [Cla54]. product [TW46a]. produites [Cla54]. professional [Bri02a]. Professor [Haf00, Hoa02]. profiles [BT97]. Profit [GS66]. Program [T^{+}78, Tuk86g, T^{+}91a]. progress [GMT76, Tuk62a, TBS77, T^{+}91b, WJT94]. Project [Buc62]. Projection [Tuk85e, Tuk85a, FT73, FT74, Tuk87d]. projections [Ars40, Nov39]. Proof [Hay84]. Propagation [Lag59, Tuk57e, Tuk58c]. Properties [JR71, DGC11, Tuk53d]. proportion [Tuk95a]. propos [Ver47b]. Proposed [T^{+}90b]. propriété [Kur40a, Lia39b]. prospects [Tuk84c]. Protection [T^{+}77]. providing [Tuk93a]. pseudo [BT40a, BHT63]. pseudo-autocovariance [BHT63]. pseudo-spherical [BT40a]. Publication [CT93]. Publications [Ano02, Lid01, Tuk47b]. puissance [Obr47]. pulse [Tuk59a, Tuk59b, Tuk59c]. purpose [Tuk68c]. Purposes [MT91e]. Pursuit [Tuk85e, Tuk85a, FT73, FT74, Tuk87d]. Pushback [CHNT93b, CHNT92b].

Qualitative [CHNT93b, HT91, CHNT92b]. quality [TAB^{+}65]. quantiles [MT00]. quantitative [AT63a, GMT76, HT91, Tuk59h, Tuk61f, Tuk86f]. quantities [Tuk59d, Tuk92b]. quefrency [BHT63]. query [PHC^{+}94, MT54a, MT54b, Tuk55b]. Question [Tuk47f, Tuk49a]. Questions [FSG^{+}78, MT49c, NGL^{+}94, PTM47, Pon38, Tuk78f]. qui [Ver47a]. Quick [Tuk51c, TA71, Tuk53c, Tuk59g]. Quicksort [Hoa61b]. Quiet [Ans03]. quite [Tuk58a].

R [AB99, Boa85, Hin73a, Hin73b, Hoa74, Kem01, Law75, Pre75, Pri85, Rob72, Tak73, Tuk52c, Tuk86e, Tuk90f, Won00]. R. [Tuk86j]. R23 [AT73a, Ley77]. Raiffa [Kem01]. Raise [JDC20]. Raleigh [BCEP94]. Rand [Tuk55d]. Random [Tuk55d, Dwi47, Tuk49m, Tuk49n, Tuk57h]. randomization [Tuk85g, Tuk85h]. Randomized [BT95, Tuk85g]. Range [BCT56, JR71, Tuk55c, TT90, Tuk49z, Tuk50a]. Ranges [KLTW66]. rank [BT78, Tuk49u, Tuk85g]. rankings [Tuk49q]. Ranks [ST60, ST61, Tuk57h]. rapports [Kra39]. rate [LT01]. Rates [BCH^{+}86, KT93, Tuk86c, Tuk88e, Tuk84k]. ratings [Tuk86g]. Räume [Fom40, Kat40, Ste39, ŠB41]. Ray [PST48, Dan40, LD39, DL42a, DL42b]. Rayleigh [Tuk59a, Tuk59b, Tuk59c]. Re [Tuk79e, Coo87b, Tuk82h, Tuk90h, Tuk92b]. re-discovery [Coo87b]. Re-expression [Tuk79e, Tuk82h, Tuk90h]. re-expressions [Tuk92b]. reading [Tuk90f]. realistic [MT84b]. reality [Pea55]. reassignment [Tuk85h]. Recapture [KMT99]. receipt [Tuk84j]. Recherches [Lag59]. Rechnung [RK24]. recognition [CBT98, CT98, Tuk82f]. Recommendations [TDD^{+}72]. recursion [Bos47]. reference [Fes46, NS56]. refined [HMT83b, HMT83a]. reflection [Cox16]. Reflections [Tuk92c]. reformulation [Tuk77a]. regions [Tuk47e, Tuk48f]. Regression [CDT93, MT77, Tuk82c, Tuk85j, Fin46, GT69c, JV85, Tuk51a, Tuk54a, Tuk57b, Tuk69b, TV85, Tuk91c, Vil47, Osb78c, Hol79, Osb78a]. Regressiya [MT82b]. réguliers [Ale40]. Rejection [BCT56]. Rejoinder [Tuk93d, Tuk93b]. Related [Tuk55c, Tuk85j, FT50, JV85, KFGT96, Tuk49v, TF49, Tuk59d, TV85, Tuk87f]. relation [Pea55]. relations [Tuk49f, Tuk67b, Tuk90d]. relationship [Tuk60c, Tuk83, Tuk90l]. Relative [ST60, ST61]. relatively [HT49]. relativity [McV56]. release [TS76]. Relevance [Tuk79c, Tuk57g, Tuk75c, Tuk76d]. relevant [Tuk95b]. Reliability [GT58, Gut45]. Remark [Ley77, Ley77]. Remarks [Tuk84j, Tuk74e]. remembered [Haf00]. Remembering [Fer03]. Reminder [Tuk84k, Tuk84l, Tuk84m]. Reminiscences [Leh08]. Repaired [JDC20]. répartitions [Kre46]. repeated [TDS50]. Reply [HIT87, KLTW65a, Sim87]. Report [KHS^{+}55, McN55, RWM^{+}72, T^{+}62a, TAB^{+}65, Tuk77e, Tuk84q, T^{+}91a, Tuk97c, CMT53, CMT54a, CMT55, Tuk62a, Tuk63f, Tuk63g]. representation [Run03, Run05, Tuk53d, Eyr40a, Kon40]. Representations [HP94, Eyr40b]. Representative [dTP46]. representing [BPT81]. requires [GMT76]. requiring [TP98b]. Research [Ake02, Tuk47k, Tuk52c, Tuk59a, Stu37, Tuk53a, Tuk62a, Tuk62d, Tuk62e, WJO95, Lag59]. Residuals [AT63b, BT95, De 64, Olm63]. Resistant [Tuk85j, HIT81, HIT86, JV85, ST82, Sim87, Tuk76e, Tuk76f, Tuk76g, Tuk76h, TBS77, Tuk78g, THI81, TV85]. resistant/robust [Tuk76f, Tuk76g, Tuk76h]. resolution [GWD17]. Resource [Tuk84a, Tuk84b]. Resources [BJT78]. respect [Cla54, TDD^{+}72]. Response [T^{+}77, TCH85, Tuk93c]. Responsibility [Tuk79c, Tuk75c, Tuk76d]. Restoring [TAB^{+}65]. Results [FSG^{+}78, Tuk81a, Tuk85j, CT89b, Mal79, MT82c, MT84b, T^{+}70a, Tuk78f, TM82, TV85, KLTW65b]. retest [Gut45]. retrieval [PHC^{+}94]. Review [AB99, Ano07, Bea86, Bla47, Boa85, Cha92, Col84, Dav88a, Dav88b, F.71, Hea83, Hea84, Hec01, Hin73a, Hin73b, Hoa74, Hod58, Ile93a, Ile93b, J.77a, J.77c, J.77b, Kan93, Kem87, Kem01, Law75, McN55, Mil74, Osb78b, Osb78a, Osb78c, Osb78d, PT50, Pre75, Pri85, Qua74, Rob72, Ros83, Sch93, SG80, Ste41, Stu84, Tak73, Tal00, Tho86, TW46c, TW46d, Tuk47i, Tuk47h, Tuk47g, Tuk47j, Tuk47k, Tuk49s, Tuk49r, Tuk52b, TW52, Tuk52c, Tuk55d, T^{+}91a, Won00, You95, Zie88, Zie89, Zie91, Zie95, Zie01, Coo00, MT49a, MT49b, MT50a, MT50b, T^{+}91b]. reviewing [CT89b]. Reviews [Ehr79, Hol79]. revisited [MHD87]. Revolutionary [Wai03]. Revolutionized [Sal01]. Rezul’tatov [Tuk81a]. Richard [Tuk93b]. Riemann [Kan39]. Rietz [Efr79]. risk [VVT16]. risoluzione [Cas43, Rom46, Rom47]. Rivers [T^{+}78]. Robb [Tuk86e]. Robin [Tuk87d]. Robust [ABH^{+}72, CDT93, Col84, Hea83, Hin73b, HMT83c, HMT00, Hub64, Hub02, Kem01, Rob72, Stu84, Tak73, Tuk76e, Tuk79f, Zie01, PT81, ST01, Tuk76f, Tuk76g, Tuk76h, TBS77, Tuk78g, Hin73a, Hoa74, Law75, Pre75]. Robust/resistant [Tuk76e]. Robustness [GWD17, Kaf03, MT91d, Tuk74b, FMT06, GT73, MT83, Tuk79h, Ros83]. rock [KT56]. Rogers [Hin73a, Hin73b, Hoa74, Law75, Pre75, Rob72, Tak73]. Role [BJT78, Tuk76c, Tuk76a, Tuk82d]. root [FT50, Tuk48g, TF49]. Roots [Ham46]. Ross [J.77c, J.77b]. rotation [Thu46]. Route [MT91d, Ros83]. routes [Tuk57a]. routines [TA71]. row [Tuk49j]. row-by-column [Tuk49j]. Rubin [Tuk86j]. rule [HIT81, McG11, Per47, THI81]. rules [HIT86, Sim87]. ruly [DRT69]. Rural [GKT98]. Russian [McG11, Tuk81a].

S [AB99, Hea84, Ros83, STM^{+}65, Tuk93b, Won00, Zie89, TDS50]. safety [MBB^{+}75]. sale [SC00]. salvo [Tuk86n]. same [PHC^{+}94]. Sample [Nai48a, BT46, Egu46, HT49, HIT81, MT50c, ST44, Tuk48j, Tuk48k, Tuk48i, Tuk59g, TM63, Tuk65j, THI81]. Samples [JR71, ST60, ST61, ABTT56, Fin46, HMTW47, Jon48, PT81, Rid39, TT56, Tuk49y, Tuk54b, Tuk58a, Tuk78g]. Sampling [CMT54b, F.71, Hod58, Tuk50b, Tuk50e, TT90, Dwi47, MT84a, Mol46, MT83, Tuk49b, Tuk50d, Tuk51b, Tuk56a, Tuk57f, Tuk59i, Tuk60d]. Samuel [STM^{+}65, Tuk64a]. Sanctification [Tuk69a]. sandwich [ST42]. saphe [BHT63]. SAT [Tuk86h]. saturation [Eyr45]. scalability [GK93]. scale [HT49, MT49a, MT49b, Tuk49v]. Scaling [Tuk49t]. scanned [BTW95]. Scatter [PKCT95, CPKT92]. Scatter-gather [PKCT95]. Scatter/Gather [CPKT92]. scattering [Dan40, DL42a, DL42b]. Scheffé [Tuk84l]. Schémas [Eyr40b]. Scheme [GWD17]. Schimmel [Tuk78f]. Schlaifer [Kem01]. School [Tuk74e]. Science [Don17, Sal01, Tuk63g, TAB^{+}65, T^{+}70a, T^{+}62a, Tuk63c, Tuk68a, Tuk80c, Tuk86f]. Sciences [SB01, T^{+}62a, T^{+}62b]. Scientific [Cra87, Nas90, Tuk47k, BMTW49, Fis55]. Scientists [TW52]. Scott [Tuk70c, Tuk70d, Tuk84e, Tuk84n]. Search [PCT91]. searching [GMT76]. Seasonality [Tuk78b]. Second [MT77, Tuk47i, Tuk70d, Tuk84n, BPT81]. seconde [Eyr44]. segments [MT01b]. Seismology [Tuk59a]. Selected [Lid01, Tuk53c]. selection [Tuk48l, Tuk49x, TT81a, Tuk81b, Tuk86e, Tuk86g, Tuk86j]. Selective [JDC20]. Self [Tuk86g]. Self-selection [Tuk86g]. semi [Kra39]. semi-ordonnés [Kra39]. semigraphic [Tuk70f, Tuk72c]. Semiology [Ber83, Ber10]. sense [Tuk59a, Tuk59b, Tuk59c]. sensible [JT00, JT01, TJ00]. sensitivity [AMT46]. sensitization [TDS50]. séparabilité [Max40b, Lia39a]. separation [Tuk42]. separations [HT89]. sequences [TT81c]. Sequential [Tuk47k, PT50]. sequentially [Tuk90o]. serial [TT66]. serials [DRT69]. serie [Odo42]. Series [Boa85, CT65, CLW67a, FSG^{+}78, Pri85, Ros63, Rud66, Tuk50f, Tuk52b, BT74b, BT74a, BHT63, Bri84a, Bri84b, BCG^{+}93, Bri02b, CT75, Que49, Tol47, Tuk53d, Tuk63e, Tuk78f, Tuk80b, Tuk84g, Tuk84q]. Service [JDC20, Tuk63g]. ses [Kra39]. Session [FSG^{+}78, SGT^{+}78, Tuk67a, Tuk77b, Tuk86k]. Session-II [FSG^{+}78]. Set [TM46]. Sets [MT75, BT40a, Cot40, Mas41a, Nov39, Tuk42]. setting [Bri64]. Several [BC93, Tuk77d, Tuk90k, Tuk93b]. Sexual [McN55, CMT54a]. Shannon [Cro01]. shape [HT85]. Shapes [HMT85, HMT06, Dav88b, Kem87, Ano07, Bea86, Dav88a, Tho86]. shaping [Tuk59a, Tuk59b, Tuk59c]. sharpening [TT81a]. sheet [Tuk90p]. sheets [Tuk84f, Tuk84k, Tuk84l, Tuk84m, Tuk84o, Tuk90q]. Sheffield [Bar81]. shifted [TT66]. Short [KLTW65b, KLTW65c, Rud66]. Short-Cut [KLTW65b, KLTW65c]. Should [GT69a, Tuk79g, Tuk80b, Tuk84r, Tuk92d, Tuk93f]. shunning [Tuk86f]. Sibson [Tuk87d]. Sierpiński [Cue47a]. signed [Tuk49u]. signed-rank [Tuk49u]. Significance [MT50c, Tuk48i, ZM08, Fes46, JT00, JT01, Mas41b, TM63, TJ00, Vil47]. significances [Tuk77d]. Silent [Car62, MHD87]. Simple [MT84a, Tuk92b, AT63a, KT93, Tuk56a, Hod58, Tuk51c]. simplest [Tuk49u]. Simplified [Tuk50e, Che43, Tuk51b]. simulation [Tuk79h]. Simultaneous [LT64, Ber45]. sine [Run03, Run05]. Singer [Tuk86g]. Single [KLTW65b, KLTW65c, PCT91, Thu46, TW46a, Tuk57i, Tuk57j, TM63]. singularities [Tor39]. Sinuswellen [Run03, Run05]. sistemi [Rom46, Rom47]. situations [BPT81]. situs [Bal39]. Six [AT73a, Ley77]. Six-Line [AT73a, Ley77]. size [Tuk49m, Tuk49n]. sizes [ST44]. Skeleton [Tuk49v]. skewed [CHT89]. skin [TDS50]. slippage [MT50c, Tuk48i]. Slopes [Tuk85j, TV85]. Small [HIT81, THI81, Bar46, HMTW47, Jon48, PT81, Tuk48j, Tuk48k, Tuk54b, Tuk86m]. Small-sample [HIT81, THI81]. smear [GGT69]. smear-and-sweep [GGT69]. smelting [Tuk82h]. smooth [Tuk56b, Tuk58d]. smoother [BP92, Mal79]. Smoothing [Tuk52b, Tuk76e, TT81d, Tuk84h]. Smoothly [CDT93]. Snippet [PCT91]. snow [Tuk82f]. Social [SB01, GMT76, TDD^{+}72, Tuk86h]. Software [Leo00, Tuk85i, Tuk85c]. soleil [Cla54]. Solution [HK47, Mol46, Tol47, Lag65]. Solutions [DGP40, Lag65, TP40, Tuk47d]. solve [Tho63]. solving [Ber45, Tuk86e]. Some [BT85, BJM^{+}20, BT46, Cox16, DL42a, DL42b, DGC11, Eps48, FSG^{+}78, JR71, Mal79, Mol46, Pen47, Per47, ST82, Smi47, Tuk42, Tuk48j, Tuk48k, Tuk50e, Tuk53c, Tuk57g, Tuk70e, Tuk70f, Tuk72c, Tuk73g, Tuk77g, Tuk81c, Tuk82g, TT82, Tuk90n, Tuk90o, Wai93, BT84, Cue47a, EHTW88b, Har46, HIT86, HT89, Nov39, RT72, Sim87, Ste42, TW49, Tuk78f, Tuk81b, TP99b]. Son [Lag59]. Sorting [MACT76, ACMT78, TACM77]. sound [Lag59]. Souslin [Max40c]. Souvenir [Tuk84o, Tuk90q, Tuk90p, Tuk84f]. Soviets [TB84]. space [BS39, DGC11]. spaces [Mor40, Tuk41]. spaceship [TB66]. Spatial [Tuk84a, Tuk84b]. Spatial-Analysis [Tuk84a]. Special [Ano01b, Fer09, WJT94, Tuk97c]. specifications [Tuk59g]. Spectra [BT58a, BT58b, BT59, Tuk59d, Tuk59e]. spectral [Bri02b, PT56a, PT56b, Tuk53d]. spectroscopic [BT74a]. Spectrum [BT84, BT85, HT66, TB82, BGT67, Tuk50d, Tuk57f, Tuk61b, Tuk61c, Tuk66c, Tuk67c, Tuk84p, Tuk84r]. Speech [Tuk97a]. SPES [Tuk89]. spherical [BT40a, Blu41]. splitting [JT87]. Spread [ST60, ST61]. Spring [Car62, MHD87]. Square [Tuk50f, Ber46, FT50, Nai48b, Tuk48g, TF49]. square-root [Tuk48g, TF49]. squared [Rea93]. squares [CT56, MPT91, Tuk67b, Tuk75a, Tuk90j]. Staircase [AMT46]. Standard [PT65, Tuk49w, Tuk51d, Tuk90r, ZM08, Ric46, Tuk50a, TDS50, Tuk51c, Tuk74b, Tuk92a]. standardization [Tuk69b]. Standards [Lid01, ST91, Tuk86d]. Stanley [Tuk64a]. stash [Tuk82b, Tuk88b, Tuk88c]. State [ALTY00, Tuk90m, WJT99]. state-to-state [WJT99]. stated [MT82c, MT84b, TM82]. Statement [Tuk66b, Tuk74f]. station [FRT92]. Stationary [Tuk52b, Tuk53d]. Statist [Sim87]. Statistic [RK57, BLA18, Bos47]. Statistical [AT73a, CMT53, CMT54a, Fis55, GKT98, Kem01, Ley77, McN55, NGL^{+}94, Par79, Pea55, She07, Tuk47c, Tuk47j, Tuk47k, Tuk49r, TW52, Tuk52c, Tuk59h, Tuk61f, Tuk77h, Tuk79g, Tuk82d, Tuk85i, TCH85, Wai03, ZM08, BK46, Cox16, Har46, Rao45, Ric46, Tuk38, Tuk53c, Tuk65a, Tuk71e, Tuk73e, Tuk77f, Tuk79a, Tuk80a, Tuk85c, T^{+}90b, Tuk47k]. Statistically [Tuk47e, Tuk48f]. statistiche [Lov42]. Statistician [JDC20, Leh08, Leo00, Lor18, Tuk79c, Tuk66a, Tuk67d, Tuk75c, Tuk76d, Tuk86a]. statisticians [FRT92, Tuk71b, Tuk86o]. Statistics [Ake02, BT99, BCH^{+}86, BJT78, DT73a, Fis50, Hub02, J.77a, J.77c, J.77b, JDC20, KJ97, MT68, MT77, NS56, Qua74, RT73, RT75, Sal01, Tak73, Tuk50c, Tuk54d, Tuk64b, T^{+}71, Tuk73f, Tuk76a, Tuk78h, CM88, DT73b, Eps48, HMTW47, Jon48, McV56, Rea93, Rid39, ST45, Ste42, TW46b, Tuk49k, Tuk49l, Tuk51c, Tuk63b, Tuk63h, TW65, Tuk65h, Tuk65i, TW66, T^{+}67, Tuk68a, Tuk71d, Tuk72b, Tuk73c, Tuk76c, Tuk88d, Tuk90a, WJO95, RWM^{+}72, Cha92, Hec01, TW46c, Tuk86c]. Statistische [TW52]. status [Tuk63f, Tuk63g]. Steady [GMT76]. Steps [Tuk90s]. steroid [TM47]. Stieltjes [Kan39]. straight [MT01b]. strategic [Tuk93a]. Stratum [KMT99]. Strengthening [T^{+}62b]. Stretch [Buc62]. Streuungsmass [Wis47b]. Structural [Bal40, GT69c]. Structure [Lus40b, Die39c, Die39b, Kel41, Tuk73a, Tuk74a]. Student [JDC20]. Studentized [TT90, Tuk90o, Nai48a, Nai48b]. Studies [KMT99]. Study [ACMT76, CMT55, KHS^{+}55, Tuk79h, Bañ46, BLA18, GT73, HMTW47, MBT^{+}45, Tuk86g, TL87, T^{+}66, Tuk74b]. studying [MT83, Tuk82g, TT82]. Styles [Tuk80a, Tuk84p, Tuk82e]. submarines [McG11]. subrange [BT83]. subsets [Blu41]. successful [Tuk73h]. Successive [Tuk50f, CT89a]. Sufficiency [Tuk48l, Tuk49x]. suggested [Tuk93b]. suggestions [Tuk90n]. Sul [Rol43]. Sulfinpyrazone [To78]. Sum [ST60, ST61]. summaries [CBT98, Tuk85h]. Summarization [TT81d, PT97]. Summary [T^{+}66]. Sums [Tuk57h]. Sun [Cla54]. Sunset [Tuk86n]. Supercomputing [Mac91]. supérieures [Kel40a]. supersonic [Tuk47d]. superstandardization [Tuk69b]. superstandardize [GT69a]. Supplement [Tuk84q]. supplemented [TT81d]. Survey [Hin73a, Hin73b, Hoa74, Law75, Pre75, Rob72, Tak73, Tuk59i, Tuk60d, ABH^{+}72]. survivors [TL87]. susceptible [Die39c]. susceptibles [Die39b]. sweep [GGT69]. Symbolic [Wat10]. Symmetric [JR71]. symmetrical [RT72, Tuk49e, Tuk62f]. symmetrizing [Tuk92b]. Symposium [Lan59, Mey56, Ros63, Tuk50c, Wat10, Tuk47c, Tuk50b, F.71, KHS^{+}55, CMT55]. System [FFT75, MBT^{+}45, Buc62, FFT74a, FFT74b, FFT88, PHC^{+}94, Tuk63f, Tuk63g, Tuk65a, TB82]. systematic [Tuk49f, Tuk90d]. Systematization [Tuk55e]. système [Max40c]. Systems [Tuk85i, Tuk85c, Tuk86a].

T [AB99, Won00]. Table [FS48, Jur47, Lev47, KT88, Ric46, Tuk74b]. Tables [Ano07, Bea86, Dav88a, Dav88b, HMT85, Kem87, MT75, MT49a, Stu39, Tho86, TM46, Bos47, HMT06, Ric46, Tuk49v, Tuk77d]. Tafeln [Stu39]. tagging [Tuk63f, Tuk63g]. tailed [RT72]. take [TM75]. taking [Tuk85h]. Task [T^{+}70b]. Tasting [Sal01]. Taylor [Tol47]. Tea [Sal01]. Teacher [Har03, Mor03]. Teaching [Tuk58e]. Team [JDC20]. Technical [MT91b, Tuk64b, Tuk65h, Tuk65i, Rud66, Tuk84q]. technique [ACMT78, Bar46, Gar69, PHC^{+}94, TACM77, Tuk78g]. Techniques [CHT89, TW65, Tuk88d, ABTT56, BGT67, MT82a, MT50a, Tuk49d, Tuk59a, Tuk59b, Tuk59c, TM75, Tuk76f, Tuk76g, Tuk76h, Tuk77f, Tuk79f, TM80, Tuk82e, Tuk86b]. Technology [Lid01, T^{+}62a, Tuk71e, Tuk73e]. Technometrics [Bri06, Buj06, Efr06, Hub06, Lan06]. Telephone [TM75]. Teletypewriter [AT73b, TA71]. teletypewriter-like [TA71]. Teoria [Kin45, Tuk47h]. Term [TB84]. termes [Ver47a]. terminal [TA71]. terms [Tuk87f]. terre [Cla54]. Test [Ram56, Gut45, JT00, JT01, Mas41b, MT50c, Nai48b, OT47, Ste45b, Tuk48i, Tuk59g, TJ00, Vil47]. test-retest [Gut45]. Testing [ETW71, TM46, TEW72, TCH85, AMT46, MBB^{+}75]. Tests [CHNT92b, CHNT93b, FRT92, JLT01, MPT91, Que49, Tuk47l, Tuk49u, TDS50]. Text [PCT91]. their [Dan40, DL42a, DL42b, Pea55, PT56a, PT56b, Tuk80a]. thematic [CBT98]. theorem [Far46, Tuk49o, Tuk58d, BP92]. théorème [Eyr47b, Kel40a, Kre46]. theorems [ST42]. Theoretical [Pon38, Mal79]. Theoria [Gau66]. Theorie [Ste39, ŠB41, Eyr47c]. Theory [Pét38, Sha48, Tuk49r, AT63a, Cox16, Dwi47, Fer09, Gau66, Hal46, Kan39, McG11, MT49a, MT49b, MT50a, MT50b, Sle74, ST51, Ste42, Tuk50d, Tuk56b, Tuk57f, Vas47, Kem01, Tuk49s]. thickness [Tuk77h]. thinking [She07, Tuk65a]. Third [Tuk57j]. Thomson [Tuk47i]. Thoughts [Tuk88f, Tuk73g, Tuk77g, Tuk81b, Tuk81c]. three [TT81b]. Thurstone [Tuk47g, Tuk47b]. Tightening [Tuk93e]. Time [Boa85, Bri84b, Coo94b, FSG^{+}78, JDC20, Pri85, Ros63, Tuk50f, Tuk52b, BHT63, BCG^{+}93, Bri02b, CT89b, Que49, Tuk53d, Tuk63e, Tuk78f, Tuk80b, Tuk84g, Bri84a]. Time-Series [FSG^{+}78, Que49]. Times [Ans03]. Timing [HMT50]. Titles [RT75, J.77c]. Today [Tuk74a, Tuk63h, Tuk73a, Tuk73d, Tuk74c]. tolerance [ST44, Tuk47e, Tuk48f]. tolerances [Tuk57e, Tuk58c]. Tomography [LSZ07]. tomorrow [Tuk73a, Tuk74a]. tool [HT89]. Tools [Tuk64b, MT91e, Tuk65h, Tuk65i]. Top [Cip00]. Topics [Tuk67b]. topological [BS39, Coh40]. topologique [Lus40b]. topologiques [Die39b]. topologischer [Fom40]. Topology [Tuk39b, Tuk40, Ste41]. touch [Tuk60b, Tuk61a]. Tract [CMAT76, ACMT78, TACM77]. tractata [Gau66]. Training [Tuk82d, T^{+}62a]. Transactions [Tuk66a]. transfini [Max39, Max40c]. transfinies [Eyr40b]. Transfinis [Eyr47c, Eyr44]. Transform [Sin67, Eps48, HR10, BJM^{+}20, CLW67a, CLW67b, Coo87b, Gar69, HJB84, HJB85, Hua71, Sch12]. Transformation [CMAT76, Mil78, TF49, ACMT78, BK46, Tuk53d, Tuk68c, TACM77]. Transformations [FT50, Kur40b, Kur41, Fis54, Tuk57d, Tuk60c, Tuk82c, Tuk90l, Tuk90q]. Transforms [GS66]. trascendenti [Cas43]. Treatment [CHNT93b, CHNT92b]. Treatment-By-Center [CHNT93b, CHNT92b]. treatments [Tuk86e]. Trends [Ano07, Bea86, Dav88a, Dav88b, HMT85, Kem87, Tho86, HMT06, KFGT96].

Trial [BT99, CHNT93a, Hec01, CHNT92a, Tuk93e, Coo00, Tal00, TL87]. Trials [Tuk78d, Tuk77g, Tuk91e]. Tribute [Cas03]. trigonometric [LD39]. trimmed [Mas04]. trimming [DT68, TM63]. trimming/Winsorization [DT68, TM63]. triumphant [McG11]. troisième [Eyr44]. troncs [Ste47]. true [Tuk68c]. truncated [HR10, Tuk49y]. truncation [Tuk48l, Tuk49x]. Tsao [JT87]. tube [Tuk77h]. Tuberculin [Tuk47l, TDS50]. Tucker [Asp84]. Tukey [Ano01a, Ans65, Asp84, Bla47, Boa85, Bri06, BJM^{+}20, Buj06, Efr06, F.71, Hub06, Jac65, JP61, JK92, Kaf01, Lan06, McC03, Mos05, Pri85, Sim87, Ste41, You95, Zie88, Zie89, Zie91, Zie95, Zie01, Ake02, Ano01b, Ano02, Ano07, Ans88, Ans03, AT85, Bea86, BC93, BLA18, BB02, Boa85, Bra94, Bri64, Bri84a, Bri84b, BFM97, Bri02a, Bri02b, Bri02c, Bri09, BP92, Cas03, Cha92, CT02, CHNT92a, CHNT93a, Cle88, Col84, Coo00, Cox92, Dav88a, Dav88b, De 64, Dem02, DGC11, Ehr79, FM97, FM00, FM03, Fer03, Fer09, FS02, GWD17, GH93, Har03, HR08, Hay84, Hea83, Hea84, Hec01, Hin73a, Hin73b, Hoa74, HIT87, Hoa02, Hoa03, Hol79, HP94, Hub02, Ile93a, Ile93b, J.77a, J.77c, J.77b, JK73, JDC20, JV85]. Tukey [JR71, Jon86a, Jon86b, Kaf03, Kan93, Kem87, Kem01, Law75, Leo00, Lor18, Mal79, Mal90, Mal03, Mal06, Mas04, McN55, Mil97, Mil78, Mor03, Mos84, Olm63, Osb78b, Osb78a, Osb78c, Osb78d, Pre75, Qua74, Ram56, RK57, Rea93, Rob72, Ros83, Sch93, She07, Sim87, SG80, SC00, Som93, Spe02, Stu84, Tak73, Tal00, Tho86, Tho01, TGH^{+}xx, VVT16, Wai90, Wai93, Wai03, YG19, Zie88, Zie91, Zie95, AB99, Won00]. Tumorigenicity [MTCH82]. Tuning [HT92]. Twentieth [Sal01]. Twin [T^{+}78]. Two [MT75, MT01b, Mas41a, McG11, MT96, MT01a, Tuk54b, Tuk59g, Tuk74b, Tuk82b, Tuk88b, Tuk88c, Wol46]. two-handed [Tuk82b, Tuk88b, Tuk88c]. two-sample [Tuk59g]. Two-Way [MT75, MT01b, MT01a, Tuk74b]. Type [LT64, Kra39]. types [Cue47b, Tuk84k].

U.K [Bar81]. U.S. [KT93, KFGT96]. U.S.A [Won00]. Un’applicazione [Lov42]. unbalanced [Tuk57i]. unbiased [Hal46, Tuk90j]. undercount [Tuk86d]. undergraduate [T^{+}67]. Understanding [Hea83, HMT83c, HMT00, RT72, BT97, Kem01, Col84, Stu84, Zie01]. unequal [Tuk48e, Tuk84d]. uniform [Mor40]. uniforme [Die39c, Die39b]. uniformes [Die39a]. Uniformity [Ste41, Tuk40]. unit [ETW71, Lev47, TEW72]. univariate [Tuk90s]. universal [Tuk90s]. University [Bar81, Bla47, LT46, Ros63, Tuk47k, Won00, Mey56]. Unpaired [ST60, ST61]. Unsolved [Tuk54d]. Upon [BCT56]. Upper [Tuk57c, Tuk49z]. Urban [GKT98]. Urbanicity [KFGT96]. Urbanicity-related [KFGT96]. urbanization [KT93]. Usable [Tuk76f, Tuk76g, Tuk76h]. Use [CT89b, Tuk66c, Tuk84b, Tuk84a, Tuk90p, Tuk91e, Gar69, MTCH82, MT83, Tuk82h]. used [Ste42, Tuk84r]. Usefulness [MT49d]. user [Tuk79f, TP98b, TP99b]. Uses [MT49d]. Using [CHNT93a, CHNT93b, GKT98, Tho63, CHNT92a, CHNT92b, Tuk81b]. utilization [AT63a].

V [Cle88]. Vald [PT50]. Validation [ST45]. valuation [Tuk84b]. values [CT56, HT89, MT84a, dTP46, Tuk74d]. Vanishing [TB66]. Vardi [She07]. Variability [BCH^{+}86, Tuk86c]. variable [Tuk38, TW46a, Ver47b]. variables [TW46a, Tuk49b, Tuk82c]. Variance [CDT93, HMT91, Spe02, Tuk85j, Tuk87e, Tuk87a, BK46, Fis54, GT60, JT87, MPT91, MT91e, Nai48b, ST01, Ste45b, Tuk48d, Tuk49c, Tuk49d, Tuk49g, Tuk51b, Tuk54c, Tuk56c, Tuk57i, Tuk57j, Tuk58b, Tuk61b, Tuk61c, TV85, TMH91, Tuk93a, Ile93b, Ile93a, Sch93]. variance-heterogeneity [BK46]. Variance. [Kan93]. Variances [Ram56, TM46, Tuk56c, Tuk57i, Tuk57j, MT84a, Tuk51b]. Variations [MTL78, TML76]. variety [Tuk78e]. various [Fis54, GT69c, Lev47, Tuk84k]. Vectors [Thu47, ETW71, Mas41a, Tuk49g, TEW72]. Velleman [Tuk85j, TV85]. verbundene [Kno43]. Versuchszahlen [GvdW43]. versus [GKT98, Tuk86j]. Verteilungsbreite [Wis47b]. Verteilungsfunktionen [Kno43]. vertices [Tuk49e]. VI [Mal90]. Vietnam [JDC20]. View [BT58a, BT58b, BT59, Tuk48c, TT81a, Tuk82i]. views [TT81c, TT81d]. VII [Cox92]. VIII [You95, Zie95, Bra94, Tuk86k]. VIII-b [Tuk86k]. Visions [Wai90]. visual [Tuk90c]. Vital [BCH^{+}86, Tuk86c]. Vital-Rates [BCH^{+}86]. Vocal [CMAT76, ACMT78, TACM77]. Vocal-Tract [CMAT76, TACM77]. Vol [You95, Zie88, Zie89, JT01]. Volume [Boa85, KJ97, Pri85, Zie91, Zie95, Bra94, Bri84a, Bri84b, Cle88, Cox92, Jon86a, Jon86b, Mal90]. Volumes [T^{+}71]. Vorlesungen [RK24]. voting [MT78]. vs [Tuk60a]. vulnerable [TM63].

W [Ake02, AB99, Ano02, Ano07, Ans88, Ans03, AT85, Bea86, BC93, BB02, Bla47, Boa85, Bra94, Bri84a, Bri84b, BFM97, Bri02a, Bri02b, Bri02c, Cle88, Col84, Coo00, Cox92, Dav88a, Dav88b, De 64, Dem02, Ehr79, F.71, FM97, FM00, Fer03, Fer09, FS02, Hea83, Hea84, Hec01, Hin73a, Hin73b, Hoa74, Hoa02, Hoa03, Hol79, Hub02, Ile93a, Ile93b, J.77a, J.77c, J.77b, JDC20, JV85, Jon86a, Jon86b, Kan93, Kem87, Kem01, Law75, Lor18, Mal90, McN55, Mor03, Mos84, Mos05, Olm63, Osb78b, Osb78a, Osb78c, Osb78d, Pre75, Pri85, Qua74, Rob72, Ros83, Sch93, Sim87, SG80, Spe02, Ste41, Stu84, Tak73, Tal00, Tho86, Tuk49s, Tuk78b, Won00, You95, Zie88, Zie89, Zie91, Zie01]. Wainer [Tuk93b, Tuk86h]. Wald [Tuk90g]. wall [Tuk77h]. Wallace [Ans65, Jac65]. water [Tuk77h]. water-wall [Tuk77h]. waves [Run03, Run05, Tuk59a, Tuk59b, Tuk59c]. Way [MT75, MT01a, MT01b, Tuk74b]. Weather [BJT78, Tuk85g, Tuk87f]. weather-related [Tuk87f]. Weights [FS48, Tuk48a]. Weil [Kra39]. WENO [GWD17]. Where [Tuk60e, Tuk92d, Tuk93f, Tuk98, Tuk80b]. Which [Tuk65j, HT49, Tuk84r, Wol46]. White [KFGT96]. whole [Tuk73b]. wide [Gar69]. widthers [TBS77]. Wiener [Tuk52b]. Wild [TB66]. Wilder [Ano01a, Asp84, Bri09, Kaf01, McC03, Wai03]. Wiley [Zie01]. Wilks [STM^{+}65, Tuk64a]. William [Tuk93b, Wai90]. Winsor [F.71]. Winsorization [DT68, TM63]. Winsorized [DT68]. Wintner [BST38, WBST38]. within [HT89]. without [BTW95, CBT98, CT98, Fes46, TP98b]. Word [Buc81, Leo00]. words [BTW95, CT98]. Work [FS02, Bri02b, Tuk69a, Tuk80c]. Workers [Tuk52c]. Working [FS48]. Works [Boa85, Pri85, You95, Zie88, Zie89, Zie91, Zie95, Bra94, Bri84a, Bri84b, Cle88, Cox92, Cue47a, Jon86a, Jon86b, Mal90]. workshop [Tuk68b, WJO95]. would [McG11]. Writings [Ano02].

X [Dan40, DL42a, DL42b, LD39]. X-ray [DL42a, DL42b, Dan40, LD39]. xx [Zie01].

Yates [Tuk82i]. Year [Tuk85f, Tuk85b, Tuk87f]. Years [Don17, Mal06, Bri06, Buj06, Efr06, Hub06, Lan06, Tuk89]. York [FSG^{+}78, Tuk78f, Zie01]. Yule [KMT99].

Zerlegung [Run03, Run05]. zero [CT89a, CHT00, TCH00]. zigzagging [Tuk73h]. Zorn [Far46]. Zulässigkeit [GvdW43]. Zur [FM48, Ric49, ŠB41]. Zwick [Tuk90f].

References

Al-Bayatti:1999:BRB

[AB99] M. F. Al-Bayatti. Book review: The Practice of Data Analysis: Essays in Honor of John W. Tukey, by D. R. Brillinger; L. T. Fernholz; S. Morgenthaler. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2):280, ???? 1999. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

           jstor.org/stable/2681198.

Andrews:1972:REL

[ABH^{+}72] D. F. Andrews, P. J. Bickel, F. R. Hampel, P. J. Huber, W. H. Rogers, and J. W. Tukey. Robust estimates of location: Survey and advances. Princeton University Press, Princeton, NJ, USA, 1972. ISBN 0-691-08113-1 (hardcover), 0-691-08116-6 (paperback). ix + 373 pp. LCCN QA276.8 .R6 1972.

Arnold:1956:MCT

[ABTT56] Harvey J. Arnold, Bradley D. Bucher, Hale F. Trotter, and John W. Tukey. Monte Carlo techniques in a complex problem about normal samples. In Meyer [Mey56], pages 80–88. LCCN QA273 .F67.

Atal:1976:ACS

[ACMT76] B. Atal, J. J. Chang, M. V. Mathews, and J. W. Tukey. Articulatory compensation — study of ambiguities in acoustic-articulatory mapping. Journal of the Acoustical Society of America, 60(??):S77, ???? 1976. CODEN JASMAN. ISSN 0001-4966.

Atal:1978:IAA

[ACMT78] B. S. Atal, J.-J. Chang, M. V. Mathews, and John W. Tukey. Inversion of articulatory-to-acoustic transformation in the vocal tract by a computer-sorting technique. Journal of the Acoustical Society of America, 63(??):1535–1555, ???? 1978. CODEN JASMAN. ISSN 0001-4966.

Akera:2002:MJW

[Ake02] Atsushi Akera. Memories of John W. Tukey: Maintained by Bell Laboratories, Statistics and Data Mining Research Group. Iterations: an interdisciplinary journal of software history, 1 (1):1–2, September 13, 2002. ISSN 1541-843X. URL http:

//www.cbi.umn.edu/iterations/akera.html;   http://www.
cbi.umn.edu/iterations/akera.pdf.

Alexits:1940:ERP

[Ale40] Georges Alexits. Les espaces réguliers et le problème de métrisation. Commentarii Mathematici Helvetici (Zurich), 13:1– 5, 1940. CODEN COMHAX. ISSN 0010-2571 (print), 1420-8946 (electronic).

Almond:2000:DCG

[ALTY00] Russell G. Almond, Charles Lewis, John W. Tukey, and Duanli Yan. Displays for comparing a given state to many others. The American Statistician, 54(2):89–93, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/Almond.htm;
 http://www.jstor.org/stable/2686023.

Anderson:1946:SMS

[AMT46] T. W. Anderson, P. J. McCarthy, and J. W. Tukey. Staircase methods of sensitivity testing. Report, Statistical Research Group, Princeton University, Princeton, NJ, USA, 1946. vii + 135 pp.

Anonymous:2001:JWT

[Ano01a] Anonymous. John Wilder Tukey *39. Princeton Alumni Weekly, ??(??):??, March 21, 2001. ISSN 0149-9270. URL http://paw.

princeton.edu/memorials/36/31/.

Anonymous:2001:STM

[Ano01b] Anonymous, editor. Special Tukey memorial issue, volume 43(3) of Technometrics. American Society for Quality Control, Rochester, NY, USA, 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). 252–384 pp. URL http://www.allbusiness.

com/technometrics/20010801/3028944-1.html.

Anonymous:2002:PWJ

[Ano02] Anonymous. The publications and writings of John W. Tukey. Annals of Statistics, 30(6):1666–1680, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://www.jstor.org/stable/1558736.

Anonymous:2007:BRBn

[Ano07] Anonymous. Book review: Exploring Data Tables, Trends, and Shapes by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Technometrics, 49(3):368, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471376.

Anscombe:1965:CKL

[Ans65] F. J. Anscombe. Comments on Kurtz–Link–Tukey–Wallace paper. Technometrics, 7(2):167–168, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:/

/www.jstor.org/stable/1266668. See [KLTW65b, KLTW65c,

KLTW65a].

Anscombe:1988:FMJ

[Ans88] Francis J. Anscombe. Frederick Mosteller and John W. Tukey: a conversation. Statistical Science, 3(1):136–144, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.ss/
1177013020;   http://www.jstor.org/stable/2245933.

Anscombe:2003:QCC

[Ans03] F. R. Anscombe. Quiet contributor: The civic career and times of John W. Tukey. Statistical Science, 18(3):287–310, August 2003.

CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1076102417.

Aroian:1947:NCF

[Aro47] Leo A. Aroian. Note on the cumulants of Fisher’s z-distribution. Biometrika, 34(3/4):359–360, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332447.

Arsenin:1940:NPC

[Ars40] V. Arsenin. Sur la nature des projections de certains ensembles mesurables B. Bull. Acad. Sci. URSS. Sér. Math. [Izvestia Akad. Nauk SSSR], 4:403–410, 1940.

Aspray:1984:OHI

[Asp84] William Aspray. Oral history interview with John Wilder Tukey and Albert Tucker, 1984 April 11. Princeton mathematics community in the 1930s: an oral history project, 1984. URL https:/

/web.math.princeton.edu/oral-history/c39.pdf. Transcript

Number 41 (PMC41).

Abelson:1959:ECN

[AT59] R. P. Abelson and John W. Tukey. Efficient conversion of nonmetric information into metric information. In Proceedings of the Social Statistics Section, volume 1959, pages 226–230. American Statistical Association, Washington, DC, USA, 1959.

Abelson:1963:EUN

[AT63a] Robert P. Abelson and John W. Tukey. Efficient utilization of non-numerical information in quantitative analysis: General theory and the case of simple order. Annals of Mathematical Statistics, 34(4):1347–1369, ???? 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703869.

Anscombe:1963:EAR

[AT63b] F. J. Anscombe and John W. Tukey. The examination and analysis of residuals. Technometrics, 5(2):141–160, May 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266059. See corrections

[Olm63, De 64].

Andrews:1973:SAA

[AT73a] David F. Andrews and John W. Tukey. Statistical algorithms: Algorithm AS 61: Six-line plots (AS R23: 77V26 p368). Applied Statistics, 22(2):265–269, June 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/61. See remark [Ley77].

Andrews:1973:TPD

[AT73b] David F. Andrews and John W. Tukey. Teletypewriter plots for data analysis can be fast: 6-line plots, including probability plots. Applied Statistics, 22(2):192–202, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://www.

jstor.org/stable/10.2307/j.ctt13x12sw.

Aspray:1985:JWT

[AT85] William Aspray and Albert Tucker. John W. Tukey interview. In The Princeton Mathematics Community in the 1930s,

page ?? ????, ????, 1985. URL http://libweb.princeton.edu/
libraries/firestone/rbsc/finding_aids/mathoral/pmO2.htm#
transcriptlist.

Balandin:1939:CIE

[Bal39] A. A. Balandin. Calculation of the intermediate equilibria in catalytic kinetics and the geometry situs. C. R. (Doklady) Acad. Sci. URSS (N.S.), 24:741–747, 1939.

Balandin:1940:SAC

[Bal40] A. A. Balandin. Structural algebra in chemistry. Acta Physicochim. URSS, 12:447–479, 1940.

Banos:1946:CSP

[Bañ46] O. Fernández Baños. Contribution to the study of Pearson’s χ^{2}. Revista Mat. Hisp.-Amer. (4), 6:66–83, 1946.

Banerjee:1949:PPI

[Ban49] D. P. Banerjee. On percentage points of incomplete beta-functions and χ^{2} distribution. Bull. Calcutta Math. Soc., 41:53–54, 1949. ISSN 0008-0659.

Banerjee:1958:II

[Ban58] D. P. Banerjee. On inverse interpolation. Bull. Calcutta Math. Soc., 1958(Supplement):54–57, 1958. ISSN 0008-0659.

Bartlett:1946:MPT

[Bar46] M. S. Bartlett. A modified probit technique for small probabilities. Suppl. J. Roy. Statist. Soc., 8:113–117, 1946.

Bartlett:1947:MA

[Bar47] M. S. Bartlett. Multivariate analysis. Suppl. J. Roy. Statist. Soc., 9:176–190; discussion 190–197, 1947.

Barnett:1981:IMD

[Bar81] Vic Barnett, editor. Interpreting Multivariate Data: Proceedings of the conference entitled “Looking at Multivariate Data” held in the University of Sheffield, U.K. from 24-27 March 1980. Wiley, New York, NY, USA, 1981. ISBN 0-471-28039-9. LCCN QA278 .I59.

Benjamini:2002:JWT

[BB02] Yoav Benjamini and Henry Braun. John W. Tukey’s contributions to multiple comparisons. Annals of Statistics, 30(6):1576–1594, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1043351247;   http://www.jstor.org/stable/1558730.

Brooks:1959:PDB

[BBB59] Frederick P. Brooks, Jr., Gerrit A. Blaauw, and Werner Buchholz. Processing data in bits and pieces. IRE Transactions on Electronic Computers, EC-8(2):118–124, June 1959. CODEN IRELAO. ISSN

0367-9950. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5219512. This paper contains on page 121 the first published reference to the term “byte”. An article of the same title appears in “Information Processing, Proceedings of the International Conference on Information Processing, UN-ESCO, Paris, 15–20 June 1959”, pp. 375–381, 1959. From [Buc62, page 40]: “Byte denotes a group of bits used to encode a character, or the number of bits transmitted in parallel to and from inputoutput units. A term other than character is used here because a given character may be represented in different applications by more than one code, and different codes may use different numbers of bits (i.e., different byte sizes). In input-output transmission the grouping of bits may be completely arbitrary and have no relation to actual characters. (The term is coined from bite, but respelled to avoid accidental mutation to bit.)”.

Becker:1993:DSG

[BC93] Richard A. Becker and William S. Cleveland. Discussion of “Graphic Comparisons of Several Linked Aspects” by John W. Tukey. Journal of Computational and Graphical Statistics, 2(1): 41–48, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/10618600.1993.10474597. See [Tuk93b].

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. SIAM Press, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Brillinger:1992:NDT

[BCG^{+}93] David R. Brillinger, P. Caines, J. Geweke, E. Parzen, M. Rosenblatt, and M. Taqqu, editors. New directions in time series analysis, Part 1, volume 45–46 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992/1993. ISBN 0-387-97896-8 (vol. 1), 0-387-97914-X (vol. 2). LCCN QA280 .N47 1992.

Breslow:1986:NVV

[BCH^{+}86] N. Breslow, C. L. Chiang, J. M. Hoem, P. Jagers, N. Keiding, L. Kish, K. G. Manton, and J. W. Tukey. The natural variability of vital-rates and associated statistics — discussion. Biometrics, 42(4):712–732, ???? 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bliss:1956:RCB

[BCT56] C. I. Bliss, W. G. Cochran, and J. W. Tukey. A rejection criterion based upon the range. Biometrika, 43(3/4):418–422, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332919.

Beaumont:1986:BRB

[Bea86] Chris Beaumont. Book review: Exploring Data Tables, Trends and Shapes, by D. C. Hoaglin; F. Mosteller; J. W. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 35(3):

397, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.org/stable/2987762.

Berry:1945:CCC

[Ber45] Clifford E. Berry. A criterion of convergence for the classical iterative method of solving linear simultaneous equations. Annals of Mathematical Statistics, 16:398–400, 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Berkson:1946:ACS

[Ber46] Joseph Berkson. Approximation of chi-square by “probits” and by “logits”. Journal of the American Statistical Association, 41 (233):70–74, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280157.

Bertin:1983:SGDj

[Ber83] Jacques Bertin. Semiology of Graphics: Diagrams, Networks, Maps. University of Wisconsin Press, Madison, WI, USA, 1983. ISBN 0-299-09060-4. xi + 415 pp. LCCN QA90 .B47513 1983.

Bertin:2010:PTF

[Ber10] Jacques Bertin. Semiology of Graphics: Diagrams, Networks, Maps. Esri Press, Redlands, CA, USA, 2010. ISBN 1-58948-261-1 (hardcover). xv + 438 pp. LCCN QA90 .B4713 2011.

Brillinger:1997:PDA

[BFM97] David R. Brillinger, Luisa Turrin Fernholz, and Stephan Morgenthaler, editors. The practice of data analysis: essays in honor of John W. Tukey. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-05782-6. LCCN QA276.16

.P73 1997. URL http://www.loc.gov/catdir/description/
prin021/97019695.html;   http://www.loc.gov/catdir/toc/

prin031/97019695.html. Two-day symposium, Princeton on June 19–20, 1995.

Bingham:1967:MTP

[BGT67] Christopher Bingham, Michael D. Godfrey, and John W. Tukey. Modern techniques of power spectrum estimation. IEEE Transactions on Audio and Electroacoustics, 15(2):56–66, ???? 1967. CO-DEN ITADAS. ISSN 0018-9278 (print), 1558-2582 (electronic). Reprinted in [Bri84a, pp. 781–810].

Burks:1946:PDL

[BGvN46] Arthur W. Burks, Herman H. Goldstine, and John von Neumann. Preliminary discussion of the logical design of an electronic computing instrument. Technical report, Institute for Advanced Study, Princeton, NJ, USA, 1946. 42 pp. Report to the U.S. Army Ordnance Department under contract W-36-034-OKD-7481. Reprinted in [?, Paper 2], [?], [?, pp. 221–259] and [?, pp. 97–146].

Bogert:1963:FAT

[BHT63] B. P. Bogert, M. J. R. Healy, and J. W. Tukey. The quefrency alanysis of time series for echoes: cepstrum, pseudo-autocovariance, cross-cepstrum, and saphe cracking. In Rosenblatt [Ros63], pages 209–243. LCCN QA280.

Brisebarre:2020:EAS

[BJM^{+}20] Nicolas Brisebarre, Mioara Joldes, Jean-Michel Muller, Ana-Maria Nanes, and Joris Picot. Error analysis of some operations involved in the Cooley–Tukey Fast Fourier Transform. ACM Transactions on Mathematical Software, 46(2):11:1–11:27, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3368619.

Brillinger:1978:MWR

[BJT78] D. R. Brillinger, L. V. Jones, and John W. Tukey. The Management of Weather Resources II: The Role of Statistics in Weather Resources Management. United States Government Printing Office, Washington, DC, USA, 1978. ???? pp. LCCN QC928.7 .U55 1978.

Bartlett:1946:SAV

[BK46] M. S. Bartlett and D. G. Kendall. The statistical analysis of variance-heterogeneity and the logarithmic transformation. Suppl. J. Roy. Statist. Soc., 8:128–138, 1946.

Boisvert:2001:HMF

[BL01] Ronald F. Boisvert and Daniel W. Lozier. Handbook of Mathematical Functions. In Lide [Lid01], pages 135–139.

URL https://nvlpubs.nist.gov/nistpubs/sp958-lide/135-
139.pdf;   https://nvlpubs.nist.gov/nistpubs/sp958-lide/

html/135-139.html. NIST Special Publication.

Black:1947:RJW

[Bla47] Max Black. Review: John W. Tukey, The Princeton University Bicentennial Conference on the Problems of Mathematics. Journal of Symbolic Logic, 12(3):89, ???? 1947. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL http://www.

jstor.org/stable/2267218. See [LT46].

Beh:2018:CAF

[BLA18] Eric J. Beh, Rosaria Lombardo, and Gianmarco Alberti. Correspondence analysis and the Freeman–Tukey statistic: a study of archaeological data. Computational Statistics & Data Analysis, 128(??):73–86, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167947318301580.

Blumenthal:1941:NCD

[Blu41] Leonard M. Blumenthal. A new concept in distance geometry with applications to spherical subsets. Bulletin of the American Mathematical Society, 47:435–443, 1941. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Bode:1949:ESG

[BMTW49] Hendrik W. Bode, Frederick Mosteller, John W. Tukey, and Charles P. Winsor. The education of a scientific generalist. Science, 109(2840):553–558, June 3, 1949. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Boardman:1985:BRBa

[Boa85] Thomas J. Boardman. Book review: The Collected Works of John W. Tukey, Volume I, Time Series: 1949–1964 by David R. Brillinger; John W. Tukey. Technometrics, 27(3):319, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269718.

Bochner:1938:LCA

[Boc38] S. Bochner. Lectures on commutative algebra, 1937–1938. Planographed report, Edwards Brothers, Inc., Ann Arbor, MI, USA, 1938. vii + 76 pp. Notes by J. W. Tukey, J. Giese, and V. Martin.

Bose:1947:RFT

[Bos47] P. K. Bose. On recursion formulae, tables and Bessel function populations associated with the distribution of classical D^{2}-statistic. Sankhyā, 8:235–248, 1947. ISSN 0972-7671.

Bryc:1992:CLT

[BP92] Włodzimierz Bryc and Magda Peligrad. The Central Limit Theorem for Tukey’s 3R smoother. Statistics & Probability Letters, 13(1):29–37, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290233U.

Bruce:1981:SRF

[BPT81] Andrew Bruce, Daryl Pregibon, and John W. Tukey. The second representing function for compound situations. Technical report, series 2 186, Princeton University, Princeton, NJ, USA, 1981. 11 pp.

Braun:1994:CWJ

[Bra94] Henry I. Braun, editor. The collected works of John W. Tukey. Volume VIII. Multiple comparisons: 1948–1983. Chapman and Hall, Ltd., London, UK, 1994. ISBN 0-412-05121-4. lxii + 475 + 10 pp. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Brillinger:1964:ABT

[Bri64] D. R. Brillinger. The asymptotic behaviour of Tukey’s general method of setting approximate confidence limits (the jackknife) when applied to maximum likelihood estimates. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):202–206, ???? 1964. CODEN ???? ISSN 0373-1138.

Brillinger:1984:CWJa

[Bri84a] David R. Brillinger, editor. The collected works of John W. Tukey. Volume I. Time series: 1949–1964. Wadsworth Statistics/Probability Series. Wadsworth, Pacific Grove, CA, USA, 1984. ISBN 0-534-03303-2. lxv + 650 + 39 pp. LCCN QA276.A12 T85 1984. With introductory material by William S. Cleveland and Frederick Mosteller.

Brillinger:1984:CWJb

[Bri84b] David R. Brillinger, editor. The collected works of John W. Tukey. Volume II. Time series: 1965–1984. Wadsworth Statistics/Probability Series. Wadsworth, Pacific Grove, CA, USA, 1984. ISBN 0-534-03304-0. lxvii + 650–1153 + 80 pp. With introductory material by William S. Cleveland and Frederick Mosteller.

Brillinger:2002:JWTa

[Bri02a] D. R. Brillinger. John W. Tukey: His life and professional contributions. Annals of Statistics, 30(6):1535–1575, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1043351246;   http://www.jstor.org/stable/1558729;
 http://www.stat.berkeley.edu/users/brill/Papers/life.
pdf.

Brillinger:2002:JWTb

[Bri02b] D. R. Brillinger. John W. Tukey’s work on time series and spectral analysis. Annals of Statistics, 30(6):1595–1618, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1043351248;   http://www.jstor.org/stable/1558731;
 http://www.stat.berkeley.edu/users/brill/Papers/time.
pdf..

Brillinger:2002:MJW

[Bri02c] David R. Brillinger. Memories of John W. Tukey. Web site.,

February 18, 2002. URL http://cm.bell-labs.com/cm/ms/
departments/sia/tukey/tributes.html.

Brillinger:2006:DTP

[Bri06] David R. Brillinger. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):325–327, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471201. See [Mal06].

Brillinger:2009:JWT

[Bri09] David R. Brillinger. John Wilder Tukey 1915–2000: a biographical memoir. National Academy of Sciences, ??(??):1–18, ????

2009. URL http://www.stat.berkeley.edu/~brill/Papers/
TukeyJohnWBMemoir.pdf.

Bebutoff:1939:DTS

[BS39] M. Bebutoff and V. Schneider. On a denumerable topological space. Uchenye Zapiski Moskov. Gos. Univ. Matematika, 30:157– 160, 1939.

Boas:1938:LAW

[BST38] Ralph P. Boas, Frank Smithies, and John W. Tukey. Lectures by Aurel Wintner on asymptotic distributions and infinite convolutions, 1937–1938. Planographed report, Edwards Brothers, Inc., Ann Arbor, MI, USA, 1938. 54 pp. With sympathetic encouragement from Cyrus C. MacDuffee.

Boas:1938:NLF

[BT38] R. P. Boas, Jr. and John W. Tukey. A note on linear functionals. Bulletin of the American Mathematical Society, 44(8):523–528, ???? 1938. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.

bams/1183500620. See correction [BT40b].

Blumenthal:1940:CPS

[BT40a] Leonard M. Blumenthal and George R. Thurman. The characterization of pseudo-spherical sets. American Journal of Mathematics, 62:835–854, 1940. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic).

Boas:1940:CNL

[BT40b] R. P. Boas, Jr. and J. W. Tukey. A correction to “A note on linear functionals”. Bulletin of the American Mathematical Society, 46(6):566, 1940. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://projecteuclid.

org/euclid.bams/1183502728. See [BT38].

Brown:1946:SDS

[BT46] George W. Brown and John W. Tukey. Some distributions of sample means. Annals of Mathematical Statistics, 17(1):1–12, ???? 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.
aop/1176993612.

Blackman:1958:MPSa

[BT58a] R. B. Blackman and J. W. Tukey. The measurement of power spectra from the point of view of communications engineering — Part I.

The Bell System Technical Journal, 37(1):185–282, January 1958. CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.bell-

labs.com/BSTJ/images/Vol37/bstj37-1-185.pdf;   http://
www.alcatel-lucent.com/bstj/vol37-1958/articles/bstj37-
1-185.pdf.

Blackman:1958:MPSb

[BT58b] R. B. Blackman and J. W. Tukey. The measurement of power spectra from the point of view of communications engineering — Part II. The Bell System Technical Journal, 37(2):485–569, March 1958. CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.bell-

labs.com/BSTJ/images/Vol37/bstj37-2-485.pdf;   http://
www.alcatel-lucent.com/bstj/vol37-1958/articles/bstj37-
2-485.pdf.

Blackman:1959:MPS

[BT59] R. B. Blackman and J. W. Tukey. The measurement of power spectra from the point of view of communications engineering. Dover Publications, Inc., New York, NY, USA, 1959. x + 190 pp.

Bingham:1966:FMF

[BT66] C. Bingham and J. W. Tukey. Fourier methods in the frequency analysis of data. ???? ??, ????, ????, 1966.

Beaton:1974:FPS

[BT74a] A. E. Beaton and J. W. Tukey. The fitting of power series, meaning polynomials, illustrated on band-spectroscopic data. Technometrics, 16(2):147–185, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267936.

Beaton:1974:CFP

[BT74b] A. E. Beaton and John W. Tukey. Comments on the fitting of power series. In D. R. Lide, Jr. and M. A. Paul, editors, Critical Evaluation of Chemical and Physical Structural Information, pages 15–35. National Academy of Sciences, Washington, DC, USA, 1974.

Breckenridge:1978:EHR

[BT78] Mary B. Breckenridge and John W. Tukey. An empirical higherrank analysis model of the age distribution of fertility. Technical report, series 2 143, Princeton University, Princeton, NJ, USA, 1978. 95 pp.

Braun:1983:MCT

[BT83] H. I. Braun and John W. Tukey. Multiple comparisons through orderly partitions: The maximum subrange procedure. In H. Wainer and S. Messick, editors, Principals of Modem Psychological Measurement: A Festschrift for Frederic M. Lord, pages 55–65. Erlbaum, Hillsdale, NJ, USA, 1983.

Brillinger:1984:SAP

[BT84] D. R. Brillinger and John W. Tukey. Spectrum analysis in the presence of noise: some issues and examples. In Brillinger [Bri84b], pages 1001–1141. ISBN 0-534-03304-0. With introductory material by William S. Cleveland and Frederick Mosteller.

Brillinger:1985:SAP

[BT85] D. R. Brillinger and J. W. Tukey. Spectrum analysis in the presence of noise: Some issues and examples. ???? ??, ????, ????, 1985. Reprinted in [Bri84b, pp. 1001–1141]. Was this a report, or some other kind of document?

Basford:1995:IRF

[BT95] K E Basford and J. W. Tukey. Intercomparing residuals to find outliers in randomized blocks. Australian Journal of Agricultural Research, 46(2):451–461, ???? 1995. CODEN AJAEA9. ISSN 0004-9409 (print), 1444-9838 (electronic).

Basford:1997:GPA

[BT97] K. E. Basford and John W. Tukey. Graphical profiles as an aid to understanding plant breeding experiments. Journal of Statistical Planning and Inference, 57(??):93–107, ???? 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Basford:1999:GAM

[BT99] Kaye E. Basford and John W. Tukey. Graphical Analysis of Multiresponse Data: Illustrated with a Plant Breeding Trial: Interdisciplinary Statistics. Chapman and Hall/CRC, Boca Raton, FL, USA, 1999. ISBN 0-8493-0384-2, 0-412-81890-6, 0-367-81184-7 (ebook), 1-00-071520-5 (PDF), 1-00-072302-X (EPUB). xvi + 587 pp. LCCN QA278 .B3745 1999.

Bloomberg:1995:DFW

[BTW95] Dan S. Bloomberg, John W. Tukey, and M. Margaret Withgott. Detecting function words without converting a scanned document to character codes. US Patent 5,455,871., 1995. URL http://

www.google.com/patents/US5455871. Filed 16 May 1994. Issued

3 October 1995.

Buchholz:1962:PCS

[Buc62] Werner Buchholz, editor. Planning a computer system: Project Stretch. McGraw-Hill, New York, NY, USA, 1962. xvii + 322

pp. LCCN 1876. URL http://ed-thelen.org/comp-hist/IBM-

7030-Planning-McJones.pdf. This important book is the primary description of the influential IBM 7030 Stretch computer, written by its architects.

Buchholz:1981:AOW

[Buc81] W. Buchholz. Anecdotes: Origin of the word byte. Annals of the History of Computing, 3(1):72, January/March 1981. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1981/pdf/a1072a.pdf;   http://www.computer.
org/annals/an1981/a1072aabs.htm. Historical notes: The word

“byte” was first used in a published article by Brooks, Blaauw, and Buchholz in 1959 [BBB59]. The word “software” was coined by John W. Tukey [Tuk58e]. The word “bit” was also coined by Tukey in 1946 at Bell Laboratories. See also [TB84].

Buja:2006:DTP

[Buj06] Andreas Buja. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):327–330, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471202. See [Mal06].

Carson:1962:SS

[Car62] Rachel Carson. Silent Spring. Fawcett Publications, Greenwich, CT, 1962. 304 pp. LCCN ????

Cassina:1943:SNM

[Cas43] Ugo Cassina. Su un nuovo metodo per la risoluzione numerica delle equazioni algebriche o trascendenti. Ist. Lombardo Sci. Lett. Cl. Sci. Mat. Nat. Rend. (3), 7(76):35–61, 1943.

Casella:2003:FTJ

[Cas03] George Casella. Foreword [to tribute to John Tukey]. Statistical Science, 18(3):283–284, August 2003. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1076102415.

Chabert:1999:HAP

[CBB^{+}99] Jean-Luc Chabert, Évelyne Barbin, Jacques Borowczyk, Michel Guillemot, Anne Michel-Pajus, Ahmed Djebbar, and Jean-Claude Martzloff, editors. A history of algorithms: from the pebble to the microchip. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-63369-3 (softcover). ix + 524 pp. LCCN QA58 .H5813

1998. URL http://www.springer.com/west/home/math/cse?
SGWID=4-10045-22-1455224-0.

Chen:1998:AMG

[CBT98] Francine R. Chen, Dan S. Bloomberg, and John W. Tukey. Automatic method of generating thematic summaries from a document image without performing character recognition. US

Patent 5,848,191., 1998. URL http://www.google.com/patents/

US5848191. Filed 14 December 1995. Issued 8 December 1998.

Cohen:1993:RSH

[CDT93] Michael Cohen, Siddhartha R. Dalal, and John W. Tukey. Robust, smoothly heterogeneous variance regression. Applied Statistics, 42 (2):339–353, 1993. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1992:BRB

[Cha92] Chris Chatfield. Book review: Computing and Graphics in Statistics, by A. Buja; P. A. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 42(3):326–327, ???? 1992. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/2348817.

Chambers:1998:CDC

[Cha98] John Chambers. Computing with data: Concepts and challenges. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1998.

URL http://cm.bell-labs.com/stat/doc/Neyman98.ps. Ex-

tended version of [Cha99].

Chambers:1999:CDC

[Cha99] John Chambers. Computing with data: Concepts and challenges. The American Statistician, 53(1):73–84, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685657. Neyman Lecture pre-

sented at the 1988 Joint Statistical Meetings. An extended version is available in [Cha98].

Cheng:1943:SFM

[Che43] Tseng-Tung Cheng. A simplified formula for mean difference. Coll. Papers Sci. Engin. Nat. Univ. Amoy, 1:69–72, 1943.

Ciminera:1992:EMC

[CHNT92a] J. L. Ciminera, J. F. Heyse, H. H. Nguyen, and John W. Tukey. Evaluation of multicenter clinical trial data using adaptations of the Mosteller–Tukey procedure. Statistics in Medicine, 12(??): 1047–1061, ???? 1992. CODEN ???? ISSN 0277-6715.

Ciminera:1992:TQT

[CHNT92b] J. L. Ciminera, J. F. Heyse, H. H. Nguyen, and John W. Tukey. Tests for qualitative treatment-by-center interaction using a “pushback” procedure. Statistics in Medicine, 12(??):1033–1045, ???? 1992. CODEN ???? ISSN 0277-6715.

Ciminera:1993:EMC

[CHNT93a] J. L. Ciminera, J. F. Heyse, H. H. Nguyen, and J. W. Tukey. Evaluation of multicenter clinical-trial data using adaptations of the Mosteller–Tukey procedure. STATISTICS IN MEDICINE, 12 (11):1047–1061, ???? 1993. ISSN 0277-6715.

Ciminera:1993:TQT

[CHNT93b] J. L. Ciminera, J. F. Heyse, H. H. Nguyen, and J. W. Tukey. Tests for qualitative treatment-by-center interaction using a pushback procedure. Statistics in Medicine, 12(11):1033–1045, ???? 1993. ISSN 0277-6715.

Ciminera:1989:TGA

[CHT89] J. L. Ciminera, J. F. Heyse, and John W. Tukey. Techniques for the graphical analysis of possibly skewed measurements. Journal of Quality Technology, 21(??):223–231, ???? 1989. CODEN JQUTAU. ISSN 0022-4065.

Cox:2000:EEP

[CHT00] J. L. Cox, J. E. Heyse, and J. W. Tukey. Efficacy estimates from parasite count data that include zero counts. Experimental Parasitology, 96(1):1–8, ???? 2000. CODEN EXPAAA. ISSN 0014-4894 (print), 1090-2449 (electronic).

Cipra:2000:BCE

[Cip00] Barry A. Cipra. The best of the 20th Century: Editors name top 10 algorithms. SIAM News, 33(4):1–2, May 2000. ISSN 0036-1437.

URL https://archive.siam.org/pdf/news/637.pdf.

Clairaut:1754:MOA

[Cla54] A. C. Clairaut. Mémoire sur l’orbite apparente du soleil autour de la terre, en ayant égard aux perturbations produites par les actions de la lune et des planètes principales. (French). [Memoir on the apparent orbit of the Sun around the Earth, with respect to perturbations produced by the Moon and the main planets]. Mémoires de mathématique et de physique de l’Académie Royale des Sciences, 9(??):521–564, 1754. This paper is credited as having the earliest known formula for the DFT (Discrete Fourier Transform) [HJB84, page 15].

Cleveland:1988:CWJ

[Cle88] William S. Cleveland, editor. The collected works of John W. Tukey. Volume V. Graphics: 1965–1985. Wadsworth & Brooks/ Cole Statistics/Probability Series. Wadsworth & Brooks/Cole Advanced Books & Software, Monterey, CA, USA, 1988. ISBN 0-534-05102-2. lxiv + 436 + i28 pp. With a biography of Tukey by Frederick Mosteller.

Cooley:1967:AFF

[CLW67a] James W. Cooley, Peter A. W. Lewis, and Peter D. Welch. Application of the Fast Fourier Transform to computation of Fourier integrals, Fourier series, and convolution integrals. IEEE Transactions on Audio and Electroacoustics, AU-15(2):79–84, June 1967. CODEN ITADAS. ISSN 0018-9278 (print), 1558-2582 (electronic).

Cooley:1967:HNF

[CLW67b] James W. Cooley, Peter A. W. Lewis, and Peter D. Welch. Historical notes on the Fast Fourier Transform. IEEE Transactions on Audio and Electroacoustics, AU-15(2):76–79, June 1967. CO-DEN ITADAS. ISSN 0018-9278 (print), 1558-2582 (electronic). Reprinted in Proc. IEEE 55(10) 1675–1677, October 1967, and in Digital Signal Processing, ed., L. R. Rabiner and C. M. Rader, IEEE Press, New York (1972), pp. 260–262.

Cleveland:1988:DGS

[CM88] William S. Cleveland and Marylyn E. McGill, editors. Dynamic graphics for statistics. The Wadsworth and Brooks/Cole statistics/probability series. Wadsworth & Brooks/Cole Advanced Books & Software, Monterey, CA, USA, 1988. ISBN 0-534-09144-X. xii + 424 + 16 pp. LCCN QA276.3 .D96 1988. URL http://

www.loc.gov/catdir/enhancements/fy0744/88026017-d.html.

Chang:1976:NIT

[CMAT76] J. J. Chang, M. V. Mathews, B. Atal, and J. W. Tukey. Numerical inversion of transformation from articulatory to acoustic parameters in vocal-tract. Journal of the Acoustical Society of America, 60(??):S77, ???? 1976. CODEN JASMAN. ISSN 0001-4966.

Cochran:1953:SPK

[CMT53] William G. Cochran, Frederick Mosteller, and John W. Tukey. Statistical problems of the Kinsey report. Journal of the American Statistical Association, 48(264):673–716, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281066.

Cochran:1954:SPK

[CMT54a] W. G. Cochran, F. Mosteller, and John W. Tukey. Statistical problems of the Kinsey report on sexual behavior in the human male. Report, American Statistical Association, Washington, DC, USA, 1954.

Cochran:1954:PS

[CMT54b] William G. Cochran, Frederick Mosteller, and John W. Tukey. Principles of sampling. Journal of the American Statistical Association, 49(265):13–35, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281032.

Cochran:1955:CMT

[CMT55] W. G. Cochran, F. Mosteller, and John W. Tukey. The Cochran– Mosteller–Tukey report on the Kinsey Study: A symposium (pp. 811–829). Journal of the American Statistical Association, 50 (271):829, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281167.

Cochran:1956:FM

[CMT56] W. G. Cochran, F. Mosteller, and John W. Tukey. Fundamentos de muestreo. J. Inter. Amer. Statist. Inst., ??(??):235–258, June 1956. CODEN ???? ISSN ????

Cohen:1940:TC

[Coh40] L. W. Cohen. On topological completeness. Bulletin of the American Mathematical Society, 46:706–710, 1940. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Collett:1984:BRB

[Col84] D. Collett. Book review: Understanding Robust and Exploratory Data Analysis, by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Journal of the Royal Statistical Society. Series A (General), 147(1):113, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2981753.

Cooley:1987:HFG

[Coo87a] J. W. Cooley. How the FFT gained acceptance. In Crane [Cra87], pages viii + 150. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Cooley:1987:RDF

[Coo87b] J. W. Cooley. The re-discovery of the Fast Fourier Transform algorithm. Microchimica Acta, 93(1–6):33–45, ???? 1987. CODEN MIACAQ. ISSN 0026-3672 (print), 1436-5073 (electronic).

Cooley:1990:HFG

[Coo90] J. W. Cooley. How the FFT gained acceptance. In Nash [Nas90], pages 133–140. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990. Nowack [Now94, page 395] quotes Cooley from this paper: it appears that Lanczos had the FFT algorithm; and if he had had an electronic computer, he would have been able to write a program permitting him to go to arbitrarily high N.

Cooley:1994:DF

[Coo94a] James W. Cooley. Development of the FFT. In Brown et al. [BCEP94], pages 393–394. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Cooley:1994:LFD

[Coo94b] James W. Cooley. Lanczos and the FFT: a discovery before its time. In Brown et al. [BCEP94], pages 3–9. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993. Plenary presentations, computational mathematics.

Cooper:2000:BRG

[Coo00] Mark Cooper. Book review: Graphical Analysis of Multiresponse Data: Illustrated with A Plant Breeding Trial, by K. E. Basford and J. W. Tukey. Crop science, 40(4):1184, ???? 2000.

Cotlar:1940:NMS

[Cot40] Mischa Cotlar. On non-measurable sets and a generalization of the Lebesgue integral. Publ. Inst. Mat. Univ. Nac. Litoral, 2:149–176, 1940.

Cox:1992:CWJ

[Cox92] D. R. Cox, editor. The collected works of John W. Tukey. Volume VII. Factorial & ANOVA (1949–1962). Wadsworth & Brooks/ Cole Statistics/Probability Series. Wadsworth & Brooks/Cole Advanced Books & Software, Monterey, CA, USA, 1992. ISBN 0-534-05104-9. lxvi + 260 + I6 pp. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Cox:2016:SPM

[Cox16] D. R. Cox. Some pioneers of modern statistical theory: a personal reflection. Biometrika, 103(4):747–759, December 2016. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://academic.oup.com/biomet/article-abstract/
103/4/747/2659040/Some-pioneers-of-modern-statistical-
theory-a.

Cutting:1992:SGC

[CPKT92] Douglas R. Cutting, Jan O. Pedersen, David Karger, and John W. Tukey. Scatter/Gather: a cluster-based approach to browsing large document collections. In Nicholas J. Belkin, Peter Ingwersen, and Annelise Mark Pejtersen, editors, Proceedings of the Fifteenth Annual International ACM SIGIR Conference on Research and Development in Information Retrieval, pages 318–329. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-523-2. LCCN Z699.A1 I659 1992g. URL http://citeseer.nj.nec.

com/cutting92scattergather.html.

Crane:1987:HAC

[Cra87] G. E. Crane, editor. HSNC’87: ACM Conference on the History of Scientific and Numeric Computation, conference proceedings: papers presented at the Conference, Princeton, New Jersey, May 13-15, 1987. ACM Press, New York, NY 10036, USA, October 1987. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Crow:2001:SBF

[Cro01] James F. Crow. Shannon’s brief foray into genetics. Genetics, 159(3):915–917, ???? 2001. CODEN GENTAE. ISSN 0016-6731 (print), 1943-2631 (electronic). URL http://www.genetics.org/

content/159/3/915.full.

Cornfield:1956:AVM

[CT56] Jerome Cornfield and John W. Tukey. Average values of mean squares in factorials. Annals of Mathematical Statistics, 27(4):907– 949, ???? 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728067.

Cooley:1965:AMC

[CT65] James W. Cooley and John W. Tukey. An algorithm for the machine calculation of complex Fourier series. Mathematics of Computation, 19(90):297–301, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www. jstor.org/stable/2003354. Reprinted in [Bri84b, pp. 651–658 or 659–675] and [KJ97]. See [CLW67a, CLW67b, Hua71, HJB85, Coo87a, Coo87b, CT93, CBB^{+}99].

Cooley:1975:AMC

[CT75] J. W. Cooley and John W. Tukey. An algorithm for the machine calculation of complex Fourier series. In Bede Liu, editor, Digital Filters and the Fast Fourier Transform, volume 12 of Benchmark papers in electrical engineering and computer science, pages 328– 332. Dowden, Hutchinson and Ross, Stroudsburg, PA, USA, 1975. ISBN 0-470-54150-4. LCCN TK7872.F5 D53.

Ciminera:1989:CCA

[CT89a] J. L. Ciminera and John W. Tukey. Control-charting automated laboratory instruments when many successive differences may be zero. Journal of Quality Technology, 21(??):7–15, ???? 1989. CO-DEN JQUTAU. ISSN 0022-4065.

Ciminera:1989:UAB

[CT89b] J. L. Ciminera and John W. Tukey. Use of augmented Box plots for reviewing measurement results collected over time. Journal of Quality Technology, 21(??):187–198, ???? 1989. CODEN JQUTAU. ISSN 0022-4065.

Cooley:1993:OPF

[CT93] James W. Cooley and John W. Tukey. On the origin and publication of the FFT paper. Current Contents: Physical, Chemical & Earth Sciences, ??(51–52):8–9, December 1993. CODEN ???? ISSN 0163-2574. See [CT65].

Chen:1998:AMI

[CT98] Francine R. Chen and John W. Tukey. Automatic method of identifying drop words in a document image without performing character recognition. US Patent 5,850,476., 1998. URL http://www.

google.com/patents/US5850476. Filed 14 December 1995. Issue

15 December 1998.

Chen:2002:IFM

[CT02] Zhiqiang Chen and David E. Tyler. The influence function and maximum bias of Tukey’s median. Annals of Statistics, 30(6): 1737–1759, ???? 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1043351255.

Cuesta:1947:NSW

[Cue47a] N. Cuesta. Notes on some works of Sierpiński. Revista Mat. Hisp.- Amer. (4), 7:128–131, 1947.

Cuesta:1947:NTO

[Cue47b] N. Cuesta. Number of types of ordering. Revista Mat. Hisp.-Amer. (4), 7:3–9, 1947.

Danielson:1940:IPF

[Dan40] Gordon Charles Danielson. Improvement in practical Fourier analysis and their application to X-ray scattering from liquids. Ph.D. thesis, Purdue University, West Lafayette, IN, USA, 1940. 95 pp.

Davison:1988:BRB

[Dav88a] A. C. Davison. Book review: Exploring Data Tables, Trends and Shapes, by D. C. Hoaglin; F. Mosteller; J. W. Tukey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):228–229, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.jstor.org/

stable/2982202.

Davison:1988:RED

[Dav88b] A. C. Davison. Review: Exploring Data Tables, Trends and Shapes, D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):228–229, ???? 1988. CODEN ???? ISSN 0964-1998 (print), 1467-985X

(electronic). URL http://www.jstor.org/stable/2982202.

DeBaun:1964:CEA

[De 64] R. M. De Baun. Correction: The Examination and Analysis of Residuals, by F. J. Anscombe and John W. Tukey. Technometrics, 6(1):127, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266059.

Dempster:2002:JWT

[Dem02] A. P. Dempster. John W. Tukey as “philosopher”. Annals of Statistics, 30(6):1619–1628, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1043351249;   http://www.
jstor.org/stable/1558732.

Dutta:2011:SIP

[DGC11] Subhajit Dutta, Anil K. Ghosh, and Probal Chaudhuri. Some intriguing properties of Tukey’s half-space depth. Bernoulli, 17 (4):1420–1434, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1320417511.

Dresden:1940:APS

[DGP40] Arnold Dresden, Margaret Gurney, and E. S. Pondiczery. Advanced problems and solutions: Solutions: 3868. American Mathematical Monthly, 47(4):249–250, April 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [?].

Dieudonne:1939:EUC

[Die39a] J. Dieudonné. Sur les espaces uniformes complets. Ann. École Norm., 56:277–291, 1939.

Dieudonne:1939:ETS

[Die39b] Jean Dieudonné. Sur les espaces topologiques susceptibles d’être munis d’une structure uniforme d’espace complet. Comptes rendus de l’Académie des sciences, Paris, 209:666–668, 1939.

Dieudonne:1939:EEN

[Die39c] Jean Dieudonné. Un exemple d’espace normal non susceptible d’une structure uniforme d’espace complet. Comptes rendus de l’Académie des sciences, Paris, 209:145–147, 1939.

Dienes:1946:CEM

[Die46] Z. P. Dienes. Sur la comparabilité des ensembles mesurables B par des prodédés dénombrables. Comptes rendus de l’Académie des sciences, Paris, 223:967–969, 1946.

Danielson:1942:SIPa

[DL42a] Gordon C. Danielson and Cornelius Lanczos. Some improvements in practical Fourier analysis and their application to X-ray scattering from liquids. Journal of The Franklin Institute, 233(4):365–380, April 1942. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Danielson:1942:SIPb

[DL42b] Gordon C. Danielson and Cornelius Lanczos. Some improvements in practical Fourier analysis and their application to X-ray scattering from liquids. Journal of The Franklin Institute, 233(5):435–452, May 1942. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Donoho:2017:YDS

[Don17] David Donoho. 50 years of data science. Journal of Computational and Graphical Statistics, 26(4):745–766, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See comments [?, ?].

Dolby:1969:RCS

[DRT69] J. L. Dolby, H. L. Resnikoff, and John W. Tukey. A ruly code for serials. Proceedings of the American Society for Information Science, 6(??):113–124, ???? 1969. CODEN AINPBX. ISSN 0044-7870.

Dixon:1968:ABD

[DT68] W. J. Dixon and John W. Tukey. Approximate behavior of the distribution of Winsorized t (trimming/Winsorization 2 ). Technometrics, 10(1):83–98, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266226.

Dolby:1973:ISP

[DT73a] James L. Dolby and John W. Tukey. Index to Statistics and Probability. The Statistics CumIndex, volume 1. R&D Press, Los Altos, CA, USA, 1973. ISBN 0-88274-001-6. ???? pp. LCCN Z6654.M33 T84 1973.

Dolby:1973:SC

[DT73b] James L. Dolby and John W. Tukey. The statistics CumIndex, volume 1 of The Information Access Series. The R & D Press, Los Altos, CA, 1973. ISBN 0-88274-000-8. xviii + 498 pp. LCCN Z7551 .D64. With the assistance of William E. Houchin and Karen B. Dolby.

Piza:1946:RVD

[dTP46] Affonso P. de Toledo Piza. Representative values of a distribution. Indices of dispersion. Anais Acad. Brasil. Ci., 28:209–235, 1946. ISSN 0001-3765 (print), 1678-2690 (electronic).

Dwinas:1947:ATR

[Dwi47] S. Dwinas. An application of the theory of random sampling to the theory of the integral. Revista Mat. Hisp.-Amer. (4), 7:234–238, 1947.

Efron:1979:RLB

[Efr79] B. Efron. The 1977 Rietz Lecture: Bootstrap methods: Another look at the jackknife. Annals of Statistics, 7(1):1–26, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://www.jstor.org/stable/2958830. See

[Que49, Que56, Tuk58a, Mil74].

Efron:2006:DTP

[Efr06] Bradley Efron. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):330–332, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471203. See [Mal06].

Egudin:1946:EMC

[Egu46] G. I. Egudin. On an effective method of calculation of the mathematical expectations of central sample moments. C. R. (Doklady) Acad. Sci. URSS (N.S.), 53:487–490, 1946.

Ehrenberg:1979:BRE

[Ehr79] A. S. C. Ehrenberg. Book reviews: Exploratory Data Analysis, by John W. Tukey. Applied Statistics, 28(1):79–83, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Emerson:1988:CAA

[EHTW88a] J. D. Emerson, D. C. Hoaglin, John W. Tukey, and G. Y. Wong. Comments on Adverbs and adjectives: A model-based analysis, by N. Cliff. Chance, New Directions for Statistics and Computers, 1(3):42–48, ???? 1988. CODEN ???? ISSN 0933-2480 (print), 1867-2280 (electronic).

Emerson:1988:ESA

[EHTW88b] J. D. Emerson, D. C. Hoaglin, John W. Tukey, and G. Y. Wong. Exploring some adverb adjective data. Chance, New Directions for Statistics and Computers, 1(3):42–48, ???? 1988. CODEN ???? ISSN 0933-2480 (print), 1867-2280 (electronic).

Ericksen:1989:ACP

[EKT89] Eugene P. Ericksen, Joseph B. Kadane, and John W. Tukey. Adjusting the 1980 census of population and housing. Journal of the American Statistical Association, 84(408):927–944, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290068.

Epstein:1948:SAM

[Eps48] Benjamin Epstein. Some applications of the Mellin transform in statistics. Annals of Mathematical Statistics, 19:370–379, 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Epp:1971:TUV

[ETW71] R. J. Epp, J. W. Tukey, and G. S. Watson. Testing unit vectors for correlation. Technical report, series 2 7, Princeton University, Princeton, NJ, USA, 1971. 10 pp.

Eyraud:1940:DRE

[Eyr40a] Henri Eyraud. D’une représentation des ensembles fermés. Ann. Univ. Lyon. Sect. A. (3), 3:33–37, 1940.

Eyraud:1940:SBR

[Eyr40b] Henri Eyraud. Schémas bifurqués et représentations transfinies. Ann. Univ. Lyon. Sect. A. (3), 3:25–32, 1940.

Eyraud:1943:PDC

[Eyr43] Henri Eyraud. Le problème du continu. Ann. Univ. Lyon. Sect. A. (3), 6:33–45, 1943.

Eyraud:1944:TOS

[Eyr44] Henri Eyraud. Les transfinis ordinaux des seconde et troisième classe. Ann. Univ. Lyon. Sect. A. (3), 7:5–13, 1944.

Eyraud:1945:PS

[Eyr45] Henri Eyraud. Le problème de la saturation. Ann. Univ. Lyon. Sect. A. (3), 8:47–48, 1945.

Eyraud:1947:DA

[Eyr47a] Henri Eyraud. De la divisibilité asymptotique. Comptes rendus de l’Académie des sciences, Paris, 224:169–171, 1947.

Eyraud:1947:TDC

[Eyr47b] Henri Eyraud. Le théorème du continu. Comptes rendus de l’Académie des sciences, Paris, 224:85–87, 1947.

Eyraud:1947:LTE

[Eyr47c] Henri Eyraud. Leçons sur la Théorie des Ensembles, les Nombres Transfinis et le Problème du Continu. Institut de Mathématiques, Lyon, 1947. 65 pp.

F:1971:RAS

[F.71] W. T. F. Review: Acceptance Sampling – A Symposium, by J. W. Tukey, F. Mosteller, and C. P. Winsor. Biometrics, 27(4):1109, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL http://www.jstor.org/stable/2528858.

Farah:1946:ZT

[Far46] Edison Farah. Zorn’s theorem. Bol. Soc. Mat. São Paulo, 1:19–34, 1946.

Fernholz:2003:RJW

[Fer03] Luisa Turrin Fernholz. Remembering John W. Tukey. Statistical Science, 18(3):336–340, 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1076102421.

Fernholz:2009:SIH

[Fer09] Luisa Turrin Fernholz, editor. Special issue in honor of John W. Tukey (1915–2000): contemporary data analysis: theory and methods, volume 122(1/2) of Journal of statistical planning and inference. Elsevier, Amsterdam, The Netherlands, 2009. CODEN JSPIDN. ISBN ???? ISSN 0378-3758 (print), 1873-1171 (electronic). vi + 294 pp. LCCN ????

Festinger:1946:SDB

[Fes46] Leon Festinger. The significance of difference between means without reference to the frequency distribution function. Psychometrika, 11:97–105, 1946. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Fisherkeller:1974:PIMa

[FFT74a] M. A. Fisherkeller, J. H. Friedman, and John W. Tukey. PRIM-9: An interactive multidimensional data display and analysis system. Technical Report SLAC-PUB-1408, Stanford Linear Accelerator Center, Stanford, CA, USA, 1974. ???? pp. Also in (1988) Dynamic Graphics for Statistics (W. S. Cleveland and M. E. McGill) 91–109. Wadsworth, Pacific Grove, CA.

Fisherkeller:1974:PIMb

[FFT74b] M. A. Fisherkeller, J. H. Friedman, and John W. Tukey. PRIM-9: An interactive multidimensional data display and analysis system. In A.E.C. Scientific Computer Information Exchange Meeting, May 2–3, page ???? ????, ????, 1974. Sound film, 25 minutes. Bin-88 Productions, Stanford Linear Accelerator Center. Video tape available through the ASA Video Library.

Fisherkeller:1975:PIM

[FFT75] Mary Anne Fisherkeller, Jerome H. Friedman, and John W. Tukey. PRIM-9: An interactive multi-dimensional data display and analysis system. In ACM Pacific, pages 140–145. ACM Press, New York, NY 10036, USA, 1975.

Fisherkeller:1988:PIM

[FFT88] M. A. Fisherkeller, J. H. Friedman, and J. W. Tukey. PRIM-9: An interactive multidimensional data display and analysis system. In Cleveland and McGill [CM88], pages xii + 424 + 16. ISBN 0-534-09144-X. LCCN QA276.3 .D96 1988. URL http://www.

loc.gov/catdir/enhancements/fy0744/88026017-d.html.

Fienberg:2006:WDB

[Fie06] Stephen E. Fienberg. When did Bayesian inference become “Bayesian”? Bayesian Analysis, 1(1):1–40, March 2006. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http:

//ba.stat.cmu.edu/journal/2006/vol01/issue01/fienberg.
pdf;   http://projecteuclid.org/euclid.ba/1340371071.

Finney:1946:FDD

[Fin46] D. J. Finney. The frequency distribution of deviates from means and regression lines in samples from a multivariate normal population. Annals of Mathematical Statistics, 17:344–349, 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Fisher:1950:CMS

[Fis50] Sir Ronald Aylmer Fisher. Contributions to Mathematical Statistics. Wiley publications in statistics. Wiley, New York, NY, USA, 1950. various pp. LCCN QA276 .F49. Index prepared by John Tukey.

Fisher:1954:AVV

[Fis54] Ronald Fisher. The analysis of variance with various binomial transformations. Biometrics, 10(??):130–139; discussion, 140–151, 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fisher:1955:SMS

[Fis55] Ronald Fisher. Statistical methods and scientific induction. Journal of the Royal Statistical Society. Series B (Methodological), 17: 69–78, 1955. CODEN JSTBAJ. ISSN 0035-9246.

Friede:1948:M

[FM48] G. Friede and H. Münzner. Zur Maximalkorrelation. Zeitschrift für Angewandte Mathematik und Mechanik, 28:158–160, 1948. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Fernholz:1997:CJW

[FM97] L. T. Fernholz and S. Morgenthaler. A conversation with John W. Tukey. In Brillinger et al. [BFM97], pages 26–45. ISBN 0-691-05782-6. LCCN QA276.16 .P73 1997. URL http://www.loc.

gov/catdir/description/prin021/97019695.html;   http:/
/www.loc.gov/catdir/toc/prin031/97019695.html. Two-day

symposium, Princeton on June 19–20, 1995.

Fernholz:2000:CJW

[FM00] Luisa T. Fernholz and Stephan Morgenthaler. A conversation with John W. Tukey and Elizabeth Tukey. Statistical Science, 15(1):79–94, February 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1009212675;   http://www.
jstor.org/stable/2676678.

Fernholz:2003:CJT

[FM03] Luisa Turrin Fernholz and Stephan Morgenthaler. A conversation with John Tukey. Statistical Science, 18(3):346–356, 2003. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1076102424.

Fernholz:2004:ONM

[FMT04] Luisa T. Fernholz, Stephan Morgenthaler, and John W. Tukey. An outlier nomination method based on the multihalver. Journal of Statistical Planning and Inference, 122(1–2):125–139, ???? 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). Contemporary data analysis: theory and methods.

Field:2006:HDI

[FMT06] Chris Field, Stephan Morgenthaler, and John W. Tukey. Highdimensional integration for robustness. Metron, 64(1):1–17, ???? 2006. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X

(electronic). URL https://ideas.repec.org/a/mtn/ancoec/
060101.html.

Fomin:1940:ETR

[Fom40] S. Fomin. Erweiterungen topologischer Räume. Rec. Math. [Mat. Sbornik] N.S., 8 (50):285–294, 1940.

Frechet:1946:GMC

[Fré46] M. Fréchet. A general method of constructing correlation indices. Proc. Math. Phys. Soc. Egypt, 3:13–20, 1946. ISSN 1110-0613.

Frechet:1948:ANG

[Fré48] Maurice Fréchet. Additional note on a general method of constructing correlation indices. Proc. Math. Phys. Soc. Egypt, 3: 73–74, 1948. ISSN 1110-0613.

Federer:1992:HME

[FRT92] W. T. Federer, D. S. Robson, and J. W. Tukey. Handout for meeting of experimental station statisticians “tests for non-additivity”. In Cox [Cox92], pages 19–73. ISBN 0-534-05104-9. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Finney:1948:TCW

[FS48] D. J. Finney and W. L. Stevens. A table for the calculation of working probits and weights in probit analysis. Biometrika, 35 (1/2):191–201, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332639.

Friedman:2002:JWT

[FS02] Jerome H. Friedman and Werner Stuetzle. John W. Tukey’s work on interactive graphics. Annals of Statistics, 30(6):1629–1639, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://www.jstor.org/stable/1558733.

Fuerst:1978:EPA

[FSG^{+}78] H. T. Fuerst, H. Schimmel, I. Goldstein, J. Tukey, R. W. Buechley, I. T. T. Higgins, and P. Bloomfield. Evidence for possible acute health effects of ambient air-pollution from time-series analysis — methodological questions and some new results based on New-York-City daily mortality, 1963–1976 — general discussion — session-II. Bulletin of the New York Academy of Medicine, 54(11): 1132–1136, ???? 1978. ISSN 0028-7091.

Freeman:1950:TRA

[FT50] Murray F. Freeman and John W. Tukey. Transformations related to the angular and the square root. Annals of Mathematical Statistics, 21(4):607–611, ???? 1950. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177729756. Reprinted in

[Mal90, pp. 149–155].

Friedman:1973:PPA

[FT73] Jerome H. Friedman and John W. Tukey. A projection pursuit algorithm for exploratory data analysis. Technical report, series 2 47, Princeton University, Princeton, NJ, USA, 1973. 17 pp.

Friedman:1974:PPA

[FT74] Jerome H. Friedman and John W. Tukey. A projection pursuit algorithm for exploratory data analysis. IEEE Transactions on Computers, C-23(9):881–890, 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Friedman:1980:AAD

[FTT80] J. H. Friedman, J. W. Tukey, and P. A. Tukey. Approaches to analysis of data that concentrate near intermediate-dimensional manifolds. In E. Diday, L. Lebart, J. P. Pages, and R. Tomassone, editors, Data Analysis and Informatics, Proceedings of the Second International Symposium on Data Analysis and Informatics, pages 3–13. North-Holland Publishing Co., Amsterdam, The Netherlands, 1980. ISBN 0-444-86005-3. LCCN QA278 .I56 1979.

Garwin:1969:FFT

[Gar69] Richard L. Garwin. The Fast Fourier Transform as an example of the difficulty in gaining wide use for a new technique. IEEE Transactions on Audio and Electroacoustics, AU-17(2):69–72, June 1969. CODEN ITADAS. ISSN 0018-9278 (print), 1558-2582 (electronic).

Gauss:1866:NTI

[Gau66] Carl Friedrich Gauss. Nachlass: Theoria Interpolationis methodo nova tractata. (German/Latin) [deduction: Interpolation theory by a new method]. In Carl Friedrich Gauss, Werke, volume 3, pages 265–303. Konigleichen Gesselschaft der Wissenschaften, Göttingen, Germany, 1866. LCCN ???? Posthumous publication of undated, and previously unpublished, work done about October/November 1805, according to historical evidence presented in [HJB85]. Gauss’ work predates Fourier’s work of 1807 on the representation of functions as infinite series of trigonometric functions, but due to opposition by Lagrange, Fourier did not publish it until 1822.

Gentleman:1969:CSS

[GGT69] W. M. Gentleman, J. P. Gilbert, and John W. Tukey. Chapter 4: The smear-and-sweep analysis. In The National Halothane Study, a study of the possible association between halothane anesthesia and postoperative hepatic necrosis, pages 287–315. National Institutes of Health, Bethesda, MD, USA, 1969. Chapter 4.

Gnanadesikan:1993:DEJ

[GH93] R. Gnanadesikan and D. C. Hoaglin. A discussion with Elizabeth and John Tukey. Video tape. Parts I and II. Amer. Statist. Assoc., Alexandria, VA., 1993.

Gupta:1993:SFP

[GK93] Anshul Gupta and Vipin Kumar. The scalability of FFT on parallel computers. IEEE Transactions on Parallel and Distributed Systems, 4(8):922–932, August 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Goodall:1998:CUR

[GKT98] Colin R. Goodall, Karen Kafadar, and John W. Tukey. Computing and using rural versus urban measures in statistical applications. The American Statistician, 52(2):101–111, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/goodall.html;   http://www.jstor.org/stable/2685467.

Goldberg:1946:AFP

[GL46] Henry Goldberg and Harriet Levine. Approximate formulas for the percentage points and normalization of t and χ^{2}. Annals of Mathematical Statistics, 17:216–225, 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Gilbert:1976:SSP

[GMT76] J. P. Gilbert, F. Mosteller, and John W. Tukey. Steady social progress requires quantitative evaluation to be searching. In C. C. Abt, editor, The Evaluation of Social Programs, pages 295–312. Sage, Beverly Hills, CA, USA, 1976.

Goldstine:1977:HNA

[Gol77] Herman Heine Goldstine. A history of numerical analysis from the 16th through the 19th century, volume 2 of Studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-90277-5. xiv + 348 pp. LCCN QA297 .G64.

Good:1958:IAP

[Goo58] I. J. Good. The interaction algorithm and practical Fourier analysis. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):361–372, 1958. CODEN JSTBAJ. ISSN 0035-9246. Addendum in same journal, 22(2) 429–430, July 1960.

Gentleman:1966:FFT

[GS66] W. Morven Gentleman and Gordon Sande. Fast Fourier Transforms—for fun and profit. In Fall Joint Computer Conference, volume 29 of AFIPS Conference Proceedings, pages 563–578. Spartan Books, New York, NY, USA, 1966.

Gulliksen:1958:RLC

[GT58] Harold O. Gulliksen and John W. Tukey. Reliability for the law of comparative judgment. Psychometrika, 23(??):95–110, 1958. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Green:1960:CAV

[GT60] Bert F. Green, Jr. and John W. Tukey. Complex analyses of variance: general problems. Psychometrika, 25(??):127–152, 1960. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Gentleman:1969:ASW

[GT69a] W. M. Gentleman and John W. Tukey. Appendix 4: Should we superstandardize agent comparisons? — general considerations and apparent conclusions. In The National Halothane Study, a study of the possible association between halothane anesthesia and postoperative hepatic necrosis, pages 372–380. National Institutes of Health, Bethesda, MD, USA, 1969.

Gentleman:1969:ADC

[GT69b] W. M. Gentleman and John W. Tukey. Appendix 5: Development of certain formulas. In The National Halothane Study, a study of the possible association between halothane anesthesia and postoperative hepatic necrosis, pages 381–386. National Institutes of Health, Bethesda, MD, USA, 1969.

Gentleman:1969:AAV

[GT69c] W. M. Gentleman and John W. Tukey. Appendix 6: Assessment of various structural regression coefficients. In The National Halothane Study, a study of the possible association between halothane anesthesia and postoperative hepatic necrosis, pages 387– 391. National Institutes of Health, Bethesda, MD, USA, 1969.

Gross:1973:EPR

[GT73] I. M. Gross and John W. Tukey. The estimators of the Princeton robustness study. Technical report, series 2 38, Princeton University, Princeton, NJ, USA, 1973. 39 pp.

Guttman:1945:BAT

[Gut45] Louis Guttman. A basis for analyzing test-retest reliability. Psychometrika, 10:255–282, 1945. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Guttman:1948:IK

[Gut48] Louis Guttman. An inequality for kurtosis. Annals of Mathematical Statistics, 19:277–278, 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Gildemeister:1943:ZKK

[GvdW43] M. Gildemeister and B. L. van der Waerden. Die Zulässigkeit des χ^{2}-Kriteriums für kleine Versuchszahlen. Ber. Verh. Sächs. Akad. Wiss. Leipzig. Math.-Nat. Kl., 95:145–150 (1944), 1943.

Gao:2017:ERH

[GWD17] Zhen Gao, Xiao Wen, and Wai Sun Don. Enhanced robustness of the hybrid compact–WENO finite difference scheme for hyperbolic conservation laws with multi-resolution analysis and Tukey’s boxplot method. Journal of Scientific Computing, 73(2– 3):736–752, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0465-0;   https://link.
springer.com/content/pdf/10.1007/s10915-017-0465-0.pdf.

Hafkin:2000:PER

[Haf00] Jessica Hafkin. Professor Emeritus remembered for lively humor and passion. The Daily Princetonian, ??(??):??, September 19,

2000. URL http://www.dailyprincetonian.com/2000/09/19/
1202/.

Halmos:1946:TUE

[Hal46] Paul R. Halmos. The theory of unbiased estimation. Annals of Mathematical Statistics, 17:34–43, 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Hamilton:1946:REF

[Ham46] Hugh J. Hamilton. Roots of equations by functional iteration. Duke Mathematical Journal, 13:113–121, 1946. CODEN DUM-JAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Hartley:1946:ASC

[Har46] H. O. Hartley. The application of some commercial calculating machines to certain statistical calculations. Suppl. J. Roy. Statist. Soc., 8:154–173, discussion, 173–183, 1946.

Hartigan:2003:MJT

[Har03] J. A. Hartigan. A memory of John Tukey as a teacher. Statistical Science, 18(3):341, 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1076102422.

Hayter:1984:PCT

[Hay84] Anthony J. Hayter. Proof of the conjecture that the Tukey– Kramer multiple comparisons procedure is conservative. Annals of Statistics, 12(1):61–75, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346392;   http://www.
jstor.org/stable/2241034.

Hooper:1978:PD

[HBK^{+}78] R. L. Hooper, C. A. Bennett, H. J. C. Kouts, F. C. Leone, and John W. Tukey. Panel discussion. In A. Gardiner and T. Truett, editors, Proceedings of the 1977 DOE Statistical Symposium, pages 175–183. National Technical Information Service, Washington, DC, USA, 1978.

Healy:1983:RUR

[Hea83] M. J. R. Healy. Review: Understanding Robust and Exploratory Data Analysis, by D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Biometrics, 39(4):1126–1127, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://

www.jstor.org/stable/2531362.

Healy:1984:BRBb

[Hea84] M. J. R. Healy. Book review: Graphical Methods for Data Analysis, by J. M. Chambers; W. S. Cleveland; B. Kleiner; P. A. Tukey. Journal of the Royal Statistical Society. Series A (General), 147 (3):513, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2981587.

Heckler:2001:BRB

[Hec01] Charles E. Heckler. Book review: Graphical Analysis of Multiresponse Data, Illustrated with a Plant Breeding Trial, Interdisciplinary Statistics, by K. E. Basford and J. W. Tukey. Technometrics, 43(1):97–98, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270862.

Hinkley:1973:BRB

[Hin73a] David V. Hinkley. Book review: Robust Estimates of Location: Survey and Advances, by D. F. Andrews; P. J. Bickel; F. R. Hampel; P. J. Huber; W. H. Rogers; J. W. Tukey. Journal of the Royal Statistical Society. Series A (General), 136(3):458–459, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2345006.

Hinkley:1973:RRE

[Hin73b] David V. Hinkley. Review: Robust Estimates of Location: Survey and Advances, by D. F. Andrews, P. J. Bickel, F. R. Hampel, P. J. Huber, W. H. Rogers, and J. W. Tukey. Journal of the Royal Statistical Society. Series A (General), 136(3):458–459, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2345006.

Hoaglin:1981:SSP

[HIT81] David C. Hoaglin, Boris Iglewicz, and John W. Tukey. Smallsample performance of a resistant rule for outlier detection. In ????, editor, 1980 Proceedings of the Statistical Computing Section, pages 148–152. American Statistical Association, Washington, DC, USA, 1981. ISSN 0149-9963. LCCN QA276.4 .A43a.

Hoaglin:1986:PSR

[HIT86] David C. Hoaglin, Boris Iglewicz, and John W. Tukey. Performance of some resistant rules for outlier labeling. Journal of the American Statistical Association, 81(396):991–999, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289073. See com-

ment [Sim87].

Hoaglin:1987:HIT

[HIT87] D. C. Hoaglin, B. Iglewicz, and J. W. Tukey. Hoaglin, Iglewicz, and Tukey — reply. Journal of the American Statistical Association, 82 (398):704, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274x (electronic).

Heideman:1984:GHF

[HJB84] M. Heideman, D. Johnson, and C. Burrus. Gauss and the history of the Fast Fourier Transform. IEEE ASSP magazine: a publication of the IEEE Acoustics, Speech, and Signal Processing Society, 1 (4):14–21, October 1984. CODEN ???? ISSN 0740-7467 (print), 1558-1284 (electronic).

Heideman:1985:GHF

[HJB85] Michael T. Heideman, Don H. Johnson, and C. Sidney Burrus. Gauss and the history of the Fast Fourier Transform. Archive for History of Exact Sciences, 34(3):265–277, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=34&issue=3&spage=265.

See [CT65, CT93].

Hartley:1947:NSP

[HK47] H. O. Hartley and S. H. Khamis. A numerical solution of the problem of moments. Biometrika, 34(3/4):340–351, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332445.

Harris:1950:TDE

[HMT50] T. E. Harris, P. Meier, and J. W. Tukey. Timing of the distribution of events between observations. Human biology, 22(??):249–270, ???? 1950. CODEN HUBIAA. ISSN 0018-7143 (print), 1534-6617 (electronic).

Tukey:1983:IMR

[HMT83a] D. C. Hoaglin, F. Mosteller, and John W. Tukey. Introduction to more refined estimators. In D. C. Hoaglin, F. Mosteller, and J. W. Tukey, editors, Understanding Robust and Exploratory Data Analysis, pages 283–296. Wiley, New York, NY, USA, 1983.

Hoaglin:1983:IMR

[HMT83b] David C. Hoaglin, Frederick Mosteller, and John W. Tukey. Introduction to more refined estimators. In Understanding Robust and Exploratory Data Analysis [HMT83c], pages 283–296. ISBN 0-471-09777-2. LCCN QA276 .U5 1983. Revised and updated reprint in 2000.

Hoaglin:1983:URE

[HMT83c] David C. Hoaglin, Frederick Mosteller, and John W. Tukey, editors. Understanding Robust and Exploratory Data Analysis. Wiley, New York, NY, USA, 1983. ISBN 0-471-09777-2. xvi + 447 pp. LCCN QA276 .U5 1983. Revised and updated reprint in 2000.

Hoaglin:1985:EDT

[HMT85] David C. Hoaglin, Frederick Mosteller, and John W. Tukey, editors. Exploring Data Tables, Trends, and Shapes. Wiley, New York, NY, USA, 1985. ISBN 0-470-04005-X. xxii + 527 pp. LCCN ????

Hoaglin:1991:FEA

[HMT91] David C. Hoaglin, Frederick Mosteller, and John W. Tukey, editors. Fundamentals of Exploratory Analysis of Variance. Wiley, New York, NY, USA, 1991. ISBN 0-471-52735-1. xvii + 430 pp. LCCN QA279 .F86 1991.

Hoaglin:2000:URE

[HMT00] David C. (David Caster) Hoaglin, Frederick Mosteller, and John Wilder Tukey, editors. Understanding Robust and Exploratory Data Analysis. Wiley classics library. Wiley, New York, NY, USA, 2000. ISBN 0-471-38491-7 (paperback). xx + 447

pp. LCCN QA276 .U5 2000. URL http://www.loc.gov/catdir/
description/wiley0310/00028322.html;   http://www.loc.
gov/catdir/toc/onix06/00028322.html.

Hoaglin:2006:EDT

[HMT06] David C. Hoaglin, Frederick Mosteller, and John W. Tukey, editors. Exploring data tables, trends, and shapes. Wiley Series in Probability and Statistics. Wiley, New York, NY, USA, 2006. ISBN 0-470-04005-X, 0-471-09776-4, 1-118-15069-4 (e-book), 1-118-15070-8 (e-book). xii + 527 pp. LCCN QA276 .E97 2006. EUR 75.00; UK £52.95.

Hastings:1947:LMS

[HMTW47] Cecil Hastings, Jr., Frederick Mosteller, John W. Tukey, and Charles P. Winsor. Low moments for small samples: a comparative study of order statistics. Annals of Mathematical Statistics, 18(3):413–426, ???? 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730388.

Hoare:1961:AP

[Hoa61a] C. A. R. Hoare. Algorithm 63: Partition. Communications of the ACM, 4(7):321, July 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoare:1961:AQ

[Hoa61b] C. A. R. Hoare. Algorithm 64: Quicksort. Communications of the ACM, 4(7):321, July 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoare:1961:AF

[Hoa61c] C. A. R. Hoare. Algorithm 65: Find. Communications of the ACM, 4(7):321–322, July 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoaglin:1974:RBE

[Hoa74] David C. Hoaglin. Review: Robust Estimates of Location: Survey and Advances, by D. F. Andrews, P. J. Bickel, F. R. Hampel, P. J. Huber, W. H. Rogers, and J. W. Tukey. Journal of the American Statistical Association, 69(345):272–274, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285544.

Hoaglin:2002:PJW

[Hoa02] David C. Hoaglin. Professor John W. Tukey, 1915–2000. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2): 279–280, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-

9884 (electronic). URL http://www.jstor.org/stable/3650326.

Hoaglin:2003:JWT

[Hoa03] David C. Hoaglin. John W. Tukey and data analysis. Statistical Science, 18(3):311–318, 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1076102418.

Hodges:1958:RKM

[Hod58] J. L. Hodges, Jr. Review: Keeping Moment-Like Sampling Computations Simple. Mathematical Tables and Other Aids to Computation, 12(64):300, October 1958. CODEN MTTCAS. ISSN

0891-6837. URL http://www.jstor.org/stable/2002407.

Holder:1979:BRD

[Hol79] R. L. Holder. Book reviews: Data Analysis and Regression, by Frederick Mosteller and John W. Tukey. Applied Statistics, 28 (2):177–178, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hsu:1994:GRT

[HP94] Jason C. Hsu and Mario Peruggia. Graphical representations of Tukey’s multiple comparison method. Journal of Computational and Graphical Statistics, 3(2):143–161, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10618600.1994.10474636.

Hassairi:2008:TDC

[HR08] Abdelhamid Hassairi and Ons Regaieg. On the Tukey depth of a continuous probability distribution. Statistics & Probability Letters, 78(15):2308–2313, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715208001259.

Harvey:2010:PTF

[HR10] David Harvey and Daniel S. Roche. An in-place truncated Fourier transform and applications to polynomial multiplication. In Watt [Wat10], pages 325–329. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Harris:1949:DLS

[HT49] Theodore E. Harris and John W. Tukey. Development of largesample measure of location and scale which are relatively insensitive to contamination. SRG Memorandum report 31, Princeton University, Princeton, NJ, USA, 1949. 29 pp.

Haubrich:1966:SAG

[HT66] R. A. Haubrich and John W. Tukey. Spectrum analysis of geophysical data. In Proceedings of the IBM Scientific Computing on Symposium Environmental Sciences, pages 115–128. IBM Corporation, San Jose, CA, USA, 1966.

Hoaglin:1985:CSD

[HT85] David C. Hoaglin and John W. Tukey. Checking the shape of discrete distributions. In D. C. Hoaglin, F. Mosteller, and J. W. Tukey, editors, Exploring Data Tables, Trends, and Shapes, pages 345–416. Wiley, New York, NY, USA, 1985.

Hoang:1989:PSW

[HT89] T. Hoang and J. W. Tukey. Procedures for separations within batches of values, I. The orderly tool kit and some heuristics. Technical Report 293, Department of Statistics, Princeton University, Princeton, NJ, USA, 1989.

Hoaglin:1991:QQC

[HT91] D. C. Hoaglin and John W. Tukey. Qualitative and quantitative confidence. In D. C. Hoaglin, F. Mosteller, and J. W. Tukey, editors, Fundamentals of Exploratory Analysis of Variance, pages 336–364. Wiley, New York, NY, USA, 1991.

Hansen:1992:TMP

[HT92] K. M. Hansen and John W. Tukey. Tuning a major part of a clustering algorithm. International Statistical Review = Revue Internationale de Statistique, 60(1):21–43, ???? 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Huang:1971:HFF

[Hua71] T. S. Huang. How the Fast Fourier Transform got its name. Computer, 4(3):15, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huber:1964:REL

[Hub64] Peter J. Huber. Robust estimation of a location parameter. Annals of Mathematical Statistics, 35(1):73–101, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703732.

Huber:2002:JWT

[Hub02] Peter J. Huber. John W. Tukey’s contributions to robust statistics. Annals of Statistics, 30(6):1640–1648, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1043351251;   http://www.jstor.org/stable/1558734.

Huber:2006:DTP

[Hub06] Peter J. Huber. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):332–334, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471204. See [Mal06].

Iles:1993:BRB

[Ile93a] Terence Iles. Book review: Fundamentals of Explaratory Analysis of Variance, by D. C. Hoaglin; F. Mosteller; J. W. Tukey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):323–324, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.jstor.

org/stable/2982745.

Iles:1993:RFE

[Ile93b] Terence Iles. Review: Fundamentals of Explaratory Analysis of Variance, D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):323–324, ???? 1993. CODEN ???? ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.jstor.org/

stable/2982745.

J:1977:BRBb

[J.77a] J. E. J. Book review: Index to Statistics and Probability: Citation Index by John W. Tukey. Technometrics, 19(1):103, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268267.

J:1977:BRBd

[J.77b] J. E. J. Book review: Index to Statistics and Probability: Locations and Authors by I. C. Ross; J. W. Tukey. Technometrics, 19(1):104, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268269.

J:1977:BRBc

[J.77c] J. E. J. Book review: Index to Statistics and Probability: Permuted Titles by I. C. Ross; J. W. Tukey. Technometrics, 19(1):104, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268268.

Jackson:1965:CPK

[Jac65] J. Edward Jackson. Comments on paper by Kurtz, Link, Tukey and Wallace. Technometrics, 7(2):163–165, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266667.

Johnson:2020:HDS

[JDC20] Tim Johnson, Christopher T. Dawes, and Dalton Conley. How does a statistician raise an Army? The time when John W. Tukey, a team of luminaries, and a statistics graduate student repaired the Vietnam Selective Service Lotteries. The American Statistician, 74(2):190–196, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2019.1677267.

Jeffreys:1946:IFP

[Jef46] Harold Jeffreys. An invariant form for the prior probability in estimation problems. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 186:453–461, 1946. CODEN PRLAAZ. ISSN 0080-4630.

Johnson:1973:EMT

[JK73] N. L. Johnson and S. Kotz. Extended and multivariate Tukey Lambda distributions. Biometrika, 60(3):655–661, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335017.

Johnsson:1992:CTF

[JK92] S. L. Johnsson and R. L. Krawitz. Cooley–Tukey FFT on the Connection Machine. Parallel Computing, 18(11):1201–1221, November 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jones:2001:HTM

[JLT01] L. V. Jones, C. Lewis, and J. W. Tukey. Hypothesis tests, multiplicity of. In Smelser and Baltes [SB01], pages 7127–7133. ISBN 0-08-043076-7 (26-volume set). LCCN H41 .I58 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0612/2001044791-d.
html.

Jones:1948:ELM

[Jon48] Howard L. Jones. Exact lower moments of order statistics in small samples from a normal distribution. Annals of Mathematical Statistics, 19:270–273, 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Jones:1986:CWJa

[Jon86a] Lyle V. Jones, editor. The collected works of John W. Tukey. Volume III. Philosophy and principles of data analysis: 1949– 1964. Wadsworth & Brooks/Cole Statistics/Probability Series. Wadsworth & Brooks/Cole Advanced Books & Software, Monterey, CA, USA, 1986. ISBN 0-534-03305-9. lxviii + 516 + 53 pp. With a biography of Tukey by Frederick Mosteller.

Jones:1986:CWJb

[Jon86b] Lyle V. Jones, editor. The collected works of John W. Tukey. Volume IV. Philosophy and principles of data analysis: 1965– 1986. Wadsworth & Brooks/Cole Statistics/Probability Series. Wadsworth & Brooks/Cole Advanced Books & Software, Monterey, CA, USA, 1986. ISBN 0-534-05101-4. i–lxviii + 517–1016 + 1–53 pp. With a biography of Tukey by Frederick Mosteller.

Jenkins:1961:CDM

[JP61] G. M. Jenkins and Emanuel Parzen. Comments on the discussions of Messrs. Tukey and Goodman. Technometrics, 3(2):229–234, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266114.

Joiner:1971:SPR

[JR71] Brian L. Joiner and Joan R. Rosenblatt. Some properties of the range in samples from Tukey’s symmetric lambda distributions. Journal of the American Statistical Association, 66(334):394–399, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283943.

Johnson:1987:GEA

[JT87] E. G. Johnson and John W. Tukey. Graphical exploratory analysis of variance illustrated on a splitting of the Johnson and Tsao data. In C. Mallows, editor, Design, Data, and Analysis by Some Friends of Cuthbert Daniel, pages 171–244. Wiley, New York, NY, USA, 1987.

Jones:2000:SFS

[JT00] L. V. Jones and John W. Tukey. A sensible formulation of the significance test. Psychological Methods, 5(4):411–414, ???? 2000. CODEN ???? ISSN 1082-989X (print), 1939-1463 (electronic).

Jones:2001:SFS

[JT01] L. V. Jones and J. W. Tukey. A sensible formulation of the significance test (vol 5, pg 411, 2000). Psychological Methods, 6(1):17, ???? 2001. ISSN 1082-989x (print), 1939-1463 (electronic).

Jurgensen:1947:TDP

[Jur47] C. E. Jurgensen. Table for determining phi coefficients. Psychometrika, 12:17–29, 1947. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Johnstone:1985:RLR

[JV85] Iain M. Johnstone and Paul F. Velleman. The resistant line and related regression methods. With a comment by John W. Tukey. Journal of the American Statistical Association, 80(392):1041– 1054, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288572.

Kafadar:2001:MJW

[Kaf01] Karen Kafadar. In memoriam: John Wilder Tukey June 16, 1915–July 26, 2000. Technometrics, 43(3):251–255, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.allbusiness.com/technology/
808081-1.html;   http://www.jstor.org/stable/1271211. Spe-

cial Tukey Memorial issue.

Kafadar:2003:JTR

[Kaf03] Karen Kafadar. John Tukey and robustness. Statistical Science, 18(3):319–331, 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1076102419.

Kantorovitch:1939:TSR

[Kan39] L. V. Kantorovitch. On the theory of Stieltjes–Riemann integrals. Leningrad State Univ. Annals [Uchenye Zapiski] Math. Ser., 6: 52–68, 1939.

Kanji:1993:RFE

[Kan93] G. K. Kanji. Review: Fundamentals of Exploratory Analysis of Variance., by D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Biometrics, 49(3):958, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://www.

jstor.org/stable/2532225.

Katetov:1940:ABR

[Kat40] Miroslav Katětov. Über H-abgeschlossene und bikompakte Räume. Časopis Pěst. Mat. Fys., 69:36–49, 1940. ISSN 0528-2195.

Keldych:1940:DDD

[Kel40a] Ludmila Keldych. Démonstration directe du théorème sur l’appartenance d’un élément canonique E_{α} à la classe α et exemples arithmétiques d’ensembles mesurables B de classes supérieures. C. R. (Doklady) Acad. Sci. URSS (N.S.), 28:675–677, 1940.

Keldych:1940:EHM

[Kel40b] Ludmila Keldych. Sur les ensembles homogènes mesurables B. C. R. (Doklady) Acad. Sci. URSS (N.S.), 26:523–525, 1940.

Keldych:1941:SEM

[Kel41] Ludmila Keldych. Sur la structure des ensembles mesurables B. C. R. (Doklady) Acad. Sci. URSS (N.S.), 31:651–653, 1941.

Kemp:1987:RED

[Kem87] A. W. Kemp. Review: Exploring Data Tables, Trends, and Shapes, by D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Biometrics, 43 (1):252, March 1987. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL http://www.jstor.org/stable/
2531968.

Kemp:2001:BRBj

[Kem01] Freda Kemp. Book review: Applied Statistical Decision Theory, by H. Raiffa; R. Schlaifer; Understanding Robust and Exploratory Data Analysis by D. C. Hoaglin; F. Mosteller; J. W. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 352–353, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-

9884 (electronic). URL http://www.jstor.org/stable/2680952.

Kafadar:1996:URT

[KFGT96] Karen Kafadar, Laurence S. Freedman, Colin R. Goodall, and John W. Tukey. Urbanicity-related trends in lung cancer mortality in U.S. counties: White females and white males, 1970–1987. International Journal of Epidemiology, 25(5):918–932, ???? 1996. CODEN IJEPBF. ISSN 0300-5771 (print), 1464-3685 (electronic).

Kruskal:1960:DPM

[KFT^{+}60] William Kruskal, Thomas S. Ferguson, John W. Tukey, E. J. Gumbel, and F. J. Anscombe. Discussion of the papers of Messrs. Anscombe and Daniel. Technometrics, 2(2):157–166, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266542.

Kinsey:1955:CMT

[KHS^{+}55] Alfred C. Kinsey, Herbert Hyman, Paul B. Sheatsley, A. H. Hobbs, R. D. Lambert, Nicholas Pastore, Jacob Goldstein, Lewis M. Terman, Paul Wallin, W. Allen Wallis, William G. Cochran, Frederick Mosteller, and John W. Tukey. The Cochran–Mosteller–Tukey Report on the Kinsey Study: A Symposium. Journal of the American Statistical Association, 50(271):811–829, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281167.

Kingston:1945:TIE

[Kin45] Jorge Kingston. A Teoria da Indução Estatística. Instituto Brasileiro de Geografia e Estatística, Rio de Janeiro, 1945. 121 pp.

Kotz:1997:BSV

[KJ97] Samuel Kotz and Norman Lloyd Johnson, editors. Breakthroughs in Statistics: Volume 3. Springer series in statistics. Perspectives in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94988-7 (hardcover), 0-387-94989-5 (softcover). xxv + 559 pp.

LCCN QA276 .B68465 1997. URL http://www.loc.gov/catdir/
enhancements/fy0815/93003854-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/93003854-t.html.

Kurtz:1965:ARA

[KLTW65a] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Authors’ reply to Anscombe’s comments. Technometrics, 7(2):169, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266669.

Kurtz:1965:SCMa

[KLTW65b] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Short-cut multiple comparisons for balanced single and double classifications: Part 1, Results. Technometrics, 7(2):95–161, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266666. Reprinted in [Cox92,

pp. 353–430]. See also [Ans65, KLTW65a].

Kurtz:1965:SCMb

[KLTW65c] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Shortcut multiple comparisons for balanced single and double classifications: Part 2. Derivations and approximations. Biometrika, 52 (3/4):485–498, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2333700. Reprinted in [Cox92, pp. 431–452]. See also [Ans65, KLTW65a].

Kurtz:1966:CRC

[KLTW66] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Correlation of ranges of correlated deviates. Biometrika, 53(1/2):191– 197, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334064.

With editorial note by Egon S. Pearson.

Kadane:1999:YAP

[KMT99] Joseph B. Kadane, Michael M. Meyer, and John W. Tukey. Yule’s association paradox and ignored stratum heterogeneity in capturerecapture studies. Journal of the American Statistical Association, 94(447):855–859, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670000.

Knoll:1943:NEG

[Kno43] F. Knoll. Über Näherungsverfahren bei empirisch gegebenen Verteilungsfunktionen und damit verbundene Korrekturformeln. Deutsche Math., 7:187–194, 1943.

Kondo:1940:RPE

[Kon40] Motokiti Kondô. Sur la représentation paramétrique des ensembles. J. Fac. Sci. Hokkaido Imp. Univ. Ser. I., 8:173–220, 1940.

Koslova:1940:EPA

[Kos40] Z. Koslova. Sur les ensembles plans analytiques ou mesurables B. Bull. Acad. Sci. URSS. Sér. Math. [Izvestia Akad. Nauk SSSR], 4: 479–500, 1940.

Krasner:1939:TES

[Kra39] M. Krasner. Un type d’ensembles semi-ordonnés et ses rapports avec une hypothèse de M. A. Weil. Bulletin de la Société Mathématique de France, 67:162–176, 1939. CODEN BSMFAA. ISSN 0037-9484 (print), 2102-622x (electronic).

Kreweras:1946:ETR

[Kre46] Germain Kreweras. Extension d’un théorème sur les répartitions en classes. Comptes rendus de l’Académie des sciences, Paris, 222: 431–432, 1946.

Krumbein:1956:MAM

[KT56] W. C. Krumbein and John W. Tukey. Multivariate analysis of mineralogic, lithologic and chemical composition of rock bodies. Journal of Sedimentary Petrology, 26(??):322–337, ???? 1956. CODEN JSEPAK. ISSN 0022-4472.

Kafadar:1988:BT

[KT88] Karen Kafadar and John W. Tukey. A bidec t table. Journal of the American Statistical Association, 83(402):532–539, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288874.

Kafadar:1993:UCD

[KT93] Karen Kafadar and John W. Tukey. U.S. cancer death rates: a simple adjustment for urbanization. International Statistical Review = Revue Internationale de Statistique, 61(2):257–281, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403628.

Kurepa:1940:PFE

[Kur40a] G. Kurepa. Une propriété des familles d’ensembles bien ordonnés linéaires. Studia Mathematica, 9:23–42, 1940. CODEN SMATAZ. ISSN 0039-3223 (print), 1730-6337 (electronic).

Kurepa:1940:TME

[Kur40b] Georges Kurepa. Transformations monotones des ensembles partiellement ordonnés. Revista Ci., Lima, 42:827–846, 1940. ISSN 0034-7760.

Kurepa:1941:TME

[Kur41] Georges Kurepa. Transformations monotones des ensembles partiellement ordonnés. Revista Ci., Lima, 43:483–500, 1941. ISSN 0034-7760.

Lagrange:1759:RNP

[Lag59] J. L. Lagrange. Recherches sur la nature et la propagation du son. (French) [Research on the nature and propagation of sound]. Miscellanea Taurinensia (Mélanges de Turin), 1(1–10):1–112, 1759.

Lagrange:1762:SDP

[Lag65] J. L. Lagrange. Solutions de différents problèmes de calcul intégral. (French) [Solution of different problems in integral calculus]. Miscellanea Taurinensia (Mélanges de Turin), 3(??):??, 1762–1765.

Langer:1959:NAP

[Lan59] R. E. Langer, editor. On numerical approximation. Proceedings of a Symposium, Madison, April 21–23, 1958. The University of Wisconsin Press, Madison, WI, USA, 1959. x + 462 pp. LCCN QA3 .U45 no. 1. Publication no. 1 of the Mathematics Research Center, U.S. Army, the University of Wisconsin.

Landwehr:2006:DTP

[Lan06] James M. Landwehr. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):334–336, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471205. See [Mal06].

Lawrence:1975:BRB

[Law75] Clive J. Lawrence. Book review: Robust Estimates of Location: Survey and Advances by D. F. Andrews; P. J. Bickel; F. R. Hampel; P. J. Huber; W. H. Rogers; J. W. Tukey. Technometrics, 17(1):138–139, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268015.

Lanczos:1939:ATI

[LD39] C. Lanczos and G. C. Danielson. Application of trigonometric interpolation to X-ray analysis. Physical Review, 55(2):242, January 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). Abstract of talk.

Lehmann:2008:RSC

[Leh08] Erich Lehmann. Reminiscences of a Statistician: the Company I Kept. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-71596-7 (paperback), 0-387-71597-5 (e-book). 318 pp. LCCN QA276.156 .L44 2008. URL

http://www.springerlink.com/content/978-0-387-71597-1.

Leonhardt:2000:JTS

[Leo00] David Leonhardt. John Tukey, 85, statistician; coined the word ‘software’. New York Times, ??(??):??, July 28, 2000. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://www.nytimes.com/2000/07/28/us/john-
tukey-85-statistician-coined-the-word-software.html.

Leverett:1947:TMD

[Lev47] Hollis M. Leverett. Table of mean deviates for various portions of the unit normal distribution. Psychometrika, 12:141–152, 1947. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Ley:1977:SAR

[Ley77] J. P. Ley. Statistical algorithms: Remark A.S. R23: a remark on Algorithm AS 61: “Six-Line Plots”. Applied Statistics, 26(3): 368, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876

(electronic). URL http://lib.stat.cmu.edu/apstat/61. See

[AT73a].

Liapounoff:1939:SMP

[Lia39a] A. Liapounoff. Séparabilité multiple pour le cas de l’opération (A). Bull. Acad. Sci. URSS. Sér. Math. [Izvestia Akad. Nauk SSSR], 1939:539–552, 1939.

Liapounoff:1939:PO

[Lia39b] A. Liapounoff. Sur une propriété des δs-opérations. Bull. Acad. Sci. URSS. Sér. Math. [Izvestia Akad. Nauk SSSR], 1939:407–412, 1939.

Lide:2001:CEM

[Lid01] D. R. Lide, editor. A Century of Excellence in Measurements, Standards, and Technology: A Chronicle of Selected NBS/NIST Publications, 1901–2000, volume 958. National Technical Information Service, Washington, DC, USA, 2001. ix + 386

pp. URL https://nvlpubs.nist.gov/nistpubs/sp958-lide/

cntsp958old.htm. NIST Special Publication.

Laderman:1954:EIP

[LLT54] J. Laderman, S. B. Littauer, and John W. Tukey. Errata: The inventory problem. Journal of the American Statistical Association, 49(268):906, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281550.

Lorenzo:2018:ASB

[Lor18] Mark Jones Lorenzo. Adventures of a Statistician: The Biography of John W. Tukey. SE Books, 5307 West Tyson Street, Philadelphia, PA 19107, USA, 2018. ISBN 1-72201-358-3 (paperback). 340 pp. LCCN QA276.157 .T854 2018.

Lovera:1942:UDC

[Lov42] G. Lovera. Un’applicazione del coefficiente di correlazione alle medie statistiche. Atti Accad. Sci. Torino. Cl. Sci. Fis. Mat. Nat., 77:341–346, 1942. ISSN 0001-4419.

Lovera:1945:MAC

[Lov45] G. Lovera. Metodo abbreviato di calcolo delle caratteristiche di una correlazione multipla. Atti Accad. Sci. Torino. Cl. Sci. Fis. Mat. Nat., 80:194–198, 1945. ISSN 0001-4419.

Liu:2007:CDI

[LSZ07] Regina Liu, William Strawderman, and Cun-Hui Zhang, editors. Complex Datasets and Inverse Problems: Tomography, Networks, and Beyond, volume 54 of Institute of mathematical statistics lecture notes — monograph series. Institute of Mathematical Statistics, Beachwood, OH, USA, 2007. ISBN 0-940600-70-6. ???? pp. LCCN ????

Lefschetz:1946:PUB

[LT46] Solomon Lefschetz and John W. Tukey, editors. The Princeton University Bicentennial Conference on the Problems of Mathematics, Fine Hall, December 17–19, 1946. Princeton University Press, Princeton, NJ, USA, 1946. LCCN ????

Luce:1964:SCM

[LT64] R. Duncan Luce and John W. Tukey. Simultaneous conjoint measurement: a new type of fundamental measurement. Journal of Mathematical Psychology, 1(1):1–27, January 1964. CODEN JMT-PAJ. ISSN 0022-2496 (print), 1096-0880 (electronic).

Lewis:2001:IMC

[LT01] C. Lewis and J. W. Tukey. Improved multiple comparison procedures for controlling the false discovery rate. Unpublished manuscript (cited in [BB02, page 1593])., 2001.

Lusternick:1940:IEL

[Lus40a] L. Lusternick. Intersections dans les espaces localement linéaires. C. R. (Doklady) Acad. Sci. URSS (N.S.), 27:771–774, 1940.

Lusternick:1940:STE

[Lus40b] L. Lusternick. Structure topologique d’un espace fonctionnel. C. R. (Doklady) Acad. Sci. URSS (N.S.), 27:775–777, 1940.

Maceda:1947:CF

[Mac47] E. Cansado Maceda. Cumulants of Fisher’s z. Revista Mat. Hisp.- Amer. (4), 7:87–89, 1947.

Maceda:1948:CGP

[Mac48] E. Cansado Maceda. On the compound and generalized Poisson distributions. Annals of Mathematical Statistics, 19:414–416, 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

MacKenzie:1991:IAL

[Mac91] Donald MacKenzie. The influence of the Los Alamos and Livermore National Laboratories on the development of supercomputing. Annals of the History of Computing, 13(2): 179–201, April/June 1991. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1991/pdf/
a2179.pdf;   http://www.computer.org/annals/an1991/a2179abs.
htm.

Mathews:1976:CSP

[MACT76] M. V. Mathews, B. Atal, J. J. Chang, and J. W. Tukey. Computersorting procedure for inverting functions. Journal of the Acoustical Society of America, 60(??):S77, ???? 1976. CODEN JASMAN. ISSN 0001-4966.

Mallows:1979:STR

[Mal79] C. L. Mallows. Some theoretical results on Tukey’s 3R smoother. Lecture Notes in Mathematics, 757:77–90, 1979. CODEN LN-MAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098491/.

Mallows:1990:CWJ

[Mal90] Colin L. Mallows, editor. The collected works of John W. Tukey. Volume VI. More mathematical: 1938–1984. Wadsworth & Brooks/Cole Statistics/Probability Series. Wadsworth & Brooks/ Cole Advanced Books & Software, Monterey, CA, USA, 1990. ISBN 0-534-05103-0. lxxii + 644 + i17 pp. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Mallows:2003:JTB

[Mal03] Colin Mallows. John Tukey at Bell Labs. Statistical Science, 18(3): 332–335, 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.ss/
1076102420.

Mallows:2006:TPA

[Mal06] Colin Mallows. Tukey’s paper after 40 years. Technometrics, 48 (3):319–325, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/25471200. See discussion [Bri06, Buj06, Efr06, Hub06, Lan06].

Masuyama:1941:CCB

[Mas41a] Motosaburô Masuyama. Correlation coefficient between two sets of complex vectors. Proc. Phys.-Math. Soc. Japan (3), 23:918–924, 1941.

Masuyama:1941:STA

[Mas41b] Motosaburô Masuyama. On the significance test of the additive correlation coefficient. Proc. Phys.-Math. Soc. Japan (3), 23:1016– 1019, 1941.

Masse:2004:ATD

[Mas04] Jean-Claude Massé. Asymptotics for the Tukey depth process, with an application to a multivariate trimmed mean. Bernoulli, 10(3):397–419, June 2004. ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.bj/
1089206404.

Maximoff:1939:EME

[Max39] Isaie Maximoff. Sur les ensembles mesurables B dans l’espace transfini. Compositio Mathematica, 7:201–213, 1939. CODEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic).

Maximoff:1940:CP

[Max40a] Isaiah Maximoff. On a continuum of power 2^{ℵ_{i}}. Annals of Mathematics (2), 41:321–327, 1940. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Maximoff:1940:SE

[Max40b] Isaie Maximoff. Sur la séparabilité d’ensembles. Acad. Roum. Bull. Sect. Sci., 22:384–389, 1940.

Maximoff:1940:SSE

[Max40c] Isaie Maximoff. Sur le système de Souslin d’ensembles dans l’espace transfini. Bulletin of the American Mathematical Society, 46:543– 550, 1940. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Mantel:1975:IMB

[MBB^{+}75] N. Mantel, N. R. Bohidar, C. C. Brown, J. L. Ciminera, and John W. Tukey. An improved Mantel–Bryan procedure for “safety” testing of carcinogens. Cancer Research, 35(??):865–872, ???? 1975. CODEN CNREA8. ISSN 0008-5472.

McNair:1945:SAG

[MBT^{+}45] W. A. McNair, H. W. Bode, G. N. Thayer, B. D. Holbrook, and J. W. Tukey. A study of an antiaircraft guided missile system. Report, Bell Laboratories, Murray Hill, NJ, USA, 1945.

McCullagh:2003:JWT

[McC03] Peter McCullagh. John Wilder Tukey: 16 June 1915–26 July 2000. Biographical Memoirs of Fellows of the Royal Society, 49 (??):537–555, December 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.jstor. org/stable/3650242. Supplement contains a 16-page biography of 431 publications by J. Wilder Tukey.

McGrayne:2011:TWH

[McG11] Sharon Bertsch McGrayne. The theory that would not die: how Bayes’ rule cracked the Enigma code, hunted down Russian submarines, and emerged triumphant from two centuries of controversy. Yale University Press, New Haven, CT, USA, 2011. ISBN 0-300-16969-8. xiii + 320 pp. LCCN QA279.5 .M415 2011.

McNemar:1955:RSP

[McN55] Quinn McNemar. Review: Statistical Problems of the Kinsey Report on Sexual Behavior in the Human Male, by W. G. Cochran, F. Mosteller, and J. W. Tukey. Science (New Series), 122(3161):206, July 29, 1955. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/1749913.

McVittie:1956:GSR

[McV56] G. C. McVittie. Galaxies, statistics and relativity. In Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability, 1954–1955, vol. III, pages 69–74. University of California Press, Berkeley, CA, USA, 1956.

Meyer:1956:SMC

[Mey56] Herbert A. Meyer, editor. Symposium on Monte Carlo Methods: held at the University of Florida, March 16 and 17, 1954. Wiley, New York, NY, USA, 1956. LCCN QA273 U577.

Marco:1987:SSR

[MHD87] Gino J. Marco, Robert M. Hollingworth, and William Durham, editors. Silent spring revisited. American Chemical Society, Washington, DC, 1987. ISBN 0-8412-0980-4, 0-8412-0981-2 (paperback).

LCCN QH545.P4 S55 1987. URL http://www.loc.gov/catdir/
enhancements/fy0604/86025952-d.html;   http://www.loc.
gov/catdir/enhancements/fy0604/86025952-t.html.

Miller:1974:JR

[Mil74] Rupert G. Miller. The jackknife — a review. Biometrika, 61(1):1– 15, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334280.

See [Que49, Que56, Efr79].

Miller:1978:IFT

[Mil78] John J. Miller. The inverse of the Freeman–Tukey double arcsine transformation. The American Statistician, 32(4):138, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682942.

Mildenberger:1997:NCG

[Mil97] Heike Mildenberger. Non-constructive Galois–Tukey connections. Journal of Symbolic Logic, 62(4):1179–1186, ???? 1997. CO-DEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

URL http://projecteuclid.org/euclid.jsl/1183745374;
http://www.jstor.org/stable/2275635.

Molina:1946:SFC

[Mol46] Edward C. Molina. Some fundamental curves for the solution of sampling problems. Annals of Mathematical Statistics, 17:325– 335, 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Morita:1940:USD

[Mor40] Kiiti Morita. On uniform spaces and the dimension of compact spaces. Proc. Phys.-Math. Soc. Japan (3), 22:969–977, 1940.

Morgenthaler:2003:JWT

[Mor03] Stephan Morgenthaler. John W. Tukey as teacher. Statistical Science, 18(3):342–345, 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1076102423.

Mosteller:1984:BJW

[Mos84] F. Mosteller. Biography of John W. Tukey. In Brillinger [Bri84a], pages xv–xvii. ISBN 0-534-03303-2. LCCN QA276.A12 T85 1984. With introductory material by William S. Cleveland and Frederick Mosteller.

Mosteller:2005:JWT

[Mos05] Frederick Mosteller. John W. Tukey: 16 June 1915–26 July 2000: Biographical memoir. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 149 (4):625–630, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.amphilsoc.

org/sites/default/files/490414.pdf.

Mosteller:1991:MST

[MPT91] F. Mosteller, A. Parunak, and John W. Tukey. Mean squares, F tests, and estimates of variance. In D. C. Hoaglin, F. Mosteller, and J. W. Tukey, editors, Fundamentals of Exploratory Analysis of Variance, chapter 7, pages 146–164. Wiley, New York, NY, USA,

Mosteller:1949:PANa

[MT49a] F. Mosteller and John W. Tukey. Practical applications of new theory and scale. A review. Part I: Location and scale: Tables. Industrial Quality Control, 6(??):5–8, September 1949. CODEN ???? ISSN 0884-822X.

Mosteller:1949:PANb

[MT49b] F. Mosteller and John W. Tukey. Practical applications of new theory and scale. A review. Part II: Counted data—graphical methods. Industrial Quality Control, 6(??):5–7, October 1949. CODEN ???? ISSN 0884-822X.

Mosteller:1949:QAd

[MT49c] Frederick Mosteller and John W. Tukey. Questions and answers. The American Statistician, 3(4):12, October 1949. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681357.

Mosteller:1949:UUB

[MT49d] Frederick Mosteller and John W. Tukey. The uses and usefulness of binomial probability paper. Journal of the American Statistical Association, 44(246):174–212, June 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2280595.

Mosteller:1950:PANa

[MT50a] F. Mosteller and John W. Tukey. Practical applications of new theory. A review. Part III — analytical techniques. Industrial Quality Control, 6(??):5–7, January 1950. CODEN ???? ISSN 0884-822X.

Mosteller:1950:PANb

[MT50b] F. Mosteller and John W. Tukey. Practical applications of new theory. A review. Part IV — gathering information. Industrial Quality Control, 6(??):5, March 1950. CODEN ???? ISSN 0884-822X.

Mosteller:1950:SLS

[MT50c] Frederick Mosteller and John W. Tukey. Significance levels for a k-sample slippage test. Annals of Mathematical Statistics, 21 (1):120–123, ???? 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729892.

Moore:1954:AQa

[MT54a] Peter G. Moore and John W. Tukey. Answer to Query 112. Biometrika, 10(??):562–568, ???? 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Moore:1954:AQb

[MT54b] Peter G. Moore and John W. Tukey. Answer to Query 112. Biometrika, 11(??):113, ???? 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Mosteller:1968:DAI

[MT68] F. Mosteller and J. W. Tukey. Data analysis including statistics. In Gardner Lindzey and Elliot Aronson, editors, The Handbook of Social Psychology, volume 2, chapter 10, pages 80–203. Addison-Wesley, Reading, MA, USA, second edition, 1968. LCCN HM251 .L486 1968. Five volumes.

McNeil:1975:HOD

[MT75] D. R. McNeil and J. W. Tukey. Higher-order diagnosis of two-way tables, illustrated on two sets of demographic empirical distributions. Biometrics, 31(2):487–510, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://

www.jstor.org/stable/2529433.

Mosteller:1977:DAR

[MT77] Frederick Mosteller and John W. Tukey. Data Analysis and Regression: a Second Course in Statistics. Addison-Wesley, Reading, MA, USA, 1977. ISBN 0-201-04854-X. xvii + 588 pp. LCCN QA276 .M698 1977.

McCarthy:1978:EAA

[MT78] John L. McCarthy and John W. Tukey. Exploratory analysis of aggregate voting behavior: Presidential elections in New Hampshire, 1896–1972. Social Science History, 2(3):292–331, Spring 1978. CODEN ???? ISSN 0145-5532. URL http://www.jstor.

org/stable/1171133.

Mallows:1982:OTD

[MT82a] Colin L. Mallows and John W. Tukey. An overview of techniques of data analysis, emphasizing its exploratory aspects. In J. Tiago de Oliveira and B. Epstein, editors, Some Recent Advances in Statistics, pages 111–172. Academic Press, New York, USA, 1982.

Mosteller:1982:ADR

[MT82b] F. Mosteller and J. W. Tukey. Analiz Dannykh i Regressiya. Financy i Statistika Press, Moskva, USSR, 1982. ISBN ???? ???? pp. LCCN ???? Russian edition of Data Analysis and Regression: A Second Course in Statistics. Translated from English by Yu. N. Blagovshchenskogo, editing and preface by Yu. P. Adleva. Volumes I and II.

Mosteller:1982:CRS

[MT82c] Frederick Mosteller and John W. Tukey. Combination of results of stated precision. I. The optimistic case. Utilitas Mathematica, 21 (A):155–178, 1982. CODEN UTMADA. ISSN 0315-3681. Special issue dedicated to Frank Yates on the occasion of his eightieth birthday, Vol. A.

Morgenthaler:1983:UDS

[MT83] Stephan Morgenthaler and John W. Tukey. The use of double sampling in studying robustness. Technical report, series 2 252, Princeton University, Princeton, NJ, USA, 1983. 16 pp.

Mendoza:1984:SAV

[MT84a] Carlos Mendoza and John W. Tukey. Simple approximate variances for better values; a sampling enquiry. Technical report, series 2 269, Princeton University, Princeton, NJ, USA, 1984. 24 pp.

Mosteller:1984:CRS

[MT84b] F. Mosteller and John W. Tukey. Combination of results of stated precision: II. A more realistic case. In P. S. R. S. Rao and J. Sedransk, editors, W. G. Cochran’s Impact on Statistics, pages 223– 252. Wiley, New York, NY, USA, 1984.

Morgenthaler:1990:NFD

[MT90] S. Morgenthaler and John W. Tukey. The next future of data analysis. In E. Diday, editor, Data Analysis, Learning Symbolic and Numeric Knowledge: Proceedings of the Conference on Data Analysis, Learning Symbolic and Numeric Knowledge, Antibes, France, September 11–14, 1989, page ?? Nova Science Publishers, New York, NY, USA, 1990. ISBN 0-941743-64-0. LCCN QA278 .C75 1989 (LC).

Morgenthaler:1991:KIO

[MT91a] S. Morgenthaler and J. W. Tukey. Key ideas and outline. In Morgenthaler and Tukey [MT91d], pages 21–36. ISBN 0-471-52372-0. LCCN QA276.6 .C66 1991.

Morgenthaler:1991:PEL

[MT91b] S. Morgenthaler and J. W. Tukey. Point estimation of location: Technical choices. In Morgenthaler and Tukey [MT91d], chapter 7, pages 87–108. ISBN 0-471-52372-0. LCCN QA276.6 .C66 1991.

Morgenthaler:1991:B

[MT91c] S. Morgenthaler and John W. Tukey. Background. In Morgenthaler and Tukey [MT91d], pages 1–8. ISBN 0-471-52372-0. LCCN QA276.6 .C66 1991.

Morgenthaler:1991:CPR

[MT91d] Stephen Morgenthaler and John W. Tukey, editors. Configural Polysampling: a Route to Practical Robustness. Wiley, New York, NY, USA, 1991. ISBN 0-471-52372-0. xiv + 228 pp. LCCN QA276.6 .C66 1991.

Mosteller:1991:PAD

[MT91e] F. Mosteller and John W. Tukey. Purposes of analyzing data that come in a form inviting us to apply tools from the analysis of variance. In D. C. Hoaglin, F. Mosteller, and J. W. Tukey, editors, Fundamentals of Exploratory Analysis of Variance, pages 24–39. Wiley, New York, NY, USA, 1991.

Morgenthaler:1996:IDL

[MT96] Stephan Morgenthaler and John W. Tukey. Inference for the direction of the larger of two eigenvectors: the case of circular elongation. In Robust statistics, data analysis, and computer intensive methods (Schloss Thurnau, 1994), volume 109 of Lecture Notes in Statistics, pages 321–352. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996.

Morgenthaler:2000:FQD

[MT00] Stephan Morgenthaler and John W. Tukey. Fitting quantiles: doubling, HR, HQ, and HHH distributions. Journal of Computational and Graphical Statistics, 9(1):180–195, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-

2715 (electronic). URL http://www.amstat.org/publications/
jcgs/abstracts00/Morgenthaler.htm;   http://www.jstor.
org/stable/1390620.

Morgenthaler:2001:MTW

[MT01a] Stephan Morgenthaler and John W. Tukey. Multipolishing and two-way plots. Metrika. International Journal for Theoretical and Applied Statistics., 53(3):245–267, 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Morgenthaler:2001:TWP

[MT01b] Stephan Morgenthaler and John W. Tukey. Two-way plots. I. The case of straight line segments. Journal of Statistical Computation and Simulation, 71(4):287–312, 2001. CODEN JSCSAJ. ISSN 0094-9655 (print), 1563-5163 (electronic).

Mantel:1982:TAI

[MTCH82] N. Mantel, John W. Tukey, J. L. Ciminera, and J. F. Heyse. Tumorigenicity assays, including use of the jackknife. Biometrical Journal, 24(6):579–596, ???? 1982. CODEN BIJODN. ISSN 0323-3847.

McGill:1978:VBP

[MTL78] Robert McGill, John W. Tukey, and Wayne A. Larsen. Variations of box plots. The American Statistician, 32(1):12–16, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683468.

Mosteller:1991:AC

[MTY91] F. Mosteller, John W. Tukey, and C. S. Youtz. Assessing changes. In D. C. Hoaglin, F. Mosteller, and J. W. Tukey, editors, Fundamentals of Exploratory Analysis of Variance, pages 295–335. Wiley, New York, NY, USA, 1991.

Nair:1948:DED

[Nai48a] K. R. Nair. The distribution of the extreme deviate from the sample mean and its Studentized form. Biometrika, 35(1/2):118– 144, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332634.

Nair:1948:SFE

[Nai48b] K. R. Nair. The Studentized form of the extreme mean square test in the analysis of variance. Biometrika, 35:16–31, 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Nelder:1994:DSQ

[NGL^{+}94] J. A. Nelder, T. Greenfield, H. J. Lenz, C. Chatfield, D. A. Preece, C. E. Lunneborg, M. C. Jones, J. Gower, R. A. Stone, M. C. Fessey, S. J. W. Evans, T. Lewis, A. S. C. Ehrenberg, D. J. Finney, A. M. Herzberg, P. Lovie, A. D. Lovie, R. J. Mackay, R. W. Oldford, I. W. Molenaar, P. C. Obrien, H. Rouanet, T. M. F. Smith, J. Tukey, M. Wise, and J. Zighera. Deconstructing statistical questions — discussion. Journal of the Royal Statistical Society Series A — Statistics in Society, 157(??):338–356, ???? 1994. ISSN 0035-9238.

Novikoff:1939:PSS

[Nov39] P. C. Novikoff. On projections of some B-sets. Doklady Akad. Nauk SSSR (N.S.), 23:863–864, 1939.

Nowack:1994:DFA

[Now94] Robert Nowack. Development of the FFT and applications in geophysics. In Brown et al. [BCEP94], pages 395–397. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993. URL http://web.ics.purdue.

edu/~nowack/nowackpubs-dir/Lanczos-nowackfft1994.pdf.

Neyman:1956:SIG

[NS56] Jerzy Neyman and Elizabeth L. Scott. Statistics of images of galaxies with particular reference to clustering. In Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability, 1954–1955, vol. III, pages 75–111. University of California Press, Berkeley, CA, USA, 1956.

Obreanu:1947:PCC

[Obr47] F. Obreanu. La puissance de certaines classes de fonctions. Duke Mathematical Journal, 14:377–380, 1947. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Odone:1942:CPS

[Odo42] Vincenzo Odone. Il collaudo di prodotti in serie ed il calcolo delle probabilità. Atti Accad. Sci. Torino. Cl. Sci. Fis. Mat. Nat., 77: 407–430, 1942. ISSN 0001-4419.

Olmstead:1963:CEA

[Olm63] Paul Olmstead. Correction: The Examination and Analysis of Residuals, by F. J. Anscombe and John W. Tukey. Technometrics, 5(4):536–537, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Osborn:1978:BRBb

[Osb78a] J. F. Osborn. Book review: Data Analysis and Regression, by F. Mosteller; John W. Tukey. Journal of the Royal Statistical Society. Series A (General), 141(4):549–550, ???? 1978. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.jstor.org/stable/2344493.

Osborn:1978:BRBa

[Osb78b] J. F. Osborn. Book review: Exploratory Data Analysis, by John W. Tukey. Journal of the Royal Statistical Society. Series A (General), 141(4):548–549, ???? 1978. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/2344492.

Osborn:1978:RDA

[Osb78c] J. F. Osborn. Review: Data Analysis and Regression, by F. Mosteller and John W. Tukey. Journal of the Royal Statistical Society. Series A (General), 141(4):549–550, ???? 1978. CODEN

JSSAEF. ISSN 0035-9238. URL http://www.jstor.org/stable/
2344493.

Osborn:1978:RED

[Osb78d] J. F. Osborn. Review: Exploratory Data Analysis, by John W. Tukey. Journal of the Royal Statistical Society. Series A (General), 141(4):548–549, ???? 1978. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/2344492.

Olmstead:1947:CTA

[OT47] Paul S. Olmstead and John W. Tukey. A corner test for association. Annals of Mathematical Statistics, 18(4):495–513, ???? 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://projecteuclid.org/euclid.aoms/
1177730341.

Otchan:1939:EFE

[Otc39] G. Otchan. Sur l’équivalence des familles d’ensembles mesurables B. C. R. (Doklady) Acad. Sci. URSS (N. S.), 23:753–755, 1939.

Parzen:1979:NSDa

[Par79] Emanuel Parzen. Nonparametric statistical data modeling. Journal of the American Statistical Association, 74(365):105–121, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2286734. See comments [Tuk79a].

Pedersen:1991:SSS

[PCT91] Jan Pedersen, Doug Cutting, and John Tukey. Snippet search: a single phrase approach to text access. In Proceedings of the 1991 Joint Statistical Meetings, page ?? American Statistical Association, 1991. Also available as Xerox PARC Technical Report SSL-91-08.

Pearson:1955:SCT

[Pea55] E. S. Pearson. Statistical concepts in their relation to reality. Journal of the Royal Statistical Society. Series B (Methodological), 17: 204–207, 1955. CODEN JSTBAJ. ISSN 0035-9246.

Penrose:1947:SND

[Pen47] L. S. Penrose. Some notes on discrimination. Ann. Eugenics, 13: 228–237, 1947.

Perks:1947:SOI

[Per47] Wilfred Perks. Some observations on inverse probability including a new indifference rule. J. Inst. Actuar., 73:285–312; discussion, 313–334, 1947.

Petard:1938:CMT

[Pét38] H. Pétard. A contribution to the mathematical theory of big game hunting. American Mathematical Monthly, 45(7):446–447, August/September 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pedersen:1994:ITP

[PHC^{+}94] Jan O. Pedersen, Per-Kristian Halvorsen, Douglass R. Cutting, John W. Tukey, Eric A. Bier, and Daniel G. Bobrow. Iterative technique for phrase query formation and an information retrieval system employing same. US Patent 5,278,980., 1994. URL http:

//www.google.com/patents/US5278980. Filed 16 August 1991.

Issued 11 January 1994.

Pedersen:1995:SGC

[PKCT95] Jan. O. Pedersen, David Karger, Douglass R. Cutting, and John W. Tukey. Scatter-gather: a cluster-based method and apparatus for browsing large document collections. US

Patent 5,442,778., 1995. URL http://www.google.com/patents/

US5442778. Field 12 November 1991. Issued 15 August 1995.

Pondiczery:1938:QDN

[Pon38] E. S. Pondiczery. Questions, discussions, and notes: a functiontheoretical paradox. American Mathematical Monthly, 45(5):307, May 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Prescott:1975:BRB

[Pre75] P. Prescott. Book review: Robust Estimates of Location: Survey and Advances, by D. R. Andrews; P. J. Bickel; F. R. Hampei; P. J. Huber; W. H. Rogers; J. W. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):308–309, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/stable/2987928.

Priestley:1985:BRB

[Pri85] M. B. Priestley. Book review: The Collected Works of John W. Tukey. Volume 1: Time Series, 1949–1964, by D. R. Brillinger. Journal of the Royal Statistical Society. Series A (General), 148 (3):280, ???? 1985. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2981973.

Pierce:1948:CRD

[PST48] John R. Pierce, Claude E. Shannon, and John W. Tukey. Cathoderay device. US Patent 2,576,040., 1948. URL http://www.google. com/patents/US2576040. Filed March 10, 1948. Issued November 20, 1951.

Peisakoff:1950:RSA

[PT50] Melvin P. Peisakoff and John W. Tukey. Review of Sequential analysis by Abraham Vald. Bulletin of the American Mathematical Society, 56(2):200–202, ???? 1950. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Press:1956:PSMa

[PT56a] Harry Press and John W. Tukey. Power spectral methods of analysis and their application to problems in airplane dynamics. In Flight Test Manual, pages 1–41. NATO Advisory Group for Aeronautical Research and Development, ????, 1956.

Press:1956:PSMb

[PT56b] Harry Press and John W. Tukey. Power spectral methods of analysis and their application to problems in airplane dynamics. Bell Telephone System Monograph, 2606(??):i + 41, 1956. CODEN ????

Pearson:1965:AMS

[PT65] E. S. Pearson and J. W. Tukey. Approximate means and standard deviations based on distances between percentage points of frequency curves. Biometrika, 52(3/4):533–546, December 1965. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333703.

Pregibon:1981:ABR

[PT81] Daryl Pregibon and John W. Tukey. Assessing the behavior of robust estimates of location in small samples: introduction to configural polysampling. Technical report, series 2 185, Princeton University, Princeton, NJ, USA, 1981. 12 pp.

Pedersen:1997:MAA

[PT97] Jan O. Pedersen and John W. Tukey. Method and apparatus for automatic document summarization. US Patent 5,638,543., 1997.

URL http://www.google.com/patents/US5638543. Filed 3 June

1993. Issued 10 June 1997.

Peach:1947:QA

[PTM47] Paul Peach, John W. Tukey, and Frederick Mosteller. Questions and answers. The American Statistician, 1(3):17–18, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682536.

Quade:1974:BRBa

[Qua74] Dana Quade. Book review: The Statistics Cum Index by James L. Dolby; John W. Tukey. Technometrics, 16(4):634, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267621.

Quenouille:1949:ATC

[Que49] M. H. Quenouille. Approximate tests of correlation in time-series. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):68–84, ???? 1949. CODEN JSTBAJ. ISSN 0035-9246.

URL http://www.jstor.org/stable/2983696. See extension in

[Que56, Tuk58a], and reviews [Mil74, Efr79].

Quenouille:1956:NBE

[Que56] M. H. Quenouille. Notes on bias in estimation. Biometrika, 43 (3–4):353–360, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2332914. This extends earlier work [Que49, Tuk58a] on what is now known as the Quenouille–Tukey jacknife technique. See reviews [Mil74, Efr79].

Ramachandran:1956:TTE

[Ram56] K. V. Ramachandran. On the Tukey test for the equality of means and the Hartley test for the equality of variances. Annals of Mathematical Statistics, 27(3):825–831, ???? 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728189.

Rao:1945:IAA

[Rao45] C. Radhakrishna Rao. Information and the accuracy attainable in the estimation of statistical parameters. Bull. Calcutta Math. Soc., 37:81–91, 1945. ISSN 0008-0659.

Read:1993:FTC

[Rea93] Campbell B. Read. Freeman–Tukey chi-squared goodness-of-fit statistics. Statistics & Probability Letters, 18(4):271–278, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390015B.

Richardson:1946:TLC

[Ric46] J. T. Richardson. A table of Lagrangian coefficients for logarithmic interpolation of standard statistical tables to obtain other probability levels. Suppl. J. Roy. Statist. Soc., 8:212–215, 1946.

Richter:1949:M

[Ric49] H. Richter. Zur Maximalkorrelation. Zeitschrift für Angewandte Mathematik und Mechanik, 29:127, 1949. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Rider:1939:CMF

[Rid39] P. Rider. Certain moment functions for Fisher’s K-statistics in samples from a finite population. Acta [Trudy] Univ. Asiae Mediae. Ser. V-a., 1939(Fasc. 30):13, 1939.

Runge:1924:VNR

[RK24] Carl Runge and H. König. Vorlesungen über Numerisches Rechnung. (German) [Lectures on numerical calculation]. In Die Grundlehren der Mathematischen Wissenschaften. (German) [The principles of mathematical science], volume 11, pages 211–237. Verlag von Julius Springer, Berlin, Germany, 1924. LCCN ????

Ramachandran:1957:DPB

[RK57] K. V. Ramachandran and C. G. Khatri. On a decision procedure based on the Tukey statistic. Annals of Mathematical Statistics, 28(3):802–806, ???? 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706897.

Robertson:1972:RRE

[Rob72] C. A. Robertson. Review: Robust Estimates of Location: Survey and Advances, by D. R. Andrews, P. J. Bickel, F. R. Hampei, P. J. Huber, W. H. Rogers, and J. W. Tukey. Biometrics, 28(4): 1144–1145, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rollero:1943:SCG

[Rol43] Aldo Rollero. Sul calcolo grafico di un limite. Atti Accad. Ligure, 3:277–282 (1946), 1943.

Roma:1946:MDP

[Rom46] Maria Sofia Roma. Il metodo dell’ortogonalizzazione per la risoluzione numerica dei sistemi di equazioni lineari algebriche. Ricerca Sci., 16:309–312, 1946.

Roma:1947:MDP

[Rom47] Maria Sofia Roma. Il metodo dell’ortogonalizzazione per la risoluzione numerica dei sistemi di equazioni algebriche. Consiglio Naz. Ricerche. Pubbl. Ist. Appl. Calcolo, 1947(189):12, 1947.

Rosenblatt:1963:PST

[Ros63] Murray Rosenblatt, editor. Proceedings of the Symposium on Time Series Analysis, held at Brown University, June 11–14, 1962. Wiley, New York, NY, USA, 1963. LCCN QA280.

Rosenberger:1983:RCP

[Ros83] J. L. Rosenberger. Review: Configural Polysampling: A Route to Practical Robustness, by S. Morgenthaler and J. W. Tukey. Biometrics, 39(4):1284, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://www.

jstor.org/stable/2532280.

Rousseeuw:1999:BBB

[RRT99] Peter J. Rousseeuw, Ida Ruts, and John W. Tukey. The bagplot: a bivariate boxplot. The American Statistician, 53(4):382–387, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686061.

Rogers:1972:USL

[RT72] William H. Rogers and John W. Tukey. Understanding some longtailed symmetrical distributions. Statistica Neerlandica. Journal of the Netherlands Society for Statistics and Operations Research, 26(3):211–226, 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). Collection of articles in honour of Professor H. C. Hamaker.

Ross:1973:ISP

[RT73] Ian C. Ross and John W. Tukey. Index to Statistics and Probability. Locations and Authors, volume 5 of The Information Access Series. R&D Press, Los Altos, CA, USA, 1973. ISBN 0-88274-004-0. xxiii + 680 + 412 pp. LCCN Z6654.M33 R6.

Ross:1975:ISP

[RT75] Ian C. Ross and John W. Tukey. Index to Statistics and Probability. Permuted Titles, volume 3–4 of The Information Access Series. R&D Press, Los Altos, CA, USA, 1975. xviii + 1588 pp. Volume 1, A–Microbiology; Volume 2, Microclimatic–Z.

Rudnick:1966:NCF

[Rud66] Philip Rudnick. Note on the calculation of Fourier series (in Technical Notes and Short Papers). Mathematics of Computation, 20 (95):429–430, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Runge:1903:ZEG

[Run03] Carl Runge. Über die Zerlegung empirisch gegebener Periodischer Funktionen in Sinuswellen. (German) [On the representation of empirically-given periodic functions by sine waves]. Zeitschrift für Mathematik und Physik, 48(??):443–456, ???? 1903.

Runge:1905:ZEF

[Run05] Carl Runge. Über die Zerlegung einer empirischen Funktion in Sinuswellen. (German) [On the representation of an empirical function by sine waves]. Zeitschrift für Mathematik und Physik, 52(??): 117–123, ???? 1905.

Rathbun:1972:RPC

[RWM^{+}72] Daniel B. Rathbun, W. Allen Wallis, Frederick Mosteller, Ansley J. Coale, Paul M. Densen, Solomon Fabricant, Robert D. Fisher, W. Braddock Hickman, William Kruskal, Stanley Lebergott, Richard M. Scammon, William H. Shaw, Frank D. Stella, James A. Suffridge, and John W. Tukey. The report of the President’s Commission on Federal Statistics. The American Statistician, 26(1):15–18, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682840.

Salsburg:2001:LTT

[Sal01] David Salsburg. The Lady Tasting Tea: How Statistics Revolutionized Science in the Twentieth Century. W. H. Freeman, New York, NY, USA, 2001. ISBN 0-7167-4106-7 (hardcover), 0-8050-7134-2 (paperback). xi + 340 pp. LCCN Q175 .S2345 2001. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0667/00049523-b.
html;   http://catdir.loc.gov/catdir/enhancements/fy0667/
00049523-d.html.

Samuelson:1945:CIP

[Sam45] Paul A. Samuelson. A convergent iterative process. J. Math. Phys. Mass. Inst. Tech., 24:131–134, 1945. ISSN 0097-1421.

Sura-Bura:1940:EBC

[ŜB40] M. R. Ŝura-Bura. Espaces bicompacts comme images des discontinus. C. R. (Doklady) Acad. Sci. URSS (N.S.), 27:431–435, 1940.

Sura-Bura:1941:TBR

[ŠB41] M. Šura-Bura. Zur Theorie der bikompakten Räume. Rec. Math. [Mat. Sbornik] N. S., 9 (51):385–388, 1941.

Smelser:2001:IES

[SB01] Neil J. Smelser and Paul B. Baltes, editors. International Encyclopedia of the Social and Behavioral Sciences. Elsevier, Amsterdam, The Netherlands, 2001. ISBN 0-08-043076-7 (26-volume set). ????

pp. LCCN H41 .I58 2001. URL http://www.loc.gov/catdir/
enhancements/fy0612/2001044791-d.html.

Solis-Cohen:2000:FTS

[SC00] L. Solis-Cohen. Freeman’s Tukey sale. Maine Antique Digest, ?? (??):??, June 2000.

Schuenemeyer:1993:BRBa

[Sch93] John H. Schuenemeyer. Book review: Fundamentals of Exploratory Analysis of Variance by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Technometrics, 35(1):89, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269297.

Schneider:2012:FFF

[Sch12] David Schneider. A faster Fast Fourier Transform. IEEE Spectrum, 49(3):12–13, March 2012. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1980:BRB

[SG80] J. T. Smith and M. P. Griffin. Book review: Exploratory Data Analysis, by John W. Tukey. Technometrics, 22(1):129–130, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268394.

Schimmel:1978:GDS

[SGT^{+}78] H. Schimmel, I. Goldstein, J. Tukey, R. W. Buechley, I. T. Higgins, and P. Bloomfield. General discussion: Session II. Bulletin of the New York Academy of Medicine, 54(11):1132–1136, ???? 1978. ISSN 0028-7091.

Shannon:1948:MTC

[Sha48] Claude E. Shannon. A mathematical theory of communication. The Bell System Technical Journal, 27(3):379–423, July 1948.

CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.bell-

labs.com/BSTJ/images/Vol27/bstj27-3-379.pdf. Reprinted in

[Sle74]. From the first page: “If the base 2 is used the resulting units may be called binary digits, or more briefly, bits, a word suggested by J. W. Tukey.”. This is the first known printed instance of the word ‘bit’ with the meaning of binary digit.

Shepp:2007:STT

[She07] Larry Shepp. Statistical thinking: From Tukey to Vardi and beyond. In Liu et al. [LSZ07], pages 268–273. ISBN 0-940600-70-

6. LCCN ???? URL http://projecteuclid.org/euclid.lnms/
1196794959.

Simaika:1946:NMF

[Sim46] J. B. Simaika. Note on M. Fréchet index of correlation. Proc. Math. Phys. Soc. Egypt, 3:21–22, 1946. ISSN 1110-0613.

Simonoff:1987:CPS

[Sim87] Jeffrey S. Simonoff. Comment on: “Performance of some resistant rules for outlier labeling” [J. Amer. Statist. Assoc. 81 (1986), no. 396, 991–999] by D. C. Hoaglin, B. Iglewicz and J. W. Tukey. With reply by Hoaglin, Iglewicz and Tukey. Journal of the American Statistical Association, 82(398):703–704, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289514. See [HIT86].

Singleton:1967:CFF

[Sin67] Richard C. Singleton. On computing the fast Fourier transform. Communications of the ACM, 10(10):647–654, October 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Slepian:1974:KPD

[Sle74] David Slepian, editor. Key papers in the development of information theory. IEEE Press selected reprint series. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1974. ISBN 0-87942-027-8, 0-87942-028-6 (paperback). vi + 463 pp. LCCN Q360 .S54.

Smith:1947:SED

[Smi47] Cedric A. B. Smith. Some examples of discrimination. Ann. Eugenics, 13:272–282, 1947.

Somerville:1993:CTK

[Som93] Paul N. Somerville. On the conservatism of the Tukey–Kramer multiple comparison procedure. Statistics & Probability Letters, 16(5):343–345, April 8, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390067S.

Speed:2002:JWT

[Spe02] T. P. Speed. John W. Tukey’s contributions to analysis of variance. Annals of Statistics, 30(6):1649–1665, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1043351252;   http://www.jstor.org/stable/1558735.

Stone:1942:GST

[ST42] A. H. Stone and John W. Tukey. Generalized “sandwich” theorems. Duke Mathematical Journal, 9(2):356–359, ???? 1942. CO-DEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

URL http://projecteuclid.org/euclid.dmj/1077493229.

Scheffe:1944:FSS

[ST44] H. Scheffé and J. W. Tukey. A formula for sample sizes for population tolerance limits. Annals of Mathematical Statistics, 15(2):217, ???? 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://www.jstor.org/stable/2236202.

Scheffe:1945:NPE

[ST45] H. Scheffé and J. W. Tukey. Non-parametric estimation. I. Validation of order statistics. Annals of Mathematical Statistics, 16(2): 187–192, ???? 1945. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic).  URL http://projecteuclid.org/
euclid.aoms/1177731119;   http://www.jstor.org/stable/
2235830.

Spitzer:1949:IPG

[ST49] Lyman Spitzer, Jr. and John W. Tukey. Interstellar polarization, galactic magnetic fields, and ferromagnetism. Science, 109(2836): 461–462, May 6, 1949. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Spitzer:1951:TIP

[ST51] L. Spitzer, Jr. and John W. Tukey. A theory of interstellar polarization. Astrophysical Journal, 114(??):187–205, ???? 1951. CO-DEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Siegel:1960:NSR

[ST60] Sidney Siegel and John W. Tukey. A nonparametric sum of ranks procedure for relative spread in unpaired samples. Journal of the American Statistical Association, 55(291):429–445, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281906. See cor-

rection in this journal, 56, 1005–1005, 1961.

Siegel:1961:CNS

[ST61] Sidney Siegel and John W. Tukey. Corrigenda: a nonparametric sum of ranks procedure for relative spread in unpaired samples. Journal of the American Statistical Association, 56(296): 1005, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282014.

Seheult:1982:SRP

[ST82] Allan Seheult and John W. Tukey. Some resistant procedures for analyzing 2^{n} factorial experiments. Utilitas Mathematica, 21(B): 57–98, 1982. CODEN UTMADA. ISSN 0315-3681. Special issue dedicated to Frank Yates on the occasion of his eightieth birthday, Vol. B.

Sivak:1991:CPE

[ST91] A. Sivak and John W. Tukey. Comments on predicting EPA’s forthcoming CO standards in light of new clinical evidence. Risk Analysis, 11(4):575–576, ???? 1991. CODEN RIANDF. ISSN 0272-4332 (print), 1539-6924 (electronic).

Seheult:2001:TRA

[ST01] Allan H. Seheult and John W. Tukey. Towards robust analysis of variance. In A. K. Md. E. Saleh, editor, Data Analysis from Statistical Foundations: a festschrift in honour of the 75th birthday of D. A. S. Fraser, pages 217–424. Nova Science Publishers, Huntington, NY, USA, 2001. URL http:/

/dro.dur.ac.uk/4357/;   http://www.novapublishers.com/
catalog/product_info.php?products_id=656.

Stebakov:1939:FTB

[Ste39] S. Stebakov. Über der Fundamentalsatz der Theorie der bikompakten Räume. Uchenye Zapiski Moskov. Gos. Univ. Matematika, 30:161–164, 1939.

Steenrod:1941:RJW

[Ste41] N. E. Steenrod. Review: J. W. Tukey, Convergence and Uniformity in Topology. Bulletin of the American Mathematical Society, 47(5):353–354, ???? 1941. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://projecteuclid.

org/euclid.bams/1183503627.

Stevens:1942:MTS

[Ste42] W. L. Stevens. Mathematical theory of some distributions used in statistics. Revista Fac. Ci. Univ. Coimbra, 10:247–288, 1942.

Steinhaus:1945:DEE

[Ste45a] Hugo Steinhaus. Sur la division des ensembles de l’espace par les plans et des ensembles plans par les cercles. Fundamenta mathematicae, 33:245–263, 1945. ISSN 0016-2736 (print), 1730-6329 (electronic).

Stevens:1945:ATA

[Ste45b] W. L. Stevens. Application of the χ^{2} test to the analysis of variance. Revista Fac. Ci. Univ. Coimbra, 13:4–17, 1945.

Steinhaus:1947:CTB

[Ste47] H. Steinhaus. Sur la cubature des troncs de bois. Colloquium Math., 1:23–28, 1947.

Stephan:1965:SW

[STM^{+}65] Frederick F. Stephan, John W. Tukey, Frederick Mosteller, Alex M. Mood, Morris H. Hansen, Leslie E. Simon, and W. J. Dixon. Samuel S. Wilks. Journal of the American Statistical Association, 60(312):939–966, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283397.

Stumpff:1937:GMP

[Stu37] Karl Stumpff. Grundlagen und Methoden der Periodenforschung. (German) [Foundation and methods of period research]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1937. ???? pp. LCCN ????

Stumpff:1939:TAH

[Stu39] Karl Stumpff. Tafeln und Aufgaben zu Harmonischen Analyse und Periodogrammrechnung. (German) [Tables and exercises in harmonic analysis and periodogram calculation]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1939. ???? pp. LCCN ????

Stuart:1984:BRB

[Stu84] Michael Stuart. Book review: Understanding Robust and Exploratory Data Analysis, by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 33(3):320–321, September 1984. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988240.

Tukey:1948:BCE

[T^{+}48] John W. Tukey et al. Biometrical clinic on entomological problems. In Chester Ittner Bliss, editor, Biometrical Clinic on Entomological Problems, New York City, December 13, 1948. Proceedings of a joint meeting of the American Association of Economic Entomologists and the Biometric Society, Eastern North American Region, page ?? Biometric Society, New Haven, CT, USA, 1948.

Tukey:1962:MMN

[T^{+}62a] John W. Tukey et al. Meeting manpower needs in science and technology. report no. 1. graduate training in engineering, mathematics and physical sciences. (A report of the President’s Science Advisory Committee). Report, The White House, Washington, DC, USA, 1962.

Tukey:1962:SBS

[T^{+}62b] John W. Tukey et al. Strengthening the behavioral sciences. Science, 136(3512):233–241, April 20, 1962. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Tukey:1966:SNH

[T^{+}66] John W. Tukey et al. Summary of the National Halothane Study. Journal of the American Medical Association, 197(??):775–788, ???? 1966. CODEN JAMAAP. ISSN 0098-7484 (print), 1538-3598 (electronic).

Tukey:1967:CUE

[T^{+}67] John W. Tukey et al. Computers in undergraduate education: Mathematics, physics, statistics, and chemistry. In Proceedings of a Conference Sponsored by The National Science Foundation, University of Maryland, December 8-9, 1967, page ?? ????, ????, 1967.

Tukey:1970:SNR

[T^{+}70a] John W. Tukey et al. 1969–1970 science: National results and illustrations of group comparisons. Report 1, National Assessment of Educational Progress, Ann Arbor, MI, USA, 1970.

Tukey:1970:CAN

[T^{+}70b] John W. Tukey et al. Cleaner Air for the Nation: President’s Task Force on Air Pollution. United States Government Printing Office, Washington, DC, USA, 1970.

Tukey:1971:FSV

[T^{+}71] John W. Tukey et al. Federal Statistics. Volumes I, II. United States Government Printing Office, Washington, DC, USA, 1971. ???? pp. LCCN ????

Tukey:1973:CH

[T^{+}73] John W. Tukey et al. Chemicals and health. Report, President’s Science Advisory Committee, Washington, DC, USA, 1973.

Tukey:1977:ROP

[T^{+}77] John W. Tukey et al. Response to ozone protection. Sections of the Clean Air Amendments, 1977.

Tukey:1978:TRP

[T^{+}78] John W. Tukey et al. The Twin Rivers Program on Energy Conservation in Housing: Highlights and conclusions. Energy and Buildings, 1(3):207–242, ???? 1978. CODEN ???? ISSN 0378-7788.

Tukey:1990:CCE

[T^{+}90a] John W. Tukey et al. 1990 Census coverage evaluation operations. In Hearing before the Subcommittee on Census and Population of the Committee on Post Office and Civil Service, House of Representatives, 101st Congress, Second Session, September 11, 1990, pages 1–142. United States Government Printing Office, Washington, DC, USA, 1990. LCCN ???? Serial No. 101-79.

Tukey:1990:PGS

[T^{+}90b] John W. Tukey et al. Proposed guidelines for statistical adjustment of the 1990 Census. In Hearing before the Subcommittee on Census and Population of the Committee on Post Office and Civil Service, House of Representatives, 101st Congress, Second Session, January 30, 1990, pages 177–190. United States Government Printing Office, Washington, DC, USA, 1990. LCCN ???? Serial No. 101-43.

Tukey:1991:ELN

[T^{+}91a] John W. Tukey et al. The experience and legacy of NAPAP (Report of the Oversight Review Board of the National Acid Precipitation Assessment Program). United States Government Printing Office, Washington, DC, USA, 1991.

Tukey:1991:OHR

[T^{+}91b] John W. Tukey et al. Oversight hearing to review the progress of coverage evaluation procedures. In Hearing before the Subcommittee on Census and Population, Committee on Post Office and Civil Service, House of Representatives, 102nd Congress, First Session, June 27, 1991, pages 1–95. United States Government Printing Office, Washington, DC, USA, 1991. LCCN ????

Tukey:1971:QPR

[TA71] John W. Tukey and D. Andrews. Quick plotting routines for teletypewriter-like terminal. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1971. ???? pp.

Tukey:1965:RQO

[TAB^{+}65] J. W. Tukey, M. Alexander, H. S. Bennett, N. C. Brady, J. C. Calhoun, Jr., J. C. Geyer, A. J. Haagen-Smit, N. Hackerman, J. B. Hartgering, D. Pimentel, R. Revelle, L. H. Roddis, W. H. Stewart, and J. L. Whittenberger. Restoring the quality of our environment: Report of the Environmental Pollution Panel, President’s Science Advisory Committee. United States Government Printing Office, Washington, DC, USA, 1965. xii + 317 pp. LCCN TD180 .U55.

Tukey:1977:IAA

[TACM77] John W. Tukey, B. Atal, J. Chang, and Max Mathews. Inversion of articulatory-to-acoustic transformation in the vocal-tract by a computer sorting technique. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1977. ???? pp.

Takeuchi:1973:RPS

[Tak73] K. Takeuchi. Review: a problem in statistics: Robust Estimates of Location. Survey and Advances, by D. F. Andrews, P. J. Bickel, F. R. Hampel, P. J. Huber, W. H. Rogers, and J. W. Tukey. Science (New Series), 179(4072):469–470, February 2, 1973. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/1735597.

Talbot:2000:RGA

[Tal00] M. Talbot. Review: Graphical Analysis of Multiresponse Data: Illustrated with a Plant Breeding Trial, by K. E. Basford and J. W. Tukey. Biometrics, 56(2):649–650, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://

www.jstor.org/stable/2677019.

Tukey:1966:EMP

[TB66] John W. Tukey and J. L. Buckley. Earth—Man’s polluted spaceship (the call of the vanishing wild—6). The Christian Science Monitor, ??(??):9, October 24, 1966. CODEN ???? ISSN 0145-8027.

Tukey:1982:SES

[TB82] John W. Tukey and David Brillinger. Spectrum estimation and system identification, particularly in the presence of noise. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1982. ???? pp.

Tropp:1984:AOT

[TB84] Henry S. Tropp and John Grist Brainerd. Anecdotes: Origin of the term Bit; The Soviets and the ENIAC. Annals of the History of Computing, 6(2):152–156, April/June 1984. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1984/pdf/a2152.pdf;   http://www.computer.org/
annals/an1984/a2152abs.htm. See [Buc81].

Tukey:1978:ADA

[TBBM78] J. W. Tukey, P. Bloomfield, H. I. Braun, and D. R. McNeill. Advances in data analysis. Unpublished manuscript cited in [BB02, page 1594]., 1978.

Tukey:1977:FPR

[TBS77] John W. Tukey, Henry I. Braun, and Michael Schwarzschild. Further progress on robust / resistant widthers. Technical report, series 2 129, Princeton University, Princeton, NJ, USA, 1977. 9 pp.

Tukey:1985:TSC

[TCH85] J. W. Tukey, J. L. Ciminera, and J. F. Heyse. Testing the statistical certainty of a response to increasing doses of a drug. Biometrics, 41(1):295–301, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://www.jstor.

org/stable/2530666.

Tukey:2000:EEP

[TCH00] John W. Tukey, J. L. Cox, and J. F. Heyse. Efficacy estimates from parasite count data that include zero counts. Experimental Parasitology, 96 (part 1)(??):1–8, ???? 2000. CODEN EXPAAA. ISSN 0014-4894 (print), 1090-2449 (electronic).

Tukey:1972:RRB

[TDD^{+}72] John W. Tukey, K. David, T. Dobzhansky, R. W. Gerard, H. B. Glass, E. R. Hilgard, J. V. Neel, and H. A. Simon. Recommendations with respect to the behavioral and social aspects of human genetics. Proceedings of the National Academy of Sciences of the United States of America, 69(??):1–3, ???? 1972. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Tukey:1950:LSF

[TDS50] John W. Tukey, E. H. Dufour, and F. Seibert. Lack of sensitization following repeated skin tests with standard tuberculin (PPD-S). American Review of Tuberculosis, 62(??):77–86, ???? 1950. CODEN ARTUA4. ISSN 0096-0381.

Tukey:1972:TUV

[TEW72] John W. Tukey, R. J. Epp, and G. S. Watson. Testing unit vectors for correlation. Journal of geophysical research, 76(??):8480–8483, ???? 1972. CODEN JGREA2. ISSN 0148-0227 (print), 2156-2202 (electronic).

Tukey:1949:TRA

[TF49] John W. Tukey and M. F. Freeman. Transformation related to the angular and the square-root. SRG Memorandum report 24, Princeton University, Princeton, NJ, USA, 1949. ?? pp.

Tukey:1979:AAD

[TFT79] John W. Tukey, Jerome Friedman, and Paul Tukey. Approaches to analysis of data that concentrate near higher-dimensional manifolds. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1979. ???? pp.

Tukey:20xx:DJT

[TGH^{+}xx] John W. Tukey, Ram Gnanadesikan, David C. Hoaglin, Elizabeth Tukey, and Jon R. Kettenring. A discussion, John Tukey, volume 034 of Distinguished statistician video series. American Statistical Association, Alexandria, VA, USA, 20xx. ISBN ???? LCCN ???? DVD video.

Tukey:1949:MNC

[TH49] John W. Tukey and R. W. Hamming. Measuring noise color. Technical memorandum, Bell Laboratories, Murray Hill, NJ, USA, 1949. 1–127 pp.

Tukey:1981:SSP

[THI81] John W. Tukey, David C. Hoaglin, and Boris Iglewicz. Smallsample performance of a resistant rule for outlier detection. In Proceedings of the Statistical Computing Section, volume 1980, pages 148–152. American Statistical Association, Washington, DC, USA, 1981.

Thomson:1947:AFP

[Tho47] D. Halton Thomson. Approximate formulae for the percentage points of the incomplete beta function and of the χ^{2} distribution. Biometrika, 34(3/4):368–372, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332451.

Thomas:1963:UCS

[Tho63] L. H. Thomas. Using a computer to solve problems in physics. In Applications of Digital Computers, page ?? Ginn and Company, Boston, MA, USA, 1963. LCCN ????

Thode:1986:BRB

[Tho86] Henry C. Thode, Jr. Book review: Exploring Data Tables, Trends, and Shapes by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Technometrics, 28(4):399–400, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268989.

Thompson:2001:AT

[Tho01] James R. Thompson. The age of Tukey. Technometrics, 43(3):256–264, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.allbusiness.com/technology/808082-1.html;   http:/
/www.jstor.org/stable/1271212. Special Tukey Memorial is-

sue.

Thurstone:1946:SPM

[Thu46] L. L. Thurstone. A single plane method of rotation. Psychometrika, 11:71–79, 1946. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Thurstone:1947:MFA

[Thu47] L. L. Thurstone. Multiple-Factor Analysis. A Development and Expansion of “The Vectors of Mind”. The University of Chicago Press, Chicago, Ill., 1947. xix + 535 pp.

Tukey:2000:SFS

[TJ00] John W. Tukey and L. V. Jones. A sensible formulation of the significance test. Psychological Methods, 6(??):17, ???? 2000. CO-DEN ???? ISSN 1082-989x (print), 1939-1463 (electronic).

Tukey:1987:LIT

[TL87] John W. Tukey and Lopressor Intervention Trial Research Group. The Lopressor Intervention Trial: Multicentre study of metoprolol in survivors of acute myocardial infarction. European Heart Journal, 8(??):1056–1064, ???? 1987. CODEN EHJODF. ISSN 0195-668X.

Thompson:1946:TTH

[TM46] Catherine M. Thompson and Maxine Merrington. Tables for testing the homogeneity of a set of estimated variances. Biometrika, 33(4):296–304, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332194. With prefatory note by H. O. Hartley and E. S. Pearson.

Tukey:1947:DMS

[TM47] John W. Tukey and T. Meites. Dipole moments and steroid identification. SRG Memorandum report 2, Princeton University, Princeton, NJ, USA, 1947. 7 pp.

Tukey:1963:LVC

[TM63] John W. Tukey and Donald H. McLaughlin. Less vulnerable confidence and significance procedures for location based on a single sample: Trimming/Winsorization. I. Sankhyā (Indian Journal of Statistics), Series A. Methods and Techniques, 25(??):331–352, 1963. CODEN SANABS. ISSN 0581-572X.

Tukey:1975:TDT

[TM75] John W. Tukey and Max Mathews. Telephone dialing techniques based on move-then-take control. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1975. ???? pp.

Tukey:1980:OTD

[TM80] John W. Tukey and Colin Mallows. An overview of techniques of data analysis, emphasizing its exploratory aspects. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1980. ???? pp.

Tukey:1982:CRS

[TM82] John W. Tukey and Frederick Mosteller. Combination of results of stated precision. I. The optimistic case. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1982. ???? pp.

Tukey:1991:CEA

[TMH91] John W. Tukey, F. Mosteller, and D. C. Hoaglin. Concepts and examples in analysis of variance (ANOVA). In D. C. Hoaglin, F. Mosteller, and J. W. Tukey, editors, Fundamentals of Exploratory Analysis of Variance, pages 1–23. Wiley, New York, NY, USA, 1991.

Tukey:1976:VBP

[TML76] John W. Tukey, R. McGill, and W. Larsen. Variations of box plots. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1976. ???? pp.

Tukey:1978:SPC

[To78] John W. Tukey and 95 others: The Anturane Reinfarction Trial Research Group. Sulfinpyrazone in the prevention of cardiac death after myocardial infarction. The New England Journal of Medicine, 298(??):289–295, ???? 1978. CODEN NEJMAG. ISSN 0028-4793 (print), 1533-4406 (electronic).

Tolstov:1947:EDS

[Tol47] Yu. G. Tolstov. An electrical device for the solution of homogeneous and inhomogeneous ordinary linear differential equations of higher order with constant coefficients, giving the solution in the form of a Taylor series. Bull. Acad. Sci. URSS. Cl. Sci. Tech. [Izvestia Akad. Nauk SSSR], 1947:319–322, 1947.

Toranzos:1939:SJC

[Tor39] Fausto I. Toranzos. On the singularities of Jordan curves. Union Mat. Argentina, Publ., 1939(13):8, 1939.

Thielman:1940:PSA

[TP40] H. P. Thielman and E. S. Pondiczery. Problems and solutions: Advanced problems: 3836. American Mathematical Monthly, 47 (2):115–117, February 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [?, ?].

Tukey:1998:MAI

[TP98a] John W. Tukey and Jan O. Pedersen. Method and apparatus for information access employing overlapping clusters. US

Patent 5,787,422., 1998. URL http://www.google.com/patents/

US5787422. Filed 11 January 1996. Issued 28 July 1998.

Tukey:1998:MOD

[TP98b] John W. Tukey and Jan O. Pedersen. Method of ordering document clusters without requiring knowledge of user interests. US

Patent 5,787,420., 1998. URL http://www.google.com/patents/

US5787420. Filed 14 December 1995. Issued 28 July 1998.

Tukey:1999:MAI

[TP99a] John W. Tukey and Jan O. Pedersen. Method and apparatus for information access employing overlapping clusters. US

Patent 5,999,927., 1999. URL http://www.google.com/patents/

US5999927. Filed 24 April 1998. Issued 7 December 1999.

Tukey:1999:MOD

[TP99b] John W. Tukey and Jan O. Pedersen. Method of ordering document clusters given some knowledge of user interests. US

Patent 5,911,140., 1999. URL http://www.google.com/patents/

US5911140. Filed 14 December 1995, issued 8 June 1999.

Tukey:1950:ABF

[TS50] John W. Tukey and H. Scheffé. Another beta-function approximation. SRG Memorandum report 28, Princeton University, Princeton, NJ, USA, 1950. 18 pp.

Tukey:1976:HEE

[TS76] John W. Tukey and Committee on Impacts of Stratospheric Change, Assembly of Mathematical and National Research Council Physical Sciences. Halocarbons: environmental effects of chlorofluoromethane release. National Academy Press, Washington, DC, USA, 1976. ISBN 0-309-02529-X. ix + 125 pp. LCCN TD885.5.C5 N37 1976.

Trotter:1956:CMC

[TT56] Hale F. Trotter and John W. Tukey. Conditional Monte Carlo for normal samples. In Meyer [Mey56], pages 64–79. LCCN QA273 .F67.

Tompkins:1966:PCS

[TT66] M. L. Tompkins and John W. Tukey. Permuted (circularly-shifted) indexes to abbreviations: a mechanically prepared aid to serial identification. In Proceedings of the 29th Annual Meeting of ADI, pages 347–355. American Documentation Institute, Washington, DC, USA, 1966.

Tukey:1981:DDV

[TT81a] P. A. Tukey and J. W. Tukey. Data-driven view selection: Agglomeration and sharpening. In Barnett [Bar81], pages 215–243. ISBN 0-471-28039-9. LCCN QA278 .I59.

Tukey:1981:GDD

[TT81b] P. A. Tukey and J. W. Tukey. Graphical display of data in three and higher dimensions. In Barnett [Bar81], pages 189–275. ISBN 0-471-28039-9. LCCN QA278 .I59.

Tukey:1981:PPS

[TT81c] P. A. Tukey and J. W. Tukey. Preparation; prechosen sequences of views. In Barnett [Bar81], pages 189–213. ISBN 0-471-28039-9. LCCN QA278 .I59.

Tukey:1981:SSS

[TT81d] P. A. Tukey and J. W. Tukey. Summarization, smoothing, supplemented views. In Barnett [Bar81], pages 245–275. ISBN 0-471-28039-9. LCCN QA278 .I59.

Tukey:1982:SGSb

[TT82] John W. Tukey and Paul A. Tukey. Some graphics for studying four-dimensional data. In Karl W. Heiner, Richard S. Sacher, and John W. Wilkinson, editors, Computer Science and Statistics: Proceedings of the 14th Symposium on the Interface, held July 5-7, 1982, at Rensselaer Polytechnic Institute, Troy, New York, pages 60–66. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90835-8, 3-540-90835-8. LCCN QA276.4 .C58 1982.

Tukey:1985:CGE

[TT85] J. W. Tukey and P. A. Tukey. Computer graphics and exploratory data analysis: An introduction. In Proceedings of the Sixth Annual Conference and Exposition: Computer Graphics ’85, Volume III, Technical Sessions, pages 773–785. National Computer Graphics Association, Fairfax, VA, USA, 1985.

Tukey:1990:SRC

[TT90] John W. Tukey and Paul A. Tukey. Studentized range with covariance: Experimental sampling and Monte Carlo. In collected papers in American Philosophical Society archive., May 29, 1990.

Tukey:1937:MIP

[Tuk37] John Wilder Tukey. Methods for ionization potentials. Thesis (Sc.M.), Department of Chemistry, Brown University, Providence,

RI, USA, 1937. iv + 30 pp. URL http://books.google.com/
books?id=ltI0GwAACAAJ.

Tukey:1938:DFP

[Tuk38] John W. Tukey. On the distribution of the fractional part of a statistical variable. Matematicheskii sbornik, 4(??):561–562, ???? 1938. CODEN ???? ISSN 1064-5616 (print), 1468-4802 (electronic).

Tukey:1939:IMP

[Tuk39a] John W. Tukey. The intrinsic metric of a polytope. Proceedings of the National Academy of Sciences of the United States of America, 25(1):51, January 15, 1939. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Tukey:1939:DT

[Tuk39b] John Wilder Tukey. On Denumerability in Topology. Thesis (Ph.D.), Department of Mathematics, Princeton University, Princeton, NJ, USA, 1939. 158 pp. URL http:/

/gateway.proquest.com/openurl?url_ver=Z39.88-2004&rft_
val_fmt=info:ofi/fmt:kev:mtx:dissertation&res_dat=xri:
pqdiss&rft_dat=xri:pqdiss:0003056.

Tukey:1940:CUT

[Tuk40] John W. Tukey. Convergence and Uniformity in Topology. Annals of Mathematics Studies, no. 2. Princeton University Press, Princeton, NJ, USA, 1940. ix + 90 pp. LCCN QA611 .T85. Republication of doctoral thesis [Tuk39b].

Tukey:1941:CGS

[Tuk41] John W. Tukey. Compactness in general spaces. In R. L. Wilder and W. L. Ayres, editors, Lectures in Topology, pages 307–309. University of Michigan Press, Ann Arbor, MI, USA, 1941.

Tukey:1942:SNS

[Tuk42] J. W. Tukey. Some notes on the separation of convex sets. Portugaliae Mathematica, 3(??):95–102, 1942. CODEN POMAAJ. ISSN 0032-5155.

Tukey:1946:IDM

[Tuk46] John W. Tukey. An inequality for deviations from medians. Annals of Mathematical Statistics, 17(1):75–78, ???? 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731025.

Tukey:1947:JMF

[Tuk47a] J. W. Tukey. The January meeting on fluid dynamics. Bulletin of the American Mathematical Society, 53(7):712–713, 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL

http://projecteuclid.org/euclid.bams/1183510861.

Tukey:1947:RPM

[Tuk47b] J. W. Tukey. Recent publications: Multiple Factor Analysis, by L. L. Thurstone. American Mathematical Monthly, 54(10 (Part 1)):613–615, December 1947. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.

org/stable/2304512.

Tukey:1947:DSP

[Tuk47c] John W. Tukey. Discussion of symposium papers (symposium on statistical methods in experimental and industrial chemistry). Analytical Chemistry (Washington, DC, USA), 19(??):956–957, ???? 1947. CODEN ANCHAM. ISSN 0003-2700 (print), 1520-6882 (electronic).

Tukey:1947:LSS

[Tuk47d] John W. Tukey. Linearization of solutions in supersonic flow. Quarterly of Applied Mathematics, 5(??):361–365, 1947. CODEN QAMAAY. ISSN 0033-569x (print), 1552-4485 (electronic).

Tukey:1947:NPE

[Tuk47e] John W. Tukey. Non-parametric estimation. II. Statistically equivalent blocks and tolerance regions–the continuous case. Annals of Mathematical Statistics, 18(4):529–539, ???? 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Tukey:1947:QIO

[Tuk47f] John W. Tukey. Question 4: Infinite observations. The American Statistician, 1(3):17–18, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682536.

Tukey:1947:RMF

[Tuk47g] John W. Tukey. Review: Multiple Factor Analysis, by L. L. Thurstone. American Mathematical Monthly, 54(10):613–615, December 1947. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tukey:1947:RBT

[Tuk47h] John W. Tukey. Review of A Teoria da Indução Estadistica, by J. Kingston. Journal of the American Statistical Association, 42 (237):190–191, March 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280205.

Tukey:1947:RBF

[Tuk47i] John W. Tukey. Review of The Factorial Analysis of Human Ability, Second Edition, by Godfrey H. Thomson. Journal of the American Statistical Association, 42(237):199–200, March 1947. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280209.

Tukey:1947:RSA

[Tuk47j] John W. Tukey. Review of Statistical Analysis in Biology, 2nd ed, by K. Mather. Quarterly of Applied Mathematics, 5(??):367–368, ???? 1947. CODEN QAMAAY. ISSN 0033-569x (print), 1552-4485 (electronic).

Tukey:1947:RSR

[Tuk47k] John W. Tukey. Review: Statistical Research Group [Columbia University, for the Office of Scientific Research and Development], Sequential Analysis of Statistical Data: Applications. Annals of Mathematical Statistics, 18(1):142–144, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730505.

Tukey:1947:TTM

[Tuk47l] John W. Tukey. Tuberculin tests of the Mercy Hospital nurses. SRG Memorandum report 1, Princeton University, Princeton, NJ, USA, 1947. 7 pp.

Tukey:1948:AW

[Tuk48a] John W. Tukey. Approximate weights. Annals of Mathematical Statistics, 19(1):91–92, ???? 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730297.

Tukey:1948:AME

[Tuk48b] John W. Tukey. Asymptotic moments and expectations. SRG Memorandum report 4, Princeton University, Princeton, NJ, USA, 1948. 44 pp.

Tukey:1948:BVP

[Tuk48c] John W. Tukey. A biometric view of the problem of diagnosis. SRG Memorandum report 3, Princeton University, Princeton, NJ, USA, 1948. 10 pp.

Tukey:1948:CIM

[Tuk48d] John W. Tukey. Comparing individual means in the analysis of variance. SRG Memorandum report 6, Princeton University, Princeton, NJ, USA, 1948. 24 pp.

Tukey:1948:EUA

[Tuk48e] John W. Tukey. Equalizing unequal accuracies of means. SRG Memorandum report 8, Princeton University, Princeton, NJ, USA, 1948. 10 pp.

Tukey:1948:NEI

[Tuk48f] John W. Tukey. Nonparametric estimation. III. Statistically equivalent blocks and multivariate tolerance regions–the discontinuous case. Annals of Mathematical Statistics, 19(1):30–39, ???? 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Tukey:1948:NSR

[Tuk48g] John W. Tukey. A note on the square-root iteration. SRG Memorandum report 10, Princeton University, Princeton, NJ, USA, 1948. 18 pp.

Tukey:1948:PB

[Tuk48h] John W. Tukey. A problem of Berkson. SRG Memorandum report 11, Princeton University, Princeton, NJ, USA, 1948. 5 pp.

Tukey:1948:SLM

[Tuk48i] John W. Tukey. Significance levels for Mosteller’s K-sample slippage test. SRG Memorandum report 9, Princeton University, Princeton, NJ, USA, 1948. 7 pp.

Tukey:1948:SEPa

[Tuk48j] John W. Tukey. Some elementary problems of importance to small sample practice. SRG Memorandum report 5, Princeton University, Princeton, NJ, USA, 1948. 10 pp.

Tukey:1948:SEPb

[Tuk48k] John W. Tukey. Some elementary problems of importance to small sample practice. Human biology, 20(??):205–214, ???? 1948. CO-DEN HUBIAA. ISSN 0018-7143 (print), 1534-6617 (electronic).

Tukey:1948:STS

[Tuk48l] John W. Tukey. Sufficiency, truncation and selection. SRG Memorandum report 12, Princeton University, Princeton, NJ, USA, 1948. 3 pp.

Tukey:1949:AQM

[Tuk49a] John W. Tukey. Answer to question 21: Maximum likelihood estimates, efficiency. The American Statistician, 3(4):12, October 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681357.

Tukey:1949:AOC

[Tuk49b] John W. Tukey. The approximate operating characteristics of combined variables and attribute sampling plans. SRG Memorandum report 22, Princeton University, Princeton, NJ, USA, 1949. 12 pp.

Tukey:1949:CIM

[Tuk49c] John W. Tukey. Comparing individual means in the analysis of variance. Biometrics, 5(??):99–114, 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1949:CPA

[Tuk49d] John W. Tukey. Comparing patterns by analyses of variance techniques. SRG Memorandum report 21, Princeton University, Princeton, NJ, USA, 1949. 14 pp.

Tukey:1949:CLP

[Tuk49e] John W. Tukey. Confidence limits for parabola vertices and mence for symmetrical minima and maxima. SRG Memorandum report 23, Princeton University, Princeton, NJ, USA, 1949. 10 pp.

Tukey:1949:DLR

[Tuk49f] John W. Tukey. Determination of linear relations between systematic parts. SRG Memorandum report 19, Princeton University, Princeton, NJ, USA, 1949. 12 pp.

Tukey:1949:DAA

[Tuk49g] John W. Tukey. Dyadic ANOVA: An analysis of variance for vectors. Human biology, 21(??):65–110, ???? 1949. CODEN HU-BIAA. ISSN 0018-7143 (print), 1534-6617 (electronic).

Tukey:1949:EAP

[Tuk49h] John W. Tukey. An empirical approximation to the 5% point of Fisher’s B. SRG Memorandum report 27, Princeton University, Princeton, NJ, USA, 1949. 2 pp.

Tukey:1949:FNP

[Tuk49i] John W. Tukey. Finite normal populations. SRG Memorandum report 35, Princeton University, Princeton, NJ, USA, 1949. 1 pp.

Tukey:1949:IRC

[Tuk49j] John W. Tukey. Interaction in a row-by-column design. SRG Memorandum report 18, Princeton University, Princeton, NJ, USA, 1949. 14 pp.

Tukey:1949:MSFa

[Tuk49k] John W. Tukey. Memorandum on statistics in the Federal Government, Part I. The American Statistician, 3(1):6–17, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681391.

Tukey:1949:MSFb

[Tuk49l] John W. Tukey. Memorandum on statistics in the Federal Government, Part II. The American Statistician, 3(2):12–16, April/ May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682739.

Tukey:1949:MRGa

[Tuk49m] John W. Tukey. Moments of random group size distributions. SRG Memorandum report 15, Princeton University, Princeton, NJ, USA, 1949. 26 pp.

Tukey:1949:MRGb

[Tuk49n] John W. Tukey. Moments of random group size distributions. Annals of Mathematical Statistics, 20(4):523–539, ???? 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729945.

Tukey:1949:NT

[Tuk49o] John W. Tukey. A nonuniqueness theorem. SRG Memorandum report 38, Princeton University, Princeton, NJ, USA, 1949. 2 pp.

Tukey:1949:ODF

[Tuk49p] John W. Tukey. One degree of freedom for non-additivity. Biometrics, 5(??):232–242, ???? 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://www.jstor. org/stable/3001938. Reprinted in [Cox92, pp. 1–13].

Tukey:1949:PDR

[Tuk49q] John W. Tukey. A problem in the distribution of rankings. SRG Memorandum report 40, Princeton University, Princeton, NJ, USA, 1949. 5 pp.

Tukey:1949:RBT

[Tuk49r] John W. Tukey. Review of Probability Theory for Statistical Methods, by F. N. David. Journal of the American Statistical Association, 44(248):567–568, December 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2279909.

Tukey:1949:RBE

[Tuk49s] John W. Tukey. Review of Theory of Experimental Inference, by C. W. Churchman. Journal of the American Statistical Association, 44(245):136–139, March 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280360.

Tukey:1949:SPE

[Tuk49t] John W. Tukey. Scaling by and for percentiles and exponential averages. SRG Memorandum report 33, Princeton University, Princeton, NJ, USA, 1949. 47 pp.

Tukey:1949:SSR

[Tuk49u] John W. Tukey. The simplest signed-rank tests. SRG Memoranda Report 17, Princeton University, Princeton, NJ, USA, 1949. ?? pp.

Tukey:1949:STR

[Tuk49v] John W. Tukey. Skeleton tables related to contaminated distributions at scale 3. SRG Memorandum report 34, Princeton University, Princeton, NJ, USA, 1949. 5 pp.

Tukey:1949:SCP

[Tuk49w] John W. Tukey. Standard confidence points. SRG Memorandum report 26, Princeton University, Princeton, NJ, USA, 1949. 37 pp.

Tukey:1949:STS

[Tuk49x] John W. Tukey. Sufficiency, truncation and selection. Annals of Mathematical Statistics, 20(2):309–311, ???? 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730042.

Tukey:1949:TMM

[Tuk49y] John W. Tukey. The truncated mean in moderately large samples. SRG Memorandum report 32, Princeton University, Princeton, NJ, USA, 1949. 10 pp.

Tukey:1949:UPN

[Tuk49z] John W. Tukey. The upper 5% point of the normal range. SRG Memorandum report 30, Princeton University, Princeton, NJ, USA, 1949. 1 pp.

Tukey:1950:ACR

[Tuk50a] John W. Tukey. Another comparison of range and standard deviation. SRG Memorandum report 49, Princeton University, Princeton, NJ, USA, 1950. 3 pp.

Tukey:1950:CCA

[Tuk50b] John W. Tukey. Chairman’s closure. Acceptance sampling. A symposium. Journal of the American Statistical Association, 45(250): 150–155, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1950:DSS

[Tuk50c] John W. Tukey. Discussion [symposium: Statistics for the clinician]. Journal of Clinical Psychology, 6(??):61–74, ???? 1950. CO-DEN JCPYAO. ISSN 0021-9762 (print), 1097-4679 (electronic).

Tukey:1950:STP

[Tuk50d] John W. Tukey. The sampling theory of power spectrum estimates. In Symposium on Applications of Autocorrelation Analysis to Physical Problems (NAVEXOSP-735), pages 47–67. Office of Naval Research, Washington, DC, USA, 1950. See also ASnumber 64.

Tukey:1950:SSS

[Tuk50e] John W. Tukey. Some sampling simplified. Journal of the American Statistical Association, 45(252):501–519, December 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280719.

Tukey:1950:WDM

[Tuk50f] John W. Tukey. What does the mean-square-successive-difference measure when applied to time series? In collected papers in American Philosophical Society archive., September 28, 1950.

Tukey:1951:CR

[Tuk51a] John W. Tukey. Components in regression. Biometrics, 7(??): 33–69, ???? 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1951:FSS

[Tuk51b] John W. Tukey. Finite sampling simplified: polykays, cumulants of variances, and cumulants of variance components. SRG Memorandum report 45, Princeton University, Princeton, NJ, USA, 1951. 112 pp.

Tukey:1951:QDM

[Tuk51c] John W. Tukey. Quick and dirty methods in statistics — Part II. Simple analyses for standard designs. In Quality Control Conference Papers 1951, pages 189–197. American Society for Quality Control, New York, NY, USA, May 1951. URL http://asq.org/

qic/display-item/?item=48.

Tukey:1951:SMA

[Tuk51d] John W. Tukey. Standard methods of analyzing data. In Proceedings of a Computation Seminar [IBM Department of Education, Endicot, NY, from December 5 to 9, 1949], pages 95–112. IBM Corporation, San Jose, CA, USA, 1951.

Tukey:1952:EAF

[Tuk52a] John W. Tukey. Estimation in the alternative family of distributions. In Proceedings of the International Congress of Mathematicians 1951, page 586. ????, ????, 1952.

Tukey:1952:RBE

[Tuk52b] John W. Tukey. Review of The Extrapolation, Interpolation and Smoothing of Stationary Time Series with Engineering Applications, by Norbert Wiener. Journal of the American Statistical Association, 47(258):319–321, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280758.

Tukey:1952:RSMa

[Tuk52c] John W. Tukey. Review of Statistical Methods for Research Workers, 11th ed, by R. A. Fisher. Econometrica, 20(??):511–512, ???? 1952. CODEN ECMTA7. ISSN 0012-9682.

Tukey:1953:GED

[Tuk53a] John W. Tukey. The growth of experimental design in a research laboratory. In Research Operations in Industry, pages 303–313. King’s Crown Press, New York, NY, USA, 1953.

Tukey:1953:PMC

[Tuk53b] John W. Tukey. The problem of multiple comparisons. This manuscript was intended for a book but was never formally published [McC03, page 549], although the manuscript was widely circulated. It was finally published in the final volume of Tukey’s Collected Works [Bra94, pp. 1–300]., 1953.

Tukey:1953:SSQ

[Tuk53c] John W. Tukey. Some selected quick and easy methods of statistical analysis. Trans. New York Acad. Sci. (2), 16(??):88–97, 1953. CODEN TNYAAE. ISSN 0028-7113 (print), 1939-2966 (electronic).

Tukey:1953:SRT

[Tuk53d] John W. Tukey. The spectral representation and transformation properties of the higher moments of stationary time series. ???? ??, ????, ????, 1953. Reprinted in [Bri84a, pp. 165–184].

Tukey:1954:CRP

[Tuk54a] John W. Tukey. Causation, regression and path analysis. In T. O. Kempthorne, A. Bancroft, J. W. Gowen, and J. L. Lush, editors, Statistics and Mathematics in Biology, pages 35–66. State College Press, Ames, IA, USA, 1954.

Tukey:1954:CTS

[Tuk54b] John W. Tukey. Comparing two small samples on many items. SRG Memorandum report 54, Princeton University, Princeton, NJ, USA, 1954. 23 pp.

Tukey:1954:NAI

[Tuk54c] John W. Tukey. Noted on the analysis of the “immensed” variance. Report RM-1317, Rand Corporation, ????, 1954.

Tukey:1954:UPE

[Tuk54d] John W. Tukey. Unsolved problems of experimental statistics. Journal of the American Statistical Association, 49(268):706–731, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281535.

Tukey:1955:MCC

[Tuk55a] J. W. Tukey. Mathematical consultants, computational mathematics and mathematical engineering. American Mathematical Monthly, 62(8):565–571, October 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2307249.

Tukey:1955:AQ

[Tuk55b] John W. Tukey. Answer to Query 113. Biometrics, 11(??):111–113, ???? 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1955:IAR

[Tuk55c] John W. Tukey. Interpolations and approximations related to the normal range. Biometrika, 42(3/4):480–485, December 1955. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333393.

Tukey:1955:RMR

[Tuk55d] John W. Tukey. Review of A Million Random Digits with 100,000 Normal Deviates, by the Rand Corporation. Journal of the Operations Research Society of America, 3(??):568–571, ???? 1955. CODEN JORSAW. ISSN 0096-3984.

Tukey:1955:SJP

[Tuk55e] John W. Tukey. Systematization of journal practices. Science, 122 (3162):246, August 5, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Tukey:1956:KML

[Tuk56a] John W. Tukey. Keeping moment-like sampling computations simple. Annals of Mathematical Statistics, 27(1):37–54, ????

1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://projecteuclid.org/euclid.aoms/
1177728349.

Tukey:1956:SIT

[Tuk56b] John W. Tukey. A smooth invertibility theory. SRG Memorandum report 57, Princeton University, Princeton, NJ, USA, 1956. 6 pp.

Tukey:1956:VVC

[Tuk56c] John W. Tukey. Variances of variance components. I. Balanced designs. Annals of Mathematical Statistics, 27(3):722–736, ???? 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://projecteuclid.org/euclid.aoms/
1177728179.

Tukey:1957:AFR

[Tuk57a] John W. Tukey. Antiquadrature formulas as routes to less biased estimate. STRG Technical report 1, Statistical Techniques Research Group, Princeton University, Princeton, NJ, USA, 1957. 17 pp.

Tukey:1957:AR

[Tuk57b] John W. Tukey. Antithesis or regression? Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 53(??):923–924, 1957. CODEN PCPSA4. ISSN 0008-1981.

Tukey:1957:AUP

[Tuk57c] John W. Tukey. Approximations to the upper 5% points of Fisher’s B distribution and non-central χ^{2}. Biometrika, 44(3/4):528–530, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332885.

Tukey:1957:CAT

[Tuk57d] John W. Tukey. On the comparative anatomy of transformations. Annals of Mathematical Statistics, 28(3):602–632, ???? 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706875.

Tukey:1957:PEF

[Tuk57e] John W. Tukey. Propagation of errors, fluctuations and tolerances. I. Basic generalized formulas. STRG Technical report 10, Princeton University, Princeton, NJ, USA, 1957. 138 pp.

Tukey:1957:STP

[Tuk57f] John W. Tukey. The sampling theory of power spectrum estimates. Journal of cycle research, 6(??):31–52, ???? 1957. CODEN ???? ISSN ????

Tukey:1957:SEF

[Tuk57g] John W. Tukey. Some examples with fiducial relevance. Annals of Mathematical Statistics, 28(3):687–695, ???? 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706880.

Tukey:1957:SRP

[Tuk57h] John W. Tukey. Sums of random partitions of ranks. Annals of Mathematical Statistics, 28(4):987–992, ???? 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706799.

Tukey:1957:VVCa

[Tuk57i] John W. Tukey. Variances of variance components. II. The unbalanced single classification. Annals of Mathematical Statistics, 28(1):43–56, ???? 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177707036.

Tukey:1957:VVCb

[Tuk57j] John W. Tukey. Variances of variance components. III. Third moments in a balanced single classification. Annals of Mathematical Statistics, 28(2):378–384, ???? 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706965.

Tukey:1958:BCQ

[Tuk58a] John W. Tukey. Bias and confidence in not-quite large samples (abstract). Annals of Mathematical Statistics, 29(2):614, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://projecteuclid.org/euclid.aoms/

1177706647. This is the other half of the Quenouille–Tukey jackknife technique; see [Que49, Que56, Mil74, Efr79].

Tukey:1958:PBM

[Tuk58b] John W. Tukey. A problem of Berkson, and minimum variance orderly estimators. Annals of Mathematical Statistics, 29 (2):588–592, ???? 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706637.

Tukey:1958:PEF

[Tuk58c] John W. Tukey. Propagation of errors, fluctuations and tolerances. 3 — an exercise in partial differentiation. Technical report 12, Princeton University, Princeton, NJ, USA, 1958. 72 pp.

Tukey:1958:SIT

[Tuk58d] John W. Tukey. A smooth invertibility theorem. Annals of Mathematical Statistics, 29(2):581–584, ???? 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706635.

Tukey:1958:TCM

[Tuk58e] John W. Tukey. The teaching of concrete mathematics. American Mathematical Monthly, 65(1):1–9, January 1958. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). This article is believed to contain the first published instance of the word ‘software’ in the meaning of instructions to a computer: “Today the ‘software’ comprising the carefully planned interpretive routines, compilers, and other aspects of automative programming are at least as important to the modern electronic calculator as its ‘hardware’ of tubes, transistors, wires, tapes and the like.” [page 2].

Tukey:1959:AEP

[Tuk59a] John W. Tukey. Appendix 9: Equalization and pulse shaping techniques applied to the determination of initial sense of Rayleigh waves: The need for fundamental research in seismology. Report, Report of the Panel on Seismic Improvement, US State Department, Washington, DC, USA, 1959.

Tukey:1959:EPSb

[Tuk59b] John W. Tukey. Equalization and pulse shaping techniques applied to the determination of initial sense of Rayleigh waves. In The Need for Fundamental Research in Seismology, pages 60–129. Report of the Panel on Seismic Improvement, U.S. State Department, Washington, DC, USA, 1959. LCCN ???? Appendix 9. Reprinted in [Bri84a, pp. 279–307].

Tukey:1959:EPSc

[Tuk59c] John W. Tukey. Equalization and pulse shaping techniques applied to the determination of initial sense of Rayleigh waves. STRG Technical report 34, Princeton University, Princeton, NJ, USA, 1959. 129 pp.

Tukey:1959:EPSa

[Tuk59d] John W. Tukey. The estimation of (power) spectra and related quantities. In Langer [Lan59], pages 389–411. LCCN QA3 .U45 no. 1. Publication no. 1 of the Mathematics Research Center, U.S. Army, the University of Wisconsin.

Tukey:1959:IMS

[Tuk59e] John W. Tukey. An introduction to the measurement of spectra. In Ulf Grenander, editor, Probability and Statistics, The Harald Cramér Volume, pages 300–330. Almqvist and Wiksell, Stockholm, Sweden, 1959.

Tukey:1959:LPM

[Tuk59f] John W. Tukey. Little pieces of mixed factorials (abstract 627). Biometrics, 15(??):641–642, ???? 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1959:QCT

[Tuk59g] John W. Tukey. A quick, compact, two-sample test to Duckworth’s specifications. Technometrics, 1(1):31–48, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266308.

Tukey:1959:SQM

[Tuk59h] John W. Tukey. Statistical and quantitative methodology. STRG Technical report 31, Princeton University, Princeton, NJ, USA, 1959. 62 pp.

Tukey:1959:SSC

[Tuk59i] John W. Tukey. A survey of sampling from contaminated distributions. STRG Technical report 33, Princeton University, Princeton, NJ, USA, 1959. 57 pp.

Tukey:1960:CVD

[Tuk60a] John W. Tukey. Conclusions vs. decisions. Technometrics, 2(4): 423–433, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266451.

Tukey:1960:CPT

[Tuk60b] John W. Tukey. Curves as parameters, and touch estimation. STRG Technical report 39, Princeton University, Princeton, NJ, USA, 1960. 30 pp.

Tukey:1960:PRB

[Tuk60c] John W. Tukey. The practical relationship between the common transformations of percentages or fractions and of amounts. STRG Technical report 36, Princeton University, Princeton, NJ, USA, 1960. 6 pp.

Tukey:1960:SSC

[Tuk60d] John W. Tukey. A survey of sampling from contaminated distributions. In I. Olkin, S. G. Ghurye, W. Hoeffding, W. G. Madow, and H. B. Mann, editors, Contributions to probability and statistics: Essays in Honor of Harold Hotelling, pages 448–485. Stanford University Press, Stanford, CA, USA, 1960.

Tukey:1960:WDW

[Tuk60e] John W. Tukey. Where do we go from here? Journal of the American Statistical Association, 55(289):80–93, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282180.

Tukey:1961:CPT

[Tuk61a] John W. Tukey. Curves as parameters, and touch estimation. In Proceedings 4th Berkeley Symposium Math. Statist. and Prob., Volume I, pages 681–694. University of California Press, Berke-

ley, CA, USA, 1961. URL http://projecteuclid.org/euclid.
bsmsp/1200512189.

Tukey:1961:DECa

[Tuk61b] John W. Tukey. Discussion, emphasizing the connection between analysis of variance and spectrum analysis. STRG Technical report 41, Princeton University, Princeton, NJ, USA, 1961. 68 pp.

Tukey:1961:DECb

[Tuk61c] John W. Tukey. Discussion, emphasizing the connection between analysis of variance and spectrum analysis. Technometrics, 3(2): 191–219, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1266112.

Tukey:1961:FDA

[Tuk61d] John W. Tukey. The future of data analysis. STRG Technical report 43, Princeton University, Princeton, NJ, USA, 1961. 99 + 3 pp.

Tukey:1961:MDP

[Tuk61e] John W. Tukey. Microwave data processing circuits. US

Patent 3,007,643., 1961. URL http://www.google.com/patents/

US3007643. Filed 31 December 1956. Issued 7 November 1961. Tukey:1961:SQM

[Tuk61f] John W. Tukey. Statistical and quantitative methodology. In D. P. Ray, editor, Trends in Social Science, pages 84–136. Philosophical Library, New York, NY, USA, 1961.

Tukey:1962:CII

[Tuk62a] John W. Tukey. The citation index and the information problem opportunities and research in progress: annual report for 1962. Report, Statistical Techniques Research Group, Princeton University, Princeton, NJ, USA, 1962. 58 pp.

Tukey:1962:CNC

[Tuk62b] John W. Tukey. Correction notes: Correction to The future of data analysis. Annals of Mathematical Statistics, 33(2):812, ???? 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://projecteuclid.org/euclid.aoms/
1177704604.

Tukey:1962:FDA

[Tuk62c] John W. Tukey. The future of data analysis. Annals of Mathematical Statistics, 33(1):1–67, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704604;
http://www.jstor.org/stable/2237638. See 40-year retrospec-

tive discussion in [Mal06] and companion papers in that journal issue (Technometrics 48(3) August 2006). See also the 50-year view in [Don17].

Tukey:1962:KRCa

[Tuk62d] John W. Tukey. Keeping research in contact with literature: Citation indices and beyond. Transactions I.R.E. Professional Group on Engineering Writing and Speech, 5(2):78–82, ???? 1962. CO-DEN ???? ISSN ????

Tukey:1962:KRCb

[Tuk62e] John W. Tukey. Keeping research in contact with literature: Citation indices and beyond. Journal of Chemical Documentation, 2 (1):34–37, ???? 1962. CODEN JCHDAN. ISSN 0021-9576.

Tukey:1962:SD

[Tuk62f] John W. Tukey. The symmetrical λ distributions. Cited as ‘in preparation’ in [Tuk62c, page 67]. See also [JR71, JK73]., 1962.

Tukey:1963:AD

[Tuk63a] John W. Tukey. The analysis of data. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1963. ???? pp.

Tukey:1963:CIS

[Tuk63b] John W. Tukey. A citation index for statistics and probability. Bulletin of the International Statistical Institute, 40(??):747–756, ???? 1963. CODEN ???? ISSN 0373-0441.

Tukey:1963:DAB

[Tuk63c] John W. Tukey. Data analysis and behavioral science (Chapters D E F G H J K). Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1963. ???? pp.

Tukey:1963:ICB

[Tuk63d] John W. Tukey. The inevitable collision between computation and data analysis. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1963. ???? pp.

Tukey:1963:MIF

[Tuk63e] John W. Tukey. Mathematics 596: An introduction to the frequency analysis of time series. Unpublished class notes., 1963.

Tukey:1963:TSJa

[Tuk63f] John W. Tukey. A tagging system for journal articles and other citable items: a status report. Report, Statistical Techniques Research Group, Princeton University, Princeton, NJ, USA, 1963. 292 pp.

Tukey:1963:TSJb

[Tuk63g] John W. Tukey. A tagging system for journal articles and other citable items: a status report. Annual Report for 1963 under National Science Foundation Grant NSF-GN-297 (from the Office of Science Information Service). Report, National Science Foundation, Washington, DC, USA, 1963.

Tukey:1963:WCD

[Tuk63h] John W. Tukey. What can data analysis and statistics offer today? In Ocean Wave Spectra: Proceedings of a Conference, pages 347– 350. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1963.

Tukey:1963:YM

[Tuk63i] John W. Tukey. You and mathematics. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1963. ???? pp.

Tukey:1964:SSW

[Tuk64a] J. W. Tukey. Samuel Stanley Wilks. In Year Book of the American Philosophical Society, pages 147–154. American Philosophical Society, ????, 1964. LCCN ????

Tukey:1964:TTS

[Tuk64b] John W. Tukey. The technical tools of statistics. Report, Bell Laboratories, Murray Hill, NJ, USA, November 1964. ????

pp. URL http://cm.bell-labs.com/cm/ms/departments/sia/
tukey/memo/techtools.html.

Tukey:1965:ATA

[Tuk65a] John W. Tukey. An approach to thinking about a statistical computing system. Unpublished notes circulated at Bell Laboratories., 1965.

Tukey:1965:DAFa

[Tuk65b] John W. Tukey. Data analysis and the frontiers of geophysics. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1965. ???? pp.

Tukey:1965:DAFb

[Tuk65c] John W. Tukey. Data analysis and the frontiers of geophysics. Science, 148(3675):1283–1289, June 4, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Tukey:1965:FPDa

[Tuk65d] John W. Tukey. The future of processes of data analysis. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1965. ???? pp.

Tukey:1965:FPDb

[Tuk65e] John W. Tukey. The future of processes of data analysis. In ????, editor, Proceedings of the Tenth Conference on Design of Experiments in Army Research, Development and Testing, pages 691–729. Army Research Office, Durham, NC, USA, 1965. LCCN UF526.3.A36.

Tukey:1965:ICB

[Tuk65f] John W. Tukey. The inevitable collision between computation and data analysis. In Proceedings of the IBM Scientific Computing Symposium on Statistics October 21–23, 1963, pages 141–152. IBM Corporation, San Jose, CA, USA, 1965.

Tukey:1965:M

[Tuk65g] John W. Tukey. Mathematics. In A. Love and J. S. Childers, editors, Listen to Leaders in Science, pages 219–232. Tupper and Love, Atlanta, GA, USA, 1965. Reprinted in AT&T Bell Laboratories booklet entitled Dimensions in Communications Engineering, Science, Mathematics.

Tukey:1965:TTSa

[Tuk65h] John W. Tukey. The technical tools of statistics. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1965. ???? pp.

Tukey:1965:TTSb

[Tuk65i] John W. Tukey. The technical tools of statistics. The American Statistician, 19(2):23–28, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682374.

Tukey:1965:WPS

[Tuk65j] John W. Tukey. Which part of the sample contains the information? Proceedings of the National Academy of Sciences of the United States of America, 53(??):127–134, 1965. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Tukey:1966:PSL

[Tuk66a] John W. Tukey. A practicing statistician looks at the Transactions. IEEE Transactions on Information Theory, 12(2):87–91, ???? 1966. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Tukey:1966:SP

[Tuk66b] John W. Tukey. Statement on pollution. In Federal Research and Development Programs, Hearings before a Sub-Committee of the Committee on Government Operations, House of Representatives. 89th Congress, Second Session, January 1966, pages 121– 126. United States Government Printing Office, Washington, DC, USA, 1966. LCCN ????

Tukey:1966:UNS

[Tuk66c] John W. Tukey. Use of numerical spectrum analysis in geophysics. Bulletin of the International Statistical Institute, 41(??):267–307, ???? 1966. CODEN ???? ISSN 0373-0441.

Tukey:1967:DS

[Tuk67a] John W. Tukey. Discussion of session 1. SIAM Review, 9(??): 308–310, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tukey:1967:DAP

[Tuk67b] John W. Tukey. Discussion on Anscombe’s paper [topics in the investigation of linear relations fitted by the method of least squares]. Journal of the Royal Statistical Society. Series B (Methodological), 29(??):47–48, ???? 1967. CODEN JSTBAJ. ISSN 0035-9246.

Tukey:1967:ICN

[Tuk67c] John W. Tukey. An introduction to the calculations of numerical spectrum analysis. In B. Harris, editor, Spectral Analysis Time Series (Proceedings Advanced Sem., Madison, Wis., 1966), pages 25–46. Wiley, New York, NY, USA, 1967. Reprinted in [Bri84a, pp. 811–835].

Tukey:1967:SC

[Tuk67d] John W. Tukey. A statistician’s comment. In A. Kent, O. E. Taulbee, J. Belzer, and G. D. Goldstein, editors, Electronic Handling of Information: Testing and Evaluation, pages 41–47. Thompson, Washington, DC, USA, 1967.

Tukey:1967:WCM

[Tuk67e] John W. Tukey. What can mathematicians do for the Federal Government? American Mathematical Monthly, 74(1 (Part 2)):101– 109, January 1967. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). Part 2: Fiftieth Anniversary Issue.

Tukey:1968:SCS

[Tuk68a] John W. Tukey. Is statistics a computing science? In D. G. Watts, editor, The Future of Statistics, pages 19–38. Academic Press, New York, USA, 1968.

Tukey:1968:PWP

[Tuk68b] John W. Tukey. Proceedings of the workshop: Practical applications of the frequency approach to EEG analysis. In Advances in EEG Analysis, Supplement 27 Electro-encephalography and Clinical Neurophysiology, pages 10–11. North-Holland Publishing Co., Amsterdam, The Netherlands, 1968.

Tukey:1968:TPT

[Tuk68c] John W. Tukey. The true purpose of transformation (abstract 1545). Biometrics, 24(??):1041, ???? 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1969:ADS

[Tuk69a] John W. Tukey. Analyzing data: Sanctification or detective work? American Psychologist, 24(2):83–91, ???? 1969. CODEN AMPSAB. ISSN 0003-066x (print), 1935-990x (electronic).

Tukey:1969:AFI

[Tuk69b] John W. Tukey. Appendix 2: a further inquiry into institutional differences by means of superstandardization (a regression adjustment beyond standardization). In The National Halothane Study, a study of the possible association between halothane anesthesia and postoperative hepatic necrosis, pages 358–369. National Institutes of Health, Bethesda, MD, USA, 1969.

Tukey:1970:EDAa

[Tuk70a] John W. Tukey. Exploratory Data Analysis 1. Addison-Wesley, Reading, MA, USA, 1970. ???? pp. LCCN ???? Limited preliminary edition.

Tukey:1970:EDAb

[Tuk70b] John W. Tukey. Exploratory Data Analysis 2. Addison-Wesley, Reading, MA, USA, 1970. ???? pp. LCCN ???? Limited preliminary edition.

Tukey:1970:FSL

[Tuk70c] John W. Tukey. First 1970 Scott lecture. Cited in [Bri02b, page 1618], 1970.

Tukey:1970:SSL

[Tuk70d] John W. Tukey. Second 1970 Scott lecture. Cited in [Bri02b, page 1618], 1970.

Tukey:1970:SFI

[Tuk70e] John W. Tukey. Some further inputs. In Daniel F. Merriam, editor, Geostatistics: A colloquium: Proceedings of a colloquium on geostatistics held on campus at the University of Kansas, Lawrence on 7–9, June 1970, pages 163–174. Plenum Press, New York, NY, USA; London, UK, 1970. ISBN 0-306-30519-4. LCCN QE33 .G47 1970.

Tukey:1970:SGS

[Tuk70f] John W. Tukey. Some graphic and semigraphic displays (abstract). Technometrics, 12(1):205, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267376.

Tukey:1971:DAC

[Tuk71a] John W. Tukey. Data analysis, computation and mathematics. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1971. ???? pp.

Tukey:1971:DSF

[Tuk71b] John W. Tukey. Do statisticians have a future? Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1971. ???? pp.

Tukey:1971:EDA

[Tuk71c] John W. Tukey. Exploratory Data Analysis 3. Addison-Wesley, Reading, MA, USA, 1971. ???? pp. LCCN ???? Limited preliminary edition.

Tukey:1971:HCS

[Tuk71d] John W. Tukey. How computing and statistics affect each other. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1971. ???? pp.

Tukey:1971:LST

[Tuk71e] John W. Tukey. Lags in statistical technology. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1971. ???? pp.

Tukey:1972:DAC

[Tuk72a] John W. Tukey. Data analysis, computation and mathematics. Quarterly of Applied Mathematics, 30(??):51–65, ???? 1972. CO-DEN QAMAAY. ISSN 0033-569x (print), 1552-4485 (electronic). Special Issue: Symposium on the Future of Applied Mathematics.

Tukey:1972:HCS

[Tuk72b] John W. Tukey. How computing and statistics affect each other. In The Babbage Memorial Meeting: Report of Proceedings, pages 21– 37. British Computer Society and Royal Statistical Society, London, UK, 1972.

Tukey:1972:SGS

[Tuk72c] John W. Tukey. Some graphic and semigraphic displays. In Statistical papers in honor of George W. Snedecor, pages 293–316. Iowa State University Press, Ames, Iowa, 1972.

Tukey:1973:DAM

[Tuk73a] John W. Tukey. Data analysis for molecular structure: today and tomorrow. Technical report, series 2 44, Princeton University, Princeton, NJ, USA, 1973. 10 pp.

Tukey:1973:EDA

[Tuk73b] John W. Tukey. Exploratory data analysis as part of a large whole. In Proceedings of the Eighteenth Conference on the Design of Experiments in Army Research, Development and Testing, Part I, pages 1–10. U.S. Army Research Office, Durham, NC, USA, 1973.

Tukey:1973:ISP

[Tuk73c] John W. Tukey. Index to statistics and probability: Citation index, volume 2 of The Information Access Series. R&D Press, Los Altos, CA, USA, 1973. ISBN 0-88274-001-6. xxx + 1269 pp. LCCN Z6654.M33 T84.

Tukey:1973:ITD

[Tuk73d] John W. Tukey. Introduction of today’s data analysis. Technical report, series 2 40, Princeton University, Princeton, NJ, USA, 1973. 24 pp.

Tukey:1973:LST

[Tuk73e] John W. Tukey. Lags in statistical technology. In C. S. Carter et al., editors, Proceedings of the First Canadian Conference in Applied Statistics, pages 96–104. American Statistical Association, Montreal chapter, ????, 1973.

Tukey:1973:LNS

[Tuk73f] John W. Tukey. Lecture notes for statistics 411 — 1973, with appendices. Technical report, Princeton University, Princeton, NJ, USA, 1973. 206 pp.

Tukey:1973:STA

[Tuk73g] John W. Tukey. Some thoughts on alternagraphic displays. Technical report, series 2 48, Princeton University, Princeton, NJ, USA, 1973. 17 pp.

Tukey:1973:ZCI

[Tuk73h] John W. Tukey. The zigzagging climb from initial observation to successful improvement. In W. E. Coffman, editor, Frontiers of Educational Measurement and Information Systems—1973, pages 113–120. Houghton Mifflin, Boston, MA, USA, 1973.

Tukey:1974:DAM

[Tuk74a] John W. Tukey. Data analysis for molecular structure. today and tomorrow. In D. R. Lide, Jr. and M. A. Paul, editors, Critical Evaluation of Chemical and Physical Structural Information, pages 48– 58. National Academy of Sciences, Washington, DC, USA, 1974.

Tukey:1974:FAF

[Tuk74b] John W. Tukey. A further analysis of the first phase of the Princeton Robustness Study: examples of less standard two-way table analysis. In W. J. Dixon and W. L. Nicholson, editors, Exploring Data Analysis: The Computer Revolution in Statistics, pages 229–294. University of California Press, Berkeley, CA, USA, 1974.

Tukey:1974:ITD

[Tuk74c] John W. Tukey. Introduction to today’s data analysis. In D. R. Lide, Jr. and M. A. Paul, editors, Critical Evaluation of Chemical and Physical Structural Information, pages 3–14. National Academy of Sciences, Washington, DC, USA, 1974.

Tukey:1974:NFV

[Tuk74d] John W. Tukey. Named and faceless values: An initial exploration in memory of Prasanta C. Mahalanobis. Sankhyā (Indian Journal of Statistics), Series A. Methods and Techniques, 36(??):125–176, ???? 1974. CODEN SANABS. ISSN 0036-4452.

Tukey:1974:NBH

[Tuk74e] John W. Tukey. New Bedford High School commencement remarks. Unpublished., 1974.

Tukey:1974:SGI

[Tuk74f] John W. Tukey. Statement on growth and its implications for the future. In Hearings before the Subcommittee on Fisheries and Wildlife Conservation and the Environment of the House of Representatives, pages 54–56. United States Government Printing Office, Washington, DC, USA, 1974. LCCN ???? Serial number 93-35.

Tukey:1975:IGM

[Tuk75a] John W. Tukey. Instead of Gauss–Markov least squares, what? In Applied statistics (Proceedings Conf., Dalhousie University, Halifax, NS, 1974), pages 351–372. North-Holland Publishing Co., Amsterdam, The Netherlands, 1975.

Tukey:1975:MPD

[Tuk75b] John W. Tukey. Mathematics and the picturing of data. In Proceedings of the International Congress of Mathematicians (Vancouver, BC, 1974), Volume 2, pages 523–531. Canadian Mathematical Congress, Montréal, Québec, Canada, 1975.

Tukey:1975:MSR

[Tuk75c] John W. Tukey. Methodology, and statisticians’ responsibility for both accuracy and relevance. Technical report, series 2 101, Bell Laboratories, Murray Hill, NJ, USA, 1975. 14 pp.

Tukey:1976:RSN

[Tuk76a] J. W. Tukey. Role of statistics in national-health policy decisions. American Journal of Epidemiology, 104(4):380–385, ???? 1976. ISSN 0002-9262.

Tukey:1976:CDA

[Tuk76b] John W. Tukey. Comments on Does air pollution cause mortality? In Proceedings of the Fourth Symposium on Statistics and the Environment: a forum for interdisciplinary interaction, March 3-5, 1976, Washington, D.C, pages 37–41. American Statistical Association, Washington, DC, USA, 1976.

Tukey:1976:CRS

[Tuk76c] John W. Tukey. Comments on Role of statistics in national health policy decisions. American Journal of Epidemiology, 104(??):380– 385, October 1976. CODEN AJEPAS. ISSN 0002-9262.

Tukey:1976:MSR

[Tuk76d] John W. Tukey. Methodology and the statistician’s responsibility for both accuracy and relevance. Statistical Reporter, 13(??):253– 262, ???? 1976. CODEN ???? ISSN 0039-050X.

Tukey:1976:RRS

[Tuk76e] John W. Tukey. Robust/resistant smoothing. In Selected Examples of ERDA Research in the Mathematical and Computer Sciences, pages 80–82. National Technical Information Service, Washington, DC, USA, 1976. ERDA 76-118.

Tukey:1976:URRa

[Tuk76f] John W. Tukey. Usable resistant/robust techniques of analysis. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1976. ???? pp.

Tukey:1976:URRb

[Tuk76g] John W. Tukey. Usable resistant/robust techniques of analysis. In W. L. Nicholson and J. L. Harris, editors, Proceedings of the First ERDA Statistical Symposium, Battelle Pacific Northwest Laboratories, Richland, WA, pages 1–31. Battelle Pacific Northwest Laboratories, Richland, WA, USA, 1976.

Tukey:1976:URRc

[Tuk76h] John W. Tukey. Usable resistant/robust techniques of analysis. In W. L. Nicholson and J. L. Harris, editors, Proceedings of the Second ERDA Statistical Symposium, Oak Ridge, Tennessee, October 25–27, 1976, pages 1–31. National Technical Information Service, Washington, DC, USA, 1976.

Tukey:1977:CRL

[Tuk77a] John W. Tukey. Comments on A reformulation of linear models. Journal of the Royal Statistical Society. Series A (General), 140 (??):72, ???? 1977. CODEN JSTAAG. ISSN 0035-9238.

Tukey:1977:DSI

[Tuk77b] John W. Tukey. Discussion of session I and III. In Proceedings of the Second ERDA Statistical Symposium, Oak Ridge, Tennessee, October 25–27, 1976, pages 128–141, 262–273. National Technical Information Service, Washington, DC, USA, 1977.

Tukey:1977:EDA

[Tuk77c] John W. Tukey. Exploratory data analysis. Addison-Wesley, Reading, MA, USA, 1977. ISBN 0-201-07616-0. xvi + 688 pp. LCCN HA29 .T916 1977.

Tukey:1977:HCI

[Tuk77d] John W. Tukey. Higher criticism for individual significances in several tables or parts of tables. Internal working paper 89-9, Princeton University, Princeton, NJ, USA, 1977.

Tukey:1977:IR

[Tuk77e] John W. Tukey. An interim report. Report, Committee on Impacts of Stratospheric Change, National Academy of Sciences, Washington, DC, USA, 1977. 1–61 pp.

Tukey:1977:MST

[Tuk77f] John W. Tukey. Modern statistical techniques in data analysis. In B. Dreyfus, editor, The Proceedings of the Fifth Biennial International CODATA Conference, pages 9–18. Pergamon, New York, NY, USA, 1977.

Tukey:1977:STC

[Tuk77g] John W. Tukey. Some thoughts on clinical trials, especially problems of multiplicity. Science, 198(4318):679–684, November 18, 1977. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Reprinted in Evaluation Studies Review Annual (T. D. Cook and Associates, editors) 3 327–332, Sage, Beverly Hills, CA, USA.

Tukey:1977:SCI

[Tuk77h] John W. Tukey. Statistical considerations important in analyzing measured changes in water-wall tube thickness. In Proceedings of the Second Stationary Source Combustion Symposium V. Addendum. U.S. Environmental Protection Agency, pages 79–92. National Technical Information Service, Washington, DC, USA,

Tukey:1978:NAA

[Tuk78a] J. W. Tukey. New approach to ARMA modeling — comment. Communications in Statistics: Simulation and Computation, 7(1): 79–92, ???? 1978. CODEN CSSCDB. ISSN 0361-0918.

Tukey:1978:CPC

[Tuk78b] John W. Tukey. Comments on a paper by C. W. J. Granger, Seasonality: causation, interpretation, and implications. In A. Zeliner, editor, Seasonal Analysis of Economic Time Series, pages 50–53. United States Government Printing Office, Washington, DC, USA, 1978. ISBN ???? LCCN ???? Reprinted in [Bri84b, page 939].

Tukey:1978:CNA

[Tuk78c] John W. Tukey. Comments on A new approach to ARMA modeling. Communications in Statistics, Part B, Simulation and Computation, 7(??):79–84, ???? 1978. CODEN ???? ISSN 0361-0918.

Tukey:1978:CCT

[Tuk78d] John W. Tukey. Controlled clinical trials. Science, 202(4372): 1105, December 8, 1978. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/202/4372/1105.2.citation.

Tukey:1978:DAC

[Tuk78e] John W. Tukey. A data analyst’s comments on a variety of points and issues. In E. Callaway, P. Tueting, and S. H. Kaslowy, editors, Event-Related Brain Potentials in Man, pages 139–154. Academic Press, New York, USA, 1978. Reprinted in [Bri84b, pp. 915–934].

Tukey:1978:DPH

[Tuk78f] John W. Tukey. Discussion of paper by Herbert Schimmel: Evidence for possible acute health effects of ambient air pollution from time series analysis: methodological questions and some new results based on New York City daily mortality, 1963–1976. Bulletin of the New York Academy of Medicine, 54(??):1111–1112, ???? 1978. CODEN BNYMAM. ISSN 0028-7091.

Tukey:1978:NTL

[Tuk78g] John W. Tukey. The ninther, a technique for low-effort robust (resistant) location in large samples. In H. A. David, editor, Contributions to Survey Sampling and Applied Statistics: Papers in Honor of H. Hartley, pages 251–258. Academic Press, New York, USA, 1978.

Tukey:1978:SP

[Tuk78h] John W. Tukey. Statistics at Princeton. In Alexander Leitch, editor, A Princeton Companion, chapter S, pages 449–451. Princeton University Press, Princeton, NJ, USA, 1978. ISBN 0-691-04654-9. LCCN LD4608 .L4 1978; LD4611 1978.

Tukey:1979:CEP

[Tuk79a] John W. Tukey. Comments on Emanuel Parzen, Nonparametric statistical data modeling. Journal of the American Statistical Association, 74(365):121–122, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2286735. See [Par79].

Tukey:1979:HFD

[Tuk79b] John W. Tukey. Harold F. Dodge, 1893–1976. Journal of the Royal Statistical Society. Series A (General), 142(3):394, ???? 1979. CO-DEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.org/

stable/2982510.

Tukey:1979:MSR

[Tuk79c] John W. Tukey. Methodology, and the statistician’s responsibility for BOTH accuracy AND relevance. Journal of the American Statistical Association, 74(368):786–793, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286400. An expanded version

of ASnumber 121.

Tukey:1979:OHF

[Tuk79d] John W. Tukey. Obituary: Harold F. Dodge, 1893–1976. Journal of the Royal Statistical Society. Series A (General), 142(3):394, 1979. CODEN JSTAAG. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982510.

Tukey:1979:REP

[Tuk79e] John W. Tukey. Re-expression and “plotting” (abstract). Bulletin — Institute of Mathematical Statistics, 8(??):274, ???? 1979. CODEN SMBCVA. ISSN 0146-3942.

Tukey:1979:RTU

[Tuk79f] John W. Tukey. Robust techniques for the user. In R. L. Launer and G. N. Wilkinson, editors, Robustness in Statistics, pages 103– 106. Academic Press, New York, USA, 1979.

Tukey:1979:SMW

[Tuk79g] John W. Tukey. Statistical mapping: What should not be plotted. In Proceedings of the 1976 Workshop on Automated Cartography and Epidemiology, March 18-19, 1976, Arlington, Virginia, volume 1976, pages 18–26. U.S. Dept. of Health, Education, and Welfare, Public Health Service, Office of Health Research, Statistics, and Technology, National Center for Health Statistics, Hyattsville, MD, USA, 1979. LCCN RA792.5 .W67 1976. DHEW Publication (PHS) 79-1254.

Tukey:1979:SRS

[Tuk79h] John W. Tukey. Study of robustness by simulation: Particularly improvement by adjustment and combination. In R. L. Launer and G. N. Wilkinson, editors, Robustness in Statistics, pages 75–102. Academic Press, New York, USA, 1979.

Tukey:1980:SDA

[Tuk80a] J. W. Tukey. Styles of data analysis, and their implications for statistical computing. In M. M. Barrett and D. Wishart, editors, Compstat 1980 (Proceedings Fourth Symposium Comput. Statist., Edinburgh, 1980), pages 21–31. Physica Verlag, Vienna, Austria / New York, NY, USA, 1980.

Tukey:1980:CWP

[Tuk80b] John W. Tukey. Can we predict where “time series” should go next? In Directions in time series (Proceedings Meeting, Iowa State Univ., Ames, Iowa, 1978), pages 1–31. Institute for Mathematical Statistics, Hayward, CA, USA, 1980. URL 2-

E&origin=MSN;   http://links.jstor.org/sici?sici=0002-
9890(196701)74:1<101:WCMDFT>2.0.CO. Reprinted in [Bri84b,

pp. 941–980].

Tukey:1980:LAB

[Tuk80c] John W. Tukey. Looking ahead to behavioral science work at Bell Laboratories. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1980. ???? pp.

Tukey:1980:MCF

[Tuk80d] John W. Tukey. Methodological comments focused on opportunities. In P. R. Monge and J. Cappella, editors, Multivariate Techniques in Human Communication Research, pages 489–528 (Chapter 16). Academic Press, New York, USA, 1980.

Tukey:1980:WNB

[Tuk80e] John W. Tukey. We need both exploratory and confirmatory. The American Statistician, 34(1):23–25, February 1980. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682991.

Tukey:1981:ARN

[Tuk81a] John W. Tukey. Analiz Rezul’tatov Nablyudeniyi. (Russian) [The Analysis of the Results of Observations]. Mir Press, Moskva, USSR, 1981. ISBN ???? ???? pp. LCCN ???? Russian edition of Exploratory Data Analysis. Translated by Ph.D. candidates A. F. Kusznira, A. L. Petrosyana and E. L. Reznikova, under the direction of B. F. Pisarenko.

Tukey:1981:KPI

[Tuk81b] John W. Tukey. Kinds of polyconfidence intervals for centers, and some thoughts on identification and selection of confidence procedures using configural polysampling. Technical report, series 2 190, Princeton University, Princeton, NJ, USA, 1981. 30 pp.

Tukey:1981:SAT

[Tuk81c] John W. Tukey. Some advance thoughts on the data analysis involved in configural polysampling directed toward high performance estimates. Technical report, series 2 189, Princeton University, Princeton, NJ, USA, 1981. 17 pp.

Tukey:1982:ALF

[Tuk82a] John W. Tukey. Another look at the future. In K. W. Heiner, R. S. Sacher, and J. W. Wilkinson, editors, Computer Science and Statistics: Proceedings of the 14th Symposium on the Interface, pages 2–8. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982.

Tukey:1982:CSP

[Tuk82b] John W. Tukey. Control and stash philosophy for two-handed, flexible and immediate control of a graphical display. Technical report, series 2 221, Bell Laboratories, Murray Hill, NJ, USA, 1982. 35 pp.

Tukey:1982:DPC

[Tuk82c] John W. Tukey. Discussion of the paper by A. C. Atkinson, Regression diagnostics, transformations and constructed variables. Journal of the Royal Statistical Society. Series B (Methodological), 44 (??):31, ???? 1982. CODEN JSTBAJ. ISSN 0035-9246.

Tukey:1982:DRS

[Tuk82d] John W. Tukey. Discussion [the role of statistical graduate training]. In Jagdish S. Rustagi and Douglas A. Wolfe, editors, Teaching of statistics and statistical consulting: proceedings of a conference held at the Ohio State University, November 24-25, 1980, pages 379–389. Academic Press, New York, USA, 1982. ISBN 0-12-604540-2. LCCN QA276.18 .T43 1982.

Tukey:1982:ISD

[Tuk82e] John W. Tukey. Introduction to styles of data analysis techniques. In R. L. Launer and A. F. Siegel, editors, Modern Data Analysis, pages 1–11. Academic Press, New York, USA, ???? 1982.

Tukey:1982:PRP

[Tuk82f] John W. Tukey. A pattern-recognition problem: curly contrails in the snow! Technical report, series 2 227, Princeton University, Princeton, NJ, USA, 1982. 32 pp.

Tukey:1982:SGSa

[Tuk82g] John W. Tukey. Some graphics for studying four-dimensional data. Technical report, Bell Laboratories, Murray Hill, NJ, USA, 1982. ???? pp.

Tukey:1982:USG

[Tuk82h] John W. Tukey. The use of smelting in guiding re-expression. In R. L. Launer and A. F. Siegel, editors, Modem Data Analysis, pages 83–102. Academic Press, New York, USA, 1982.

Tukey:1982:CFY

[Tuk82i] Paul A. Tukey. Collaboration with Frank Yates — a personal view. Utilitas Mathematica, 21(B):377–379, ???? 1982. CODEN UTMADA. ISSN 0315-3681.

Tukey:1983:ARE

[Tuk83] John W. Tukey. Appendix A: The relationship of empirical analysis to more narrowly modeled analysis. In Mary B. Breckenridge, editor, Age, Time, and Fertility: Applications of Exploratory Data Analysis, pages 274–281. Academic Press, New York, USA, 1983. ISBN 0-12-128750-5. LCCN HB1027 .B73 1983.

Tukey:1984:USA

[Tuk84a] J. W. Tukey. Use of spatial-analysis in mineral-resource evaluation — comments. Journal of the International Association for Mathematical Geology, 16(6):591–594, ???? 1984. CODEN IMGJBS. ISSN 0020-5958.

Tukey:1984:CUS

[Tuk84b] John W. Tukey. Comments on Use of spatial analysis in mineral resource’s valuation. Journal of the International Association for Mathematical Geology, 16(??):591–594, ???? 1984. CODEN IMGJBS. ISSN 0020-5958.

Tukey:1984:DAH

[Tuk84c] John W. Tukey. Data analysis: History and prospects. In H. A. David and David H. I, editors, Statistics: An Appraisal: proceedings of a conference marking the 50th anniversary of the Statistical Laboratory, Iowa State University, pages 183–202. Iowa State University Press, Ames, IA, USA, ???? 1984.

Tukey:1984:EUA

[Tuk84d] John W. Tukey. Equalizing unequal accuracies of means. In Braun [Bra94], pages 301–313. ISBN 0-412-05121-4. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1984:FSL

[Tuk84e] John W. Tukey. First Scott lecture. In Brillinger [Bri84b], pages 857–884. ISBN 0-534-03304-0. With introductory material by William S. Cleveland and Frederick Mosteller.

Tukey:1984:FMF

[Tuk84f] John W. Tukey. Fourier methods in the frequency analysis of data, and souvenir sheets. In Brillinger [Bri84b], pages 755–780. ISBN 0-534-03304-0. With introductory material by William S. Cleveland and Frederick Mosteller.

Tukey:1984:MIF

[Tuk84g] John W. Tukey. Mathematics 596: An introduction to the frequency analysis of time series. In Brillinger [Bri84a], pages 503– 650. ISBN 0-534-03303-2. LCCN QA276.A12 T85 1984. With introductory material by William S. Cleveland and Frederick Mosteller.

Tukey:1984:NNM

[Tuk84h] John W. Tukey. Nonlinear (nonsuperposable) methods for smoothing data. In Brillinger [Bri84b], pages 837–856. ISBN 0-534-03304-0. With introductory material by William S. Cleveland and Frederick Mosteller.

Tukey:1984:PMC

[Tuk84i] John W. Tukey. The problem of multiple comparisons. In Braun [Bra94], pages 1–300. ISBN 0-412-05121-4. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1984:RRJ

[Tuk84j] John W. Tukey. Remarks on receipt of James Madison Medal. Unpublished., 1984.

Tukey:1984:RSA

[Tuk84k] John W. Tukey. Reminder sheets for Allowances for various types of error rates. In Braun [Bra94], pages 335–339. ISBN 0-412-05121-4. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1984:RSD

[Tuk84l] John W. Tukey. Reminder sheets for Discussion of paper on multiple comparisons by Henry Scheffé. In Braun [Bra94], pages 469– 475. ISBN 0-412-05121-4. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1984:RSM

[Tuk84m] John W. Tukey. Reminder sheets for Multiple Comparisons. In Braun [Bra94], pages 341–345. ISBN 0-412-05121-4. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1984:SSL

[Tuk84n] John W. Tukey. Second Scott lecture. In Brillinger [Bri84b], pages 885–914. ISBN 0-534-03304-0. With introductory material by William S. Cleveland and Frederick Mosteller.

Tukey:1984:SSC

[Tuk84o] John W. Tukey. Souvenir sheets for Confidence procedures are better. In Braun [Bra94], pages 347–351. ISBN 0-412-05121-4. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1984:SSA

[Tuk84p] John W. Tukey. Styles of spectrum analysis. In A Celebration in Geophysics and Oceanography—1982, in Honor of Walter Munk, pages 100–103. Scripps Institute of Oceanography, La Jolla, CA, USA, 1984. Reprinted in [Bri84b, 1143–1153].

Tukey:1984:STR

[Tuk84q] John W. Tukey. Supplement to Technical Report 269, series 2. Technical report, series 2 269, Princeton University, Princeton, NJ, USA, 1984. 6 pp.

Tukey:1984:WSW

[Tuk84r] John W. Tukey. When should which spectrum approach be used? In Brillinger [Bri84b], pages 981–1000. ISBN 0-534-03304-0. Reprinted in [Bri84b, 981–1000].

Tukey:1985:PPD

[Tuk85a] J. Tukey. Projection pursuit — discussion. Annals of Statistics, 13(2):517–518, ???? 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Tukey:1985:CEP

[Tuk85b] John W. Tukey. Comments on Estimating the population in a census year: 1980 and beyond. Journal of the American Statistical Association, 80(389):127–128, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288061.

Tukey:1985:CMI

[Tuk85c] John W. Tukey. Comments on More intelligent statistical software and statistical expert systems: Future directions. The American Statistician, 39(1):12–14, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683898.

Tukey:1985:DPC

[Tuk85d] John W. Tukey. Discussion of the paper by C. Chatfield, The initial examination of data. Journal of the Royal Statistical Society. Series A (General), 148(??):246–247, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

Tukey:1985:DPP

[Tuk85e] John W. Tukey. Discussion: Projection pursuit. The American Statistician, 39(2):517–518, June 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349533;   http://www.
jstor.org/stable/2241189.

Tukey:1985:EPC

[Tuk85f] John W. Tukey. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):127–128, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288061.

Tukey:1985:ICR

[Tuk85g] John W. Tukey. Improving crucial randomized experiments— especially in weather modification—by double randomization and rank combination. In Proceedings of the Berkeley conference in honor of Jerzy Neyman and Jack Kiefer, Volume I (Berkeley, California, 1983), Wadsworth & Brooks/Cole Statistics/Probability Series, pages 79–108. Wadsworth, Pacific Grove, CA, USA, 1985.

Tukey:1985:LRD

[Tuk85h] John W. Tukey. Limited randomization with detailed reassignment as the key to taking advantage of modern summaries. In Design of experiments in army research, development and testing, Proc. 30th Conf., Las Cruces/N.M. 1984, ARO Rep. 85-2, pages 479– 490. ????, ????, 1985.

Tukey:1985:MIS

[Tuk85i] John W. Tukey. [more intelligent statistical software and statistical expert systems: Future directions]: Comment. The American Statistician, 39(1):12–14, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683898.

Tukey:1985:RLR

[Tuk85j] John W. Tukey. The resistant line and related regression methods: Comment: The variance of slopes of lines fitted by groups: An analysis of the Johnston and Velleman Monte Carlo results. Journal of the American Statistical Association, 80(392):1055– 1059, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288573.

Tukey:1986:ASE

[Tuk86a] John W. Tukey. An alphabet for statisticians’ expert systems. In W. A. Gale, editor, Artificial Intelligence and Statistics, pages 401–409. Addison-Wesley, Reading, MA, USA, 1986.

Tukey:1986:CTA

[Tuk86b] John W. Tukey. Choosing techniques for the analysis of data. In Jones [Jon86b], pages 869–874. ISBN 0-534-05101-4. With a biography of Tukey by Frederick Mosteller.

Tukey:1986:CNV

[Tuk86c] John W. Tukey. Comments on The natural variability of vital rates and associated statistics. Biometrics, 42(??):729–732, ???? 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1986:CDA

[Tuk86d] John W. Tukey. Comments on Deriving appropriate standards for Census undercount adjustment. ????, ??(??):??, ???? 1986. CODEN ???? ISSN ????

Tukey:1986:CPJ

[Tuk86e] John W. Tukey. Comments on paper by J. Heckman and R. Robb, Alternative methods for solving the problem of selection bias in evaluating the impact of treatments on outcomes. In Drawing Inferences from Self-Selected Samples, pages 108–110. Lawrence Erlbaum Associates, Mahwah, NJ, USA, 1986. ISBN 0-8058-3802-

3. LCCN HA31.2 .D7 2000. URL http://www.loc.gov/catdir/
enhancements/fy0634/00268843-d.html.

Tukey:1986:DAB

[Tuk86f] John W. Tukey. Data analysis and behavioral science or learning to bear the quantitative man’s burden by shunning badmandments. In Jones [Jon86a], pages 187–389. ISBN 0-534-03305-9. With a biography of Tukey by Frederick Mosteller.

Tukey:1986:DPB

[Tuk86g] John W. Tukey. Discussion of paper by B. Singer, Self-selection and performance-based ratings: a case study in program evaluation. In Drawing Inferences from Self-Selected Samples, pages 52–55.

Lawrence Erlbaum Associates, Mahwah, NJ, USA, 1986. ISBN 0-8058-3802-3. LCCN HA31.2 .D7 2000. URL http://www.loc.

gov/catdir/enhancements/fy0634/00268843-d.html.

Tukey:1986:DPH

[Tuk86h] John W. Tukey. Discussion of paper by H. Wainer, The SAT as a social indicator: a pretty bad idea. In Drawing Inferences from Self-Selected Samples, pages 24–25. Lawrence Erlbaum Associates, Mahwah, NJ, USA, 1986. ISBN 0-8058-3802-3. LCCN HA31.2

.D7 2000. URL http://www.loc.gov/catdir/enhancements/
fy0634/00268843-d.html.

Tukey:1986:DPJ

[Tuk86i] John W. Tukey. Discussion of paper by J. Heckman, Alternative methods for evaluating the impact of intervention. In Drawing Inferences from Self-Selected Samples, pages 58–60. Lawrence Erlbaum Associates, Mahwah, NJ, USA, 1986. ISBN 0-8058-3802-

3. LCCN HA31.2 .D7 2000. URL http://www.loc.gov/catdir/
enhancements/fy0634/00268843-d.html.

Tukey:1986:DPR

[Tuk86j] John W. Tukey. Discussion of paper by R. Glynn, N. Laird and D. Rubin, Mixture modeling versus selection modeling with nonignorable nonresponse. In Drawing Inferences from Self-Selected Samples, page 143. Lawrence Erlbaum Associates, Mahwah, NJ, USA, 1986. ISBN 0-8058-3802-3. LCCN HA31.2 .D7 2000. URL http://

www.loc.gov/catdir/enhancements/fy0634/00268843-d.html.

Tukey:1986:DSV

[Tuk86k] John W. Tukey. Discussion of Session VIII-b. In Proceedings of the Second Annual Research Conference, pages 408–410. Bureau of the Census, Washington, DC, USA, 1986. ISBN ????

Tukey:1986:DDC

[Tuk86l] John W. Tukey. Do derivations come from heaven? In Jones [Jon86b], pages 875–880. ISBN 0-534-05101-4. With a biography of Tukey by Frederick Mosteller.

Tukey:1986:ICS

[Tuk86m] John W. Tukey. The interface with computing: In the small or in the large. In T. J. Boardman, editor, Computer Science and Statistics: Proceedings of the 18th Symposium the Interface, pages 3–7. American Statistical Association, Washington, DC, USA, 1986.

Tukey:1986:SS

[Tuk86n] John W. Tukey. Sunset salvo. The American Statistician, 40 (1):72–76, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/ stable/2683137. Reprinted in [Jon86b, 1003–1016].

Tukey:1986:WSB

[Tuk86o] John W. Tukey. What have statisticians been forgetting? In Jones [Jon86b], pages 587–599. ISBN 0-534-05101-4. With a biography of Tukey by Frederick Mosteller.

Tukey:1987:WAV

[Tuk87a] J. W. Tukey. What is an analysis of variance — discussion. Annals of Statistics, 15(3):936–937, ???? 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Tukey:1987:CDG

[Tuk87b] John W. Tukey. Comments on Dynamic graphics for data analysis. Statistical Science, 2(4):383–385, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013105;   http:
//www.jstor.org/stable/2245524.

Tukey:1987:CP

[Tuk87c] John W. Tukey. Configural polysampling. SIAM Review, 29(1): 1–20, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). 1985 John von Neumann Lecture.

Tukey:1987:DPM

[Tuk87d] John W. Tukey. Discussion of paper by M. C. Jones and Robin Sibson, What is projection pursuit? Journal of the Royal Statistical Society. Series A (General), 150(1):33, ???? 1987. CODEN

JSSAEF. ISSN 0035-9238. URL http://www.jstor.org/stable/
2981662.

Tukey:1987:DWA

[Tuk87e] John W. Tukey. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):936–937, ???? 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350483.

Tukey:1987:KBK

[Tuk87f] John W. Tukey. Kinds of bootstraps and kinds of jackknives, discussed in terms of a year of weather-related data. Technical report, series 2 292, Princeton University, Princeton, NJ, USA, 1987. 56 pp.

Tukey:1988:CDG

[Tuk88a] John W. Tukey. Comments on Dynamic graphics for data analysis. In Cleveland and McGill [CM88], pages 50–54. ISBN 0-534-09144-X. LCCN QA276.3 .D96 1988. URL http://www.loc.gov/

catdir/enhancements/fy0744/88026017-d.html.

Tukey:1988:CSPa

[Tuk88b] John W. Tukey. Control and stash philosophy for two-handed, flexible, and immediate control of a graphic display. In Cleveland and McGill [CM88], pages 133–178. ISBN 0-534-09144-X.

LCCN QA276.3 .D96 1988. URL http://www.loc.gov/catdir/
enhancements/fy0744/88026017-d.html.

Tukey:1988:CSPb

[Tuk88c] John W. Tukey. Control and stash philosophy for two-handed, flexible, and immediate control of a graphic display. In Cleveland [Cle88], pages 329–382. ISBN 0-534-05102-2. With a biography of Tukey by Frederick Mosteller.

Tukey:1988:DAS

[Tuk88d] John W. Tukey. Data analysis and statistics: Techniques and approaches. In Cleveland [Cle88], pages 1–22. ISBN 0-534-05102-2. With a biography of Tukey by Frederick Mosteller.

Tukey:1988:NPC

[Tuk88e] John W. Tukey. Notch plots for counted rates. In Cleveland [Cle88], pages 79–92. ISBN 0-534-05102-2. With a biography of Tukey by Frederick Mosteller.

Tukey:1988:TED

[Tuk88f] John W. Tukey. Thoughts on the evolution of dynamic graphics for data-modification display. In Cleveland [Cle88], pages 383– 401. ISBN 0-534-05102-2. With a biography of Tukey by Frederick Mosteller.

Tukey:1989:SYA

[Tuk89] John W. Tukey. SPES in the years ahead. In Proceedings of the American Statistical Association: Sesquicentennial Invited Paper Sessions, pages 175–182. American Statistical Association, Washington, DC, USA, 1989.

Tukey:1990:BCC

[Tuk90a] John W. Tukey. Better coordinates for combinations of order statistics. In Mallows [Mal90], pages 429–433. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:BFM

[Tuk90b] John W. Tukey. Biography of Frederick Mosteller. In S. E. Fienberg, W. H. Kruskal, D. C. Hoaglin, and J. M. Tanur, editors, A Statistical Model. Frederick Mosteller’s Contributions to Statistics, Science, and Public Policy, pages 1–5. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990.

Tukey:1990:DBG

[Tuk90c] John W. Tukey. Data-based graphics: visual display in the decades to come. Statistical Science, 5(3):327–339, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.ss/
1177012101;   http://www.jstor.org/stable/2245820.

Tukey:1990:DLR

[Tuk90d] John W. Tukey. Determination of linear relations between systematic parts. In Mallows [Mal90], pages 317–329. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:FCP

[Tuk90e] John W. Tukey. The finite case of the Problem of the Nile. In Mallows [Mal90], pages 35–40. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:FEC

[Tuk90f] John W. Tukey. Foreword for The effect of changes in national assessment: Disentangling the NAEP 1985–86 reading anomaly, by A. E. Beaton and R. Zwick. Report 17-TR21, National Assessment of Educational Progress, Educational Testing Service, Riverdale Road, Princeton, NJ 08541-0001, USA, 1990.

Tukey:1990:HWL

[Tuk90g] John W. Tukey. Handouts for the Wald lectures 1958. In Mallows [Mal90], pages 119–148. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:IGR

[Tuk90h] John W. Tukey. Introduction to guided re-expression. In Mallows [Mal90], pages 221–236. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:NF

[Tuk90i] John W. Tukey. NkNS forests. In Mallows [Mal90], pages 553–556. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:NLS

[Tuk90j] John W. Tukey. A note on least squares and unbiased estimates. In Mallows [Mal90], pages 421–427. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:ODF

[Tuk90k] John W. Tukey. One degree of freedom or several? Parsimony in detection of an effect. In Mallows [Mal90], pages 407–419. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:PRB

[Tuk90l] John W. Tukey. The practical relationship between the common transformations of percentages or fractions and of amounts. In Mallows [Mal90], pages 211–219. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:PSF

[Tuk90m] John W. Tukey. The present state of fiducial probability. In Mallows [Mal90], pages 55–118. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:SFM

[Tuk90n] John W. Tukey. Some further multivariate suggestions. In Mallows [Mal90], pages 557–583. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:SGP

[Tuk90o] John W. Tukey. Some general principles and approximations for sequentially studentized procedures. In Mallows [Mal90], pages 367–371. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:SSU

[Tuk90p] John W. Tukey. Souvenir sheet for Use of control classifications: adjustment for inadequacy of broad classes. In Mallows [Mal90], pages 435–443. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:SSC

[Tuk90q] John W. Tukey. Souvenir sheets for The criticism of transformations. In Mallows [Mal90], pages 157–165. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:SCP

[Tuk90r] John W. Tukey. Standard confidence points. In Mallows [Mal90], pages 331–365. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1990:STU

[Tuk90s] John W. Tukey. Steps toward a universal univariate distribution analyzer. In Mallows [Mal90], pages 585–590. ISBN 0-534-05103-0. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Tukey:1991:ACA

[Tuk91a] John W. Tukey. Appendix on a circular approach to cubature. In Morgenthaler and Tukey [MT91d], pages 213–221. ISBN 0-471-52372-0. LCCN QA276.6 .C66 1991.

Tukey:1991:CD

[Tuk91b] John W. Tukey. Consumer datesware. In W. Stahel and S. Weisberg, editors, Directions in Robust Statistics and Diagnostics, Part II, pages 297–308. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Tukey:1991:GDA

[Tuk91c] John W. Tukey. Graphical displays for alternate regression fits. In W. Stahel and S. Weisberg, editors, Directions in Robust Statistics and Diagnostics, Part II, pages 309–325. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Tukey:1991:PMC

[Tuk91d] John W. Tukey. The philosophy of multiple comparisons. Statistical Science, 6(1):100–116, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177011945;   http://www.

jstor.org/stable/2245714. 1989 Miller Lecture presented at Stanford University.

Tukey:1991:UMC

[Tuk91e] John W. Tukey. Use of many covariates in clinical trials. International Statistical Review = Revue Internationale de Statistique, 59(2):123–137, ???? 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tukey:1992:DPB

[Tuk92a] John W. Tukey. Discussion of paper by Bradley Efron Jackknifeafter-bootstrap standard errors and influence functions. Journal of the Royal Statistical Society. Series B (Methodological), 54(??): 120–121, ???? 1992. CODEN JSTBAJ. ISSN 0035-9246.

Tukey:1992:ESN

[Tuk92b] John W. Tukey. Exbrids: Simple nearly symmetrizing reexpressions for exponentially distributed quantities. In K. V. Mardia, editor, The Art of Statistical Science: A Tribute to G. S. Watson, pages 183–187. Wiley, New York, NY, USA, 1992.

Tukey:1992:R

[Tuk92c] John W. Tukey. Reflections. In Brillinger et al. [BCG^{+}93], page 387. ISBN 0-387-97896-8 (vol. 1), 0-387-97914-X (vol. 2). LCCN QA280 .N47 1992.

Tukey:1992:WSM

[Tuk92d] John W. Tukey. Where should multiple comparisons go next? In F. M. Hoppe, editor, Multiple Comparisons, Selection, and Applications in Biometry. A Festschrift in Honor of Charles W Dunnett, pages 187–207. Marcel Dekker, New York, NY, USA, 1992.

Tukey:1993:EAV

[Tuk93a] John W. Tukey. Exploratory analysis of variance as providing examples of strategic choices. In S. Morgenthaler, E. Ronchetti, and W. Stahel, editors, New Directions in Statistical Data Analysis and Robustness, pages 205–258. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1993.

Tukey:1993:GCS

[Tuk93b] John W. Tukey. Graphic comparisons of several linked aspects: alternatives and suggested principles. with comments by Howard Wainer, Richard A. Becker and William S. Cleveland and a rejoinder by the author. Journal of Computational and Graphical Statistics, 2(1):1–33, March 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). Discussion pp. 35–49. With comments by Howard Wainer, Richard A. Becker and William S. Cleveland and a rejoinder by the author.

Tukey:1993:MCM

[Tuk93c] John W. Tukey. Major challenges for multiple-response (and multiple-adjustment) analysis. In C. R. Rao, editor, Multivariate Analysis: Future Directions, pages 401–421. North-Holland Publishing Co., Amsterdam, The Netherlands, 1993.

Tukey:1993:R

[Tuk93d] John W. Tukey. Rejoinder. Journal of Computational and Graphical Statistics, 2(1):49, March 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10618600.1993.10474598. Reply to com-

ments on Graphic comparisons of several linked aspects: Alternatives and suggested principles. See [Tuk93b].

Tukey:1993:TCT

[Tuk93e] John W. Tukey. Tightening the clinical trial. Controlled Clinical Trials, 14(??):266–285, ???? 1993. CODEN CCLTDH. ISSN 0197-2456 (print), 1879-050X (electronic).

Tukey:1993:WSM

[Tuk93f] John W. Tukey. Where should multiple comparisons go next? In F. M. Hoppe, editor, Statistics (New York); Multiple comparisons, selection, and applications, volume 134, pages 187–207. Marcel Dekker, New York, NY, USA, 1993. ISBN 0-8247-8895-8. LCCN ????

Tukey:1995:CPF

[Tuk95a] John W. Tukey. Controlling the proportion of false discoveries for multiple comparison — future directions. In Williams et al. [WJO95], page ?? LCCN ???? Technical report 35.

Tukey:1995:IRH

[Tuk95b] John W. Tukey. Issues relevant to an honest account of data-based inference, partially in the light of Laurie Davies’ paper. Preprint, Princeton University, Princeton, NJ, USA, 1995.

Tukey:1997:SJB

[Tuk97a] E. R. Tukey. Speech at JWT’s 80th birthday celebration. In Brillinger et al. [BFM97], pages 46–47. ISBN 0-691-05782-6. LCCN QA276.16 .P73 1997. URL http://www.loc.

gov/catdir/description/prin021/97019695.html;   http:/
/www.loc.gov/catdir/toc/prin031/97019695.html. Two-day

symposium, Princeton on June 19–20, 1995.

Tukey:1997:MHF

[Tuk97b] John W. Tukey. More honest foundations for data analysis. Journal of Statistical Planning and Inference, 57(??):21–28, ???? 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Tukey:1997:RSM

[Tuk97c] John W. Tukey. Report to the Special Master. In R. S. Cohen, editor, Report to the New Jersey Supreme Court, page ?? ????, ????, 1997. ISBN ???? LCCN ????

Tukey:1998:WMW

[Tuk98] John W. Tukey. Where might we be going? Unpublished., 1998.

Tukey:1985:VSL

[TV85] John W. Tukey and P. Velleman. The variance of slopes of lines fitted by groups: An analysis of the Johnstone and Velleman Monte Carlo results. comment on “the resistant line and related regression methods,” by I. Johnstone and P. Velleman. Journal of the American Statistical Association, 80(392):1055–1059, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288573.

Tukey:1946:ADP

[TW46a] John W. Tukey and S. S. Wilks. Approximation of the distribution of the product of beta variables by a single beta variable. Annals of Mathematical Statistics, 17(3):318–324, ???? 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730944.

Tukey:1946:IS

[TW46b] John W. Tukey and C. P. Winsor. Industrial statistics. Journal of the American Statistical Association, 41(??):406–411, ???? 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1946:BRB

[TW46c] John W. Tukey and Charles P. Winsor. Book review: Industrial Statistics. Journal of the American Statistical Association, 41(235):406–411, September 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280270.

Tukey:1946:RBE

[TW46d] John W. Tukey and J. Wolfowitz. Review: Industrial Experimentation, by K. A. Brownlee. Journal of the American Statistical Association, 41(233):125–128, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2280164.

Tukey:1949:NSN

[TW49] John W. Tukey and Charles P. Winsor. Note on some χ^{2} normalizations. SRG Memorandum report 29, Princeton University, Princeton, NJ, USA, 1949. 3 pp.

Tukey:1952:RBM

[TW52] John W. Tukey and J. Wolfowitz. Review of Statistische Methoden für Naturwissenschafter, Mediziner, und Ingenieure [Statistical Methods for Natural Scientists, Physicians, and Engineers], by A. Linder. Journal of the American Statistical Association, 47 (259):554–556, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281326.

Tukey:1965:DAS

[TW65] John W. Tukey and M. B. Wilk. Data analysis and statistics: Techniques and approaches. In P. W. Nye, editor, Proceedings of the Symposium on Information Processing in Sight Sensory Systems, pages 7–27. California Institute of Technology, Pasadena, CA, USA, 1965.

Tukey:1966:DAS

[TW66] J. W. Tukey and M. B. Wilk. Data analysis and statistics: an expository overview. In Proceedings of AFIPS ’66 Fall Joint Computer Conference (November 7–10, 1966), pages 695–709. AFIPS Press, Montvale, NJ, USA, 1966. LCCN ????

Vaswani:1947:PCT

[Vas47] S. P. Vaswani. A pitfall in correlation theory. Nature, 160(4064): 405–406, September 20, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Vernotte:1947:TLD

[Ver47a] Pierre Vernotte. Les termes librement déterminés qui doivent compléter les développements limités. Comptes rendus de l’Académie des sciences, Paris, 224:32–34, 1947.

Vernotte:1947:PDC

[Ver47b] Pierre Vernotte. A propos du calcul pratique de la limite d’une variable. Comptes rendus de l’Académie des sciences, Paris, 225: 1130–1132, 1947.

Villars:1947:STE

[Vil47] D. S. Villars. A significance test and estimation in the case of exponential regression. Annals of Mathematical Statistics, 18:596– 600, 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Vijverberg:2016:LTL

[VVT16] Chu-Ping C. Vijverberg, Wim P. M. Vijverberg, and Süleyman Taspinar. Linking Tukey’s legacy to financial risk measurement. Computational Statistics & Data Analysis, ??(??):595–615, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167947315002066.

Wainer:1990:GVW

[Wai90] Howard Wainer. Graphical visions from William Playfair to John Tukey. Statistical Science, 5(3):340–346, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012102;   http:
//www.jstor.org/stable/2245821.

Wainer:1993:GMC

[Wai93] Howard Wainer. Graphing multiple comparisons: Some comments on Tukey. Journal of Computational and Graphical Statistics, 2 (1):35–40, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/10618600.1993.10474596. See [Tuk93b].

Wainer:2003:JWT

[Wai03] Howard Wainer. John Wilder Tukey: Statistical inventor, discoverer and revolutionary. Statistical Science, 18(3):285–286, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.ss/
1076102416.

Watt:2010:IPI

[Wat10] Stephen M. Watt, editor. ISSAC 2010: Proceedings of the 2010 International Symposium on Symbolic and Algebraic Computation, July 25–28, 2010, Munich, Germany. ACM Press, New York, NY 10036, USA, 2010. ISBN 1-4503-0150-9. LCCN QA76.95 .I59 2010.

Wintner:1938:LAD

[WBST38] Aurel Wintner, Ralph Philip Boss, Frank Smithies, and John Wilder Tukey. Lectures by Aurel Wintner on asymptotic distributions and infinite convolutions, 1937–1938. Edwards Brothers, Inc., Ann Arbor, Mich., 1938. 2 + 54 pp. LCCN QA273 .W65.

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1988/pdf/a4348.pdf;   http://www.computer.org/
annals/an1988/a4348abs.htm.

Wheeler:1962:FTN

[Whe62] John Archibald Wheeler. Fission then and now. International Atomic Energy Agency Bulletin, 4(0):33–36, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (elec-

tronic). URL http://www.iaea.org/Publications/Magazines/
Bulletin/Bull040su/04005093336su.pdf.

Wiener:1956:NPD

[Wie56] Norbert Wiener. Nonlinear prediction and dynamics. In Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability, 1954–1955, vol. III, pages 247–252. University of California Press, Berkeley, CA, USA, 1956.

Wishart:1947:CLD

[Wis47a] John Wishart. The cumulants of the Z and of the logarithmic χ^{2} and t distributions. Biometrika, 34(1/2):170–178, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332520.

Wisseroth:1947:GVN

[Wis47b] K. Wisseroth. Die günstigste Verteilungsbreite, ein neues Streuungsmass. Zeitschrift für Angewandte Mathematik und Mechanik, 25/27:126–127, 1947. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Williams:1995:PSE

[WJO95] Valerie S. L. Williams, Lyle V. Jones, and Ingram Olkin, editors. Perspectives on statistics for educational research: proceedings of a workshop. National Institute of Statistical Sciences, Research Triangle Park, NC, USA, 1995. LCCN ???? Technical report 35.

Williams:1994:CEM

[WJT94] V. S. L. Williams, L. V. Jones, and J. W. Tukey. Controlling error in multiple comparisons, with special attention to the National Assessment of Educational Progress. Technical Report 33, National Institute of Statistical Sciences, Research Triangle Park, NC, USA, 1994.

Williams:1999:CEM

[WJT99] V. S. L. Williams, L. V. Jones, and John W. Tukey. Controlling error in multiple comparisons, with examples from state-to-state differences in educational achievement. Journal of Educational and Behavioral Statistics, 24(1):42–69, Spring 1999. CODEN ???? ISSN 1076-9986 (print), 1935-1054 (electronic). URL http://www.

jstor.org/stable/1165261.

Wolfowitz:1946:CLF

[Wol46] J. Wolfowitz. Confidence limits for the fraction of a normal population which lies between two given limits. Annals of Mathematical Statistics, 17:483–488, 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Wong:2000:BRP

[Won00] Weng Kee Wong. Book review: The Practice of Data Analysis: Essays in Honor of John W. Tukey. D. R. Brillinger, L. T. Fernholz and S. Morgenthaler (eds.), Princeton University Press, U.S.A., 1998. No. of pages: 337. Price: $49.50. ISBN 0-691-05782-6. Statistics in Medicine, 19(6):884–885, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yates:1937:DAF

[Yat37] Frank Yates. The design and analysis of factorial experiments. Technical communication 35, Imperial Bureau of Soil Science, Harpenden, UK, 1937. ??–?? pp. URL http://en.wikipedia.

org/wiki/Frank_Yates.

Yan:2019:TGH

[YG19] Yuan Yan and Marc G. Genton. The Tukey g-and-h distribution. Significance (Oxford, England), 16(3):12–13, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Youden:1959:DPM

[YKT^{+}59] W. J. Youden, Oscar Kempthorne, John W. Tukey, G. E. P. Box, J. S. Hunter, F. E. Satterthwaite, and Thomas A. Budne. Discussion of the papers of Messrs. Satterthwaite and Budne. Technometrics, 1(2):157–193, May 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266468.

Young:1995:BRB

[You95] Karen Young. Book review: The Collected Works of John W. Tukey: Vol. VIII, Multiple Comparisons: 1948–1983, by John W. Tukey; H. I. Braun. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):632, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983453.

Ziegel:1988:BRBs

[Zie88] Eric R. Ziegel. Book review: The Collected Works of John W. Tukey [Vol. 3 (1949–1964) and Vol. 4 (1965–1986): Philosophy and Principles of Data Analysis] by Lyle Jones; John W. Tukey. Technometrics, 30(3):363, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270109.

Ziegel:1989:BRBs

[Zie89] Eric Ziegel. Book review: The Collected Works of John W. Tukey (Vol. 5, Graphics, 1965–1985) by W. S. Cleveland. Technometrics, 31(3):398, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/3556173.

Ziegel:1991:BRBj

[Zie91] Eric R. Ziegel. Book review: Collected Works of John W. Tukey, Volume 6: More Mathematical 1938–1984 by Colin Mallows; John W. Tukey. Technometrics, 33(2):247, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269066.

Ziegel:1995:BRBz

[Zie95] Eric R. Ziegel. Book review: The Collected Works of John Tukey: Volume VIII, Multiple Comparisons by H. Braun; John Tukey. Technometrics, 37(2):242, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269652.

Ziegel:2001:BRBar

[Zie01] Eric R. Ziegel. Book review: Understanding Robust and Exploratory Data Analysis, edited by David C. Hoaglin, Frederick Mosteller, and John W. Tukey, New York: Wiley, 2000, ISBN 0-471-38491-7, xx + 445 pp., $69.95. Technometrics, 43(3):384, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.allbusiness.com/technology/808121-1.html;   http:/
/www.jstor.org/stable/1271251. Special Tukey Memorial is-

sue.

Ziliak:2008:CSS

[ZM08] Stephen Thomas Ziliak and Deirdre N. McCloskey. The Cult of Statistical Significance: How the Standard Error Costs Us Jobs, Justice, and Lives. Economics, cognition, and society. University of Michigan Press, Ann Arbor, MI, USA, 2008. ISBN 0-472-07007-X (cloth), 0-472-05007-9 (paperback). xxiii + 321 pp. LCCN HB137

.Z55 2008. URL http://www.loc.gov/catdir/enhancements/
fy0809/2007035401-d.html.